### v1

# Boundedness in general type MMP and fivefold effective termination

Abstract.
We show that in any sequence of a general type MMP, the minimal log discrepancy of singularities takes at most finitely many values, and the fibers of all the extremal contractions and flips belong to a bounded family. As an application, we prove the effective termination of general type MMPs in dimension at most $5$. A key ingredient in the proof is an analysis of the behavior of local volumes in the MMP.
Throughout this paper, we work over $\mathbb{C}$, the field of complex numbers.
1. Introduction
The Minimal Model Program (MMP) aims to construct, for any projective variety, a birational model that is as simple as possible. One of the main open problems in the MMP is the termination of the program. Conjecturally, every sequence of MMP terminates after finitely many steps; in particular, most natural invariants of the variety are expected to be bounded in any given MMP sequence. The purpose of this article is to prove some boundedness statements of this kind, without assuming termination, for varieties of general type, or more generally, for any MMP that starts with a Kawamata log terminal (klt) pair $(X,\Delta)$ such that either $K_{X}+\Delta$ or $\Delta$ is big. As applications, we prove some effective and explicit termination results in dimension at most five.
1.1. Boundedness
Our first result concerns the boundedness of the index and minimal log discrepancy. Recall that the minimal log discrepancy (mld), denoted $\mathrm{mld}(x,X,\Delta)$, of a klt pair $(X,\Delta)$ at a point $x\in X$ is the smallest log discrepancy of prime divisors over $x\in X$. It is one of the most important invariants in the study of the MMP, partly due to its monotonicity under the MMP.
Theorem 1.1 (see Theorem 3.1).
Let $(X,\Delta)$ be a projective $\mathbb{Q}$-factorial klt pair such that either $K_{X}+\Delta$ or $\Delta$ is big. Then there exist a positive integer $r$ and a finite set $S\subseteq\mathbb{R}_{+}$, depending only on the pair $(X,\Delta)$, such that for any sequence of steps of a $(K_{X}+\Delta)$-MMP $(X,\Delta)\dashrightarrow(Y,\Delta_{Y})$,
  1. (1)
    the Cartier index of any Weil divisor on $Y$ is at most $r$, and
  2. (2)
    for any (not necessarily closed) point $y\in Y$, we have $\mathrm{mld}(y,Y,\Delta_{Y})\in S$.
Shokurov [Sho-mld-conj] has shown that if the minimal log discrepancy of singularities that appear in the MMP satisfies both lower semi-continuity (LSC) and ascending chain condition (ACC), then the MMP must terminate after finitely many steps. Theorem 1.1 confirms the ACC conjecture for minimal log discrepancies in general type MMPs, thus verifies this part of Shokurov’s approach. For some past works on the ACC conjecture in special cases, see for example [Ale93, Amb06, Kaw14, Jia21, HLL22, NS22, HLS24, Kaw24, HL25].
Our second result shows that when running the MMP on varieties of general type (or when the boundary is big), the fibers of all possible extremal contractions and flips belong to a bounded family. This is certainly expected if the MMP terminates. Indeed, one can also show that a log canonical MMP terminates if and only if the pairs that appear in the MMP belong to a bounded family, see Appendix A.1.
Theorem 1.2 (see Theorem 4.1).
Let $(X,\Delta)$ be a projective $\mathbb{Q}$-factorial klt pair such that either $K_{X}+\Delta$ or $\Delta$ is big. Then there exists a projective family $\mathcal{W}\to\mathcal{B}$ over a finite type base $\mathcal{B}$, such that in any sequence of $(K_{X}+\Delta)$-MMP, every fiber of the extremal contractions or the flips is isomorphic to $\mathcal{W}_{b}$ for some $b\in\mathcal{B}$.
Recall that a flip $X\dashrightarrow X^{+}$ in the MMP is a diagram $X\rightarrow Z\leftarrow X^{+}$ where $X\to Z$ is the flipping contraction and $X^{+}\to Z$ is (also called) the flip [^1] In some literature it is called the flipped contraction.. We refer to the fiber of $X^{+}\to Z$ as the fiber of the flip.
More generally, our method can be used to show that in the setting of Theorem 1.2, any fixed order infinitesimal neighborhood of the fibers belong to a bounded family. This suggests that there are only finitely many “analytic types” of flips in the MMP, and most invariants of the pair can only change in finitely many different ways in the MMP.
1.2. Termination in dimension at most five
Termination of the MMP is known in full generality in dimension three by [Kaw-termination, Sho96]. Several special cases are also known in dimension four [KMM87, Fuj05, AHK07, Bir07, Sho09, Bir10, CT23], and in higher dimensions we have the termination of MMP with scaling (a special kind of MMP) in the general type or big boundary case [BCHM]. Beyond these, the termination conjecture remains extremely difficult. In particular, it is an open question whether every fivefold general type MMP terminates. A major application of our main results is the effective termination of general type MMPs, or more generally, of pseudo-effective MMP with a big boundary, in dimension five.
Theorem 1.3 (see Theorem 5.8).
Let $(X,\Delta)$ be a projective klt pair of dimension $5$. Assume that at least one of the following holds.
  1. (1)
    $K_{X}+\Delta$ is big.
  2. (2)
    $K_{X}+\Delta$ is pseudo-effective and $\Delta$ is big.
Then there exists a positive integer $m$ depending only on $(X,\Delta)$ such that every $(K_{X}+\Delta)$-MMP terminates after at most $m$ steps.
In dimensions $3$ and $4$, we also give some explicit bounds on the number of steps in the MMP. A sample result is the following. Here $h_{4}^{\mathrm{alg}}(X)$ is the dimension of the subspace in the singular homology group $H_{4}(X,\mathbb{R})$ generated by algebraic cycles.
Theorem 1.4 (see Theorems 6.15).
Let $N$ be a positive integer and $(X,\Delta)$ a projective log smooth klt pair of dimension $4$ such that $N\Delta$ has integer coefficients. Assume that
  1. (1)
    $(X,\Delta)$ has at most $N$ strata, each stratum has Picard number at most $N$, and $h_{4}^{\mathrm{alg}}(X)\leq N$, and
  2. (2)
    $\Delta$ is big, $\mathrm{vol}(\Delta)\geq\frac{1}{N}$, and $(\Delta\cdot H^{3})\leq N$ for some very ample divisor $H$ on $X$.
Then any $(K_{X}+\Delta)$-MMP terminates after at most $M^{M}$ steps, where $M:=(2N)^{9}!$.
By passing to log resolutions, Theorem 1.4 gives explicit termination results for klt MMP with a big boundary (see Lemma 2.15).
We remark that even the termination of such fourfold MMPs was not previously known. We also prove similar explicit termination results for general type fourfold MMPs (Theorem 6.14), terminal fourfold MMPs (Corollary 6.9) and arbitrary threefold MMPs (Theorem 6.16). For some past works on the explicit or effective termination problem for threefolds, see for example [CH11, Mar22, CC24, Kim25].
1.3. Local volume
A key player in our proof is the local volume of klt singularities, an invariant introduced by Chi Li [Li-normalized-volume] with motivation from questions in K-stability. A natural question is how the local volume behaves in the MMP. In general, this invariant does not seem to satisfy any monotonicity in the MMP, see Remark 3.6. Nonetheless, one of our observations is that the local volumes are bounded away from zero in the MMP, at least when the boundary is big. By standard argument (as in [BCHM]), any general type MMP can be turned into an MMP with a big boundary.
Theorem 1.5 (see Theorem 3.2).
Let $(X,\Delta)$ be a projective $\mathbb{Q}$-factorial klt pair such that $\Delta$ is big. Then there exists some $\varepsilon>0$ such that for any sequence of steps of a $(K_{X}+\Delta)$-MMP $(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ and any closed point $x^{\prime}\in X^{\prime}$, we have $\widehat{\rm vol}(x^{\prime},X^{\prime})\geq\varepsilon$.
Here $\widehat{\rm vol}(x^{\prime},X^{\prime})$ denotes the local volume of the klt singularity $x^{\prime}\in X^{\prime}$ (see Section 2.3). To prove this theorem, we further introduce the log canonical volume (see Section 3.2) of the big $\mathbb{R}$-divisor $\Delta$, defined as the largest volume of a graded linear series $V_{\bullet}$ of $\Delta$ such that $\left(X,\frac{1}{m}\Gamma\right)$ is log canonical for any positive integer $m$ and any $\mathbb{R}$-divisor $\Gamma$ in the linear system $|V_{m}|$. It turns out that the log canonical volume is always non-decreasing in the MMP (Lemma 3.11), and the local volume of singularities is always bounded from below by the log canonical volume (Lemma 3.12), which implies Theorem 1.5. On the other hand, the local volume controls many other invariants of the singularity. In particular, by the local K-stability theory of klt singularities (see [Z-SDC-survey] for a recent survey), the index and minimal log discrepancy of a klt singularity are bounded from above once the local volume is bounded away from zero, and Theorem 1.1 follows from this.
Local volume also plays an important role in the proof of Theorem 1.2. A basic idea is that if $(X^{\prime},\Delta^{\prime})\to Z$ is an extremal contraction in a $(K_{X}+\Delta)$-MMP and $\Delta$ is big, then the local volume of the relative cone (see Section 4.2 for the construction) of $(X^{\prime},\Delta^{\prime})$ over $Z$ is also bounded from below by the log canonical volume of $\Delta$ (up to a suitable multiplicative constant that only depends on the coefficients of $\Delta$), see Theorem 4.10. The local K-stability theory then implies that the relative cones are at least specially bounded (namely, they are bounded up to degenerations induced by special test configurations), though in this case a significant amount of work is still needed to deduce the boundedness in Theorem 1.2 from the above special boundedness.
1.4. Outline of the article
In Section 2 we summarize some preliminary results that will be used in the later part of this paper. In Section 3 we prove Theorems 1.1 and 1.5. In Section 4 we prove Theorem 1.2. In Section 5 we prove Theorem 1.3. In Section 6 we prove Theorem 1.4 and its variants. In the appendix, we prove that if the pairs that appear in a log canonical MMP are bounded, then the MMP terminates after finitely many steps.
Acknowledgment
The authors would like to thank Jungkai A. Chen, Christopher D. Hacon, James McKernan, Vyacheslav Shokurov, Chenyang Xu for helpful discussions and comments. JH is supported by NSFC for Excellent Young Scientists (#12322102), and the National Key R&D Program of China (#2023YFA1010600, #2020YFA0713200). JH is a member of LMNS, Fudan University. JL is supported by the National Key R&D Program of China (#2024YFA1014400). LQ is partially supported by Shanghai Sailing program 24YF2709800. ZZ is partially supported by the NSF Grants DMS-2240926, DMS-2234736, a Sloan research fellowship, a Packard fellowship, and the Simons Collaboration on Moduli of Varieties.
2. Preliminary
In this section, we collect some preliminary materials that will be used throughout the article. We suggest skimming through most of this section on first reading and returning to the relevant subsections for more details when certain definitions and statements are needed.
2.1. Notation and conventions
In this paper, varieties are assumed to be quasi-projective unless otherwise specified. We follow the standard terminology from [KM98, Kol13].
Given two $\mathbb{R}$-divisors $\Delta$ and $\Delta^{\prime}$ on a variety $X$, we write $\Delta^{\prime}\geq\Delta$ if $\mathrm{mult}_{D}\Delta^{\prime}\geq\mathrm{mult}_{D}\Delta$ for any prime divisor $D$ on $X$.
A sub-pair $(X,\Delta)$ consists of a normal variety $X$ together with an $\mathbb{R}$-divisor $\Delta$ on $X$ (a priori, we do not require that $K_{X}+\Delta$ is $\mathbb{R}$-Cartier). It is called a pair if $\Delta\geq 0$. We denote by $\mathrm{Coeff}(\Delta)$ the (finite) set of coefficients of $\Delta$ (in particular, we require that $0\in\mathrm{Coeff}(\Delta)$). A log (sub-)pair is a (sub-)pair $(X,\Delta)$ such that $K_{X}+\Delta$ is $\mathbb{R}$-Cartier. A singularity $x\in(X,\Delta)$ consists of a pair $(X,\Delta)$ and a closed point $x\in X$. If there is no confusion, we shall assume that $X$ is affine and $x$ is contained in any irreducible component of $\mathrm{Supp}(\Delta)$.
A sub-pair $(X,\Delta)$ is called log smooth if $X$ is smooth and $\mathrm{Supp}(\Delta)$ is a simple normal crossing divisor. A stratum (resp. open stratum) of a log smooth pair $(X,\Delta=\sum_{i\in I}b_{i}B_{i})$, where $B_{i}$ are the irreducible components of $\Delta$, is defined to be a connected component of $B_{J}:=\cap_{j\in J}B_{j}$ (resp. $B^{\circ}_{J}:=\cap_{j\in J}B_{j}\setminus\cup_{j\not\in J}B_{j}$) for some $J\subseteq I$ (by convention, $B_{\emptyset}=X$ and $B^{\circ}_{\emptyset}=X\setminus\mathrm{Supp}(\Delta)$).
Let $(X,\Delta)$ be a log sub-pair. A proper birational morphism $\pi\colon(Y,\Delta_{Y})\rightarrow(X,\Delta)$ from a sub-pair is called crepant if $K_{Y}+\Delta_{Y}=\pi^{*}(K_{X}+\Delta)$, in which case we also call $(Y,\Delta_{Y})$ the crepant pullback of $(X,\Delta)$ (via $\pi$).
An $\mathbb{R}$-divisor $D$ on a variety $X$ is big if $D=A+E$ for some ample $\mathbb{R}$-divisor $A$ and some $E\geq 0$. Let $X\to Z$ be a morphism. We say $D$ is big over $Z$ if $D\sim_{\mathbb{R},Z}D^{\prime}$ for some big $\mathbb{R}$-divisor $D^{\prime}$ on $X$.
Let $f\colon X\to Z$ be a morphism and $z\in Z$ a point. We denote by $X_{z}:=f^{-1}(z)$ the (scheme-theoretic) fiber over $z$.
An $\mathbb{R}$-divisor $D$ on a projective variety $X$ is pseudo-effective if its numerical class is a limit of $\mathbb{R}$-divisors $D_{i}\geq 0$ ($i=1,2,\dots$). Given a projective morphism $X\to Z$, we say $D$ is pseudo-effective over $Z$ if its restriction to the fiber $X_{\eta}$ is pseudo-effective, where $\eta\in Z$ is the image of the generic point of $X$.
A test configuration of a pair $(X,\Delta)$ consists of the following data:
  1. (1)
    A pair $(\mathcal{X},\mathcal{D})$ and a flat morphism $\pi\colon\mathcal{X}\to\mathbb{A}^{1}$.
  2. (2)
    A $\mathbb{G}_{m}$-action on $(\mathcal{X},\mathcal{D})$ lifting the canonical multiplication action of $\mathbb{G}_{m}$ on $\mathbb{A}^{1}$.
  3. (3)
    An isomorphism $(\mathcal{X}_{t},\mathcal{D}_{t})\cong(X,\Delta)$ for all $0\neq t\in\mathbb{A}^{1}$.
A fibration is a proper surjective morphism $f\colon X\to Z$ of normal varieties such that $f_{*}\mathcal{O}_{X}=\mathcal{O}_{Z}$. A fibration germ, denoted by $f\colon X\to Z\ni z$, consists of a fibration $f\colon X\to Z$ where $Z$ is affine and a closed point $z\in Z$. If $X\to Z$ is a fibration and $(X,\Delta)$ is a pair, by abuse of notation we also call $(X,\Delta)\to Z$ a fibration.
Let $f:X\dashrightarrow Y$ be a birational map and let $D$ be an $\mathbb{R}$-Cartier $\mathbb{R}$-divisor on $Y$. The birational pullback $f^{*}D$ is defined as $p_{*}q^{*}D$ where $p:Z\to X$, $q:Z\to Y$ is a resolution of the indeterminacy of $f$ (one can check that the definition is independent of the resolution).
Let $X$ be a variety. For any proper birational morphism $\pi\colon Y\to X$ and any prime divisor $E$ on $Y$, we call $E$ a (prime) divisor over $X$ and denote by $\mathrm{Center}_{X}(E):=\pi(E)$ the center of $E$ on $X$. We also denote by $c_{X}(E)$ the generic point of $\mathrm{Center}_{X}(E)$ and sometimes also call it the center of $E$ if no confusion arises. We also call $E$ as a (prime) divisor over $x\in X$ where $x$ is the generic point of $\pi(E)$. We denote by $\mathrm{Ex}(\pi)\subseteq Y$ the exceptional locus of $\pi$. The discrepancy of $E$ with respect to a log sub-pair $(X,\Delta)$, denoted by $a(E,X,\Delta)$, is defined by
$a(E,X,\Delta):=\mathrm{mult}_{E}(K_{Y}-\pi^{*}(K_{X}+\Delta)).$
The log discrepancy of $E$ with respect to $(X,\Delta)$ is defined by $A_{X,\Delta}(E):=1+a(E,X,\Delta)$.
The notation of terminal, klt, plt, dlt, and log canonical (sub-)pairs are defined as in [KM98, Definitions 2.34 and 2.37]. The non-klt locus of a log pair $(X,\Delta)$ is defined as
$\mathrm{Nklt}(X,\Delta):=\{x\in X\,|\,(X,\Delta)\mbox{ is not klt near }x\}.$
Let $(X,\Delta)$ be a log canonical pair. A subvariety $W\subseteq X$ is called a log canonical center if either $W=X$, or $W$ is the center of some divisor $E$ over $X$ with $A_{X,\Delta}(E)=0$.
2.2. Log discrepancy
In this and the next two subsections, we recall some invariants of pairs and their basic properties.
Definition 2.1.
For any log canonical sub-pair $(X,\Delta)$ and a (not necessarily closed) point $x\in X$, we define
$\mathrm{mld}(x,X,\Delta):=\inf\{A_{X,\Delta}(E)\mid E\text{ is a prime divisorover }x\in X\}$
as the minimal log discrepancy (mld) of $(X,\Delta)$ at $x$, and define
$\mathrm{mld}(X,\Delta):=\inf\{A_{X,\Delta}(E)\mid E\text{ is a prime divisor over }X\}$
as the total minimal log discrepancy (total mld) of $(X,\Delta)$.
Lemma 2.2.
Let $n$ and $r$ be two positive integers and $S_{0}\subseteq[0,1]$ a finite set. Then there exists a discrete set $S\subseteq\mathbb{R}_{\geq 0}$ depending only on $n,r$ and $S_{0}$ satisfying the following. Assume that
  1. (1)
    $(X,\Delta)$ is a dlt pair of dimension $n$,
  2. (2)
    $\mathrm{Coeff}(\Delta)\subseteq S_{0}$,
  3. (3)
    $rL$ is Cartier for any $\mathbb{Q}$-Cartier Weil divisor $L$ on $X$, and
  4. (4)
    $W$ is a log canonical center of $(X,\Delta)$ of dimension $\geq 1$ and $(W,\Delta_{W})$ is the dlt pair induced by adjunction $K_{W}+\Delta_{W}:=(K_{X}+\Delta)|_{W}.$
Then for any prime divisor $E$ over $W$, we have $A_{W,\Delta_{W}}(E)\in S$.
Proof.
By [HLS24]*Theorem 5.6, there exist real numbers $a_{1},\dots,a_{k}\in(0,1]$ and a positive integer $N$ depending only on $n$ and $S_{0}$, and $\mathbb{Q}$-divisors $\Delta_{1},\dots,\Delta_{k}\geq 0$ such that $\sum_{i=1}^{k}a_{i}=1$, $\Delta=\sum_{i=1}^{k}a_{i}\Delta_{i}$, each pair $(X,\Delta_{i})$ is dlt, and $N\Delta_{i}$ is a Weil divisor for each $i$. In particular, $W$ is a log canonical center of each $(X,\Delta_{i})$. Let $K_{W}+\Delta_{W,i}:=(K_{X}+\Delta_{i})|_{W}$ for each $i$, then each $(W,\Delta_{W,i})$ is dlt and $\Delta_{W}=\sum_{i=1}^{k}a_{i}\Delta_{W,i}$. For each $i$, since $Nr(K_{X}+\Delta_{i})$ is Cartier by assumption (3), we have $Nr(K_{W}+\Delta_{W,i})$ is also Cartier. Thus the set
$S=\frac{1}{Nr}\left\{\sum_{i=1}^{k}m_{i}a_{i}\middle|\,m_{i}\in\mathbb{N}\right\}$
satisfies the statement of the lemma. ∎
2.3. Local volume
Let $X$ be a variety and $x\in X$ a closed point. A valuation on the function field $\mathbb{C}(X)$ of a variety $X$, also called a valuation on $X$, is an $\mathbb{R}$-valued function $v:\mathbb{C}(X)^{*}\rightarrow\mathbb{R}$ such that $v|_{\mathbb{C}^{*}}=0$, $v(ab)=v(a)+v(b)$ for all $a,b\in\mathbb{C}(X)^{*}$, and $v(a+b)\geq\min\{v(a),v(b)\}$ if $a+b\neq 0$. We say that $v$ is a valuation over $x\in X$ if $v$ is centered at $x$. The set of valuations on $X$ is denoted by $\mathrm{Val}_{X}$. The set of valuations over $x\in X$ is denoted by $\mathrm{Val}_{X,x}$.
Given a klt sub-pair $(X,\Delta)$, the log discrepancy function
$A_{X,\Delta}\colon\mathrm{Val}_{X}\to\mathbb{R}\cup\{+\infty\},$
is defined as in [JM-val-ideal-seq] and [BdFFU-log-discrepancy]*Theorem 3.1. For a divisorial valuation $v=\lambda\cdot\mathrm{ord}_{E}$ (where $\lambda>0$ and $E$ is a prime divisor over $X$), one has $A_{X,\Delta}(v)=\lambda\cdot A_{X,\Delta}(E)$. Since $(X,\Delta)$ is klt, $A_{X,\Delta}(v)>0$ for all $v\in\mathrm{Val}_{X,x}$. However, it is possible that $A_{X,\Delta}(v)=+\infty$ for some $v\in\mathrm{Val}_{X}$, see e.g. [JM-val-ideal-seq]*Remark 5.12.
The volume of a valuation $v\in\mathrm{Val}_{X,x}$ is defined as
$\mathrm{vol}(v)=\mathrm{vol}_{X,x}(v)=\limsup_{m\to+\infty}\frac{\ell(\mathcal{O}_{X,x}/\mathfrak{a}_{m}(v))}{m^{n}/n!},$
where $n=\dim X$ and $\mathfrak{a}_{m}(v)$ denotes the valuation ideal, i.e.
$\mathfrak{a}_{m}(v):=\{f\in\mathcal{O}_{X,x}\mid v(f)\geq m\}.$
Thanks to the works of [ELS03, LM-Okounkov-body, Cut13], the above limsup is actually a limit.
Definition 2.3.
Let $(X,\Delta)$ be a klt pair of dimensional $n$ and $x\in X$ a closed point. For any $v\in\mathrm{Val}_{X,x}$, we define the normalized volume of $v$ as
$\widehat{\rm vol}_{X,\Delta}(v):=\begin{cases}A_{X,\Delta}(v)^{n}\cdot\mathrm{vol}_{X,x}(v),&\mbox{if }A_{X,\Delta}(v)<+\infty,\\ +\infty,&\mbox{if }A_{X,\Delta}(v)=+\infty.\end{cases}$
The local volume of $(X,\Delta)$ at $x$ is defined as
$\widehat{\rm vol}(x,X,\Delta):=\inf_{v\in\mathrm{Val}_{X,x}}\widehat{\rm vol}_{X,\Delta}(v).$
By [Li-normalized-volume]*Theorem 1.2, the local volume of a klt singularity is always positive. By [Blu-minimizer-exist], there always exists some valuation $v$ such that $\widehat{\rm vol}(x,X,\Delta)=\widehat{\rm vol}_{X,\Delta}(v)$. We also need the following properties of the local volume.
Theorem 2.4 ([XZ-minimizer-unique]*Corollary 1.4).
Let $(X,\Delta)$ be a klt pair of dimensional $n$, let $x\in X$ be a closed point, and let $D$ be a $\mathbb{Q}$-Cartier Weil divisor on $X$. Then $rD$ is Cartier near $x$ for some positive integer $r\leq\frac{n^{n}}{\widehat{\rm vol}(x,X,\Delta)}$.
Theorem 2.5.
Fix $n\in\mathbb{N}$ and $\varepsilon>0$. Then there exists some constant $A>0$ depending only on $n$ and $\varepsilon$, such that for any $n$-dimensional klt pair $(X,\Delta)$ satisfying $\widehat{\rm vol}(x,X,\Delta)\geq\varepsilon$ for all closed points $x\in X$, we have
$\mathrm{mld}(x,X,\Delta)\leq A$
for all (not necessarily) closed points $x\in X$.
Proof.
By [Amb-mld]*Proposition 2.3, it suffices to prove that $\mathrm{mld}(x,X,\Delta)\leq A$ for all closed point $x\in X$. We are done by [XZ-local-bdd, Theorem 1.3] or [HLQ24]*Corollary 2.7. ∎
2.4. Log canonical threshold
For any log canonical sub-pair $(X,\Delta)$ and any $\mathbb{R}$-Cartier $\mathbb{R}$-divisor $D$ on $X$, we denote by
$\mathrm{lct}(X,\Delta;D):=\sup\{\,t\,|\,(X,\Delta+tD)\text{ is log canonical}\,\}\in\mathbb{R}\cup\{+\infty\}$
the log canonical threshold of $D$ with respect to $(X,\Delta)$. For any $x\in X$, we also denote by
$\mathrm{lct}_{x}(X,\Delta;D):=\sup\{\,t\,|\,(X,\Delta+tD)\text{ is log canonical at }x\,\}$
the log canonical threshold at $x$. The following result shows how the log canonical thresholds behave when we perturb the boundary.
Lemma 2.6.
Let $(X,\Delta)$ be a klt pair and let $D,D^{\prime}\geq 0$ be $\mathbb{R}$-Cartier $\mathbb{R}$-divisors on $X$. Then
$\mathrm{lct}(X,\Delta;D+D^{\prime})\geq\left(\frac{1}{\mathrm{lct}(X,\Delta;D)}+\frac{1}{\mathrm{lct}(X,\Delta;D^{\prime})}\right)^{-1}.$
If $(X,\Delta+D^{\prime})$ is log canonical, then we also have
$\mathrm{lct}(X,\Delta+D^{\prime};D)\geq\left(1-\frac{1}{\mathrm{lct}(X,\Delta;D^{\prime})}\right)\cdot\mathrm{lct}(X,\Delta;D).$
Proof.
Let $a:=\mathrm{lct}(X,\Delta;D)$ and $a^{\prime}:=\mathrm{lct}(X,\Delta;D^{\prime})$. Then $(X,\Delta+aD)$ and $(X,\Delta+a^{\prime}D^{\prime})$ are log canonical, so the convex combination
$\left(X,\Delta+\frac{a^{\prime}}{a+a^{\prime}}\cdot aD+\frac{a}{a+a^{\prime}}\cdot a^{\prime}D^{\prime}\right)=\left(X,\Delta+\frac{aa^{\prime}}{a+a^{\prime}}(D+D^{\prime})\right)$
is log canonical, which gives the first inequality $\mathrm{lct}(X,\Delta;D+D^{\prime})\geq\frac{aa^{\prime}}{a+a^{\prime}}=\left(\frac{1}{a}+\frac{1}{a^{\prime}}\right)^{-1}$. If $a^{\prime}\geq 1$, then the convex combination
$\left(X,\Delta+\frac{a^{\prime}-1}{a^{\prime}}\cdot aD+\frac{1}{a^{\prime}}\cdot a^{\prime}D^{\prime}\right)=\left(X,\Delta+D^{\prime}+\frac{a^{\prime}-1}{a^{\prime}}\cdot aD\right)$
is log canonical, which implies the second inequality $\mathrm{lct}(X,\Delta+D^{\prime};D)\geq\left(1-\frac{1}{a^{\prime}}\right)a$. ∎
The next statement is a version of Izumi’s inequality and should be well-known to experts. It will be useful in proving explicit termination results in Section 6.
Lemma 2.7.
Let $\varepsilon\in[0,1]$, let $(X,\Delta)$ be a log smooth sub-pair such that the coefficients of $\Delta$ are at most $1-\varepsilon$, and let $D\geq 0$ be an $\mathbb{R}$-divisor on $X$. Then for any closed point $x\in\mathrm{Supp}(D)$, we have
$\mathrm{lct}_{x}(X,\Delta;D)\geq\frac{\varepsilon}{\mathrm{mult}_{x}D}.$
Proof.
By removing the components of $\Delta$ with negative coefficients, we obtain an $\mathbb{R}$-divisor $\Delta^{\geq 0}$ with $\mathrm{lct}_{x}(X,\Delta;D)\geq\mathrm{lct}_{x}(X,\Delta^{\geq 0};D)$. Possibly replacing $\Delta$ with $\Delta^{\geq 0}$, it suffices to prove the lemma when $\Delta\geq 0$. Possibly replacing $D$ with a multiple, we may also assume that $\mathrm{mult}_{x}D=1$. We need to show that $x\in(X,\Delta+\varepsilon D)$ is log canonical.
Let $\pi\colon\widetilde{X}\to X$ be the blowup of $x$ and let $E$ be the exceptional divisor. Let $\widetilde{\Delta}$ (resp. $\widetilde{D}$) be the strict transform of $\Delta$ (resp. $D$) on $\tilde{X}$. Since $\mathrm{mult}_{x}D=1$, we have $\widetilde{D}|_{E}\sim_{\mathbb{R}}H$ where $H$ is the hyperplane class on $E$ under the natural isomorphism $E\cong\mathbb{P}^{\dim X-1}$. Note that $(E,\widetilde{\Delta}|_{E})$ is a toric pair and the coefficients of $\widetilde{\Delta}|_{E}$ are at most $1-\varepsilon$, thus $\mathrm{lct}(E,\widetilde{\Delta}|_{E};G)\geq\varepsilon$ for any torus invariant $\mathbb{R}$-divisor $0\leq G\sim_{\mathbb{R}}H$ on $E$. By [BJ-delta]*Theorem F, we deduce that $\mathrm{lct}(E,\widetilde{\Delta}|_{E};G)\geq\varepsilon$ for any $\mathbb{R}$-divisor $0\leq G\sim_{\mathbb{R}}H$ on $E$. In particular, the pair $(E,(\widetilde{\Delta}+\varepsilon\widetilde{D})|_{E})$ is log canonical. By inversion of adjunction (cf. [KM98]*Theorem 5.50), the pair $(\widetilde{X},\widetilde{\Delta}+\varepsilon\widetilde{D}+E)$ is log canonical near $E$. Since the coefficients of $\Delta$ are at most $1-\varepsilon$, we have $A_{X,\Delta}(E)\geq\varepsilon$, hence $A_{X,\Delta+\varepsilon D}(E)=A_{X,\Delta}(E)-\varepsilon\cdot\mathrm{mult}_{x}D\geq 0$. Then $K_{\widetilde{X}}+\widetilde{\Delta}+\varepsilon\widetilde{D}+E\geq\pi^{*}(K_{X}+\Delta+\varepsilon D)$, so $(X,\Delta+\varepsilon D)$ is log canonical at $x$. ∎
2.5. Linear series
For any $\mathbb{R}$-divisor $D$ on a normal variety $X$, we set
$H^{0}(X,D):=\{0\neq s\in\mathbb{C}(X)\,|\,\mathrm{div}(s)+D\geq 0\}\cup\{0\}.$
Any finite dimensional subspace $V$ of $H^{0}(X,D)$ is called a linear series of $D$. It defines a linear system
[…]
A graded linear series $V_{\bullet}$ of $D$ is a sequence of finite dimensional subspaces $V_{m}\subseteq H^{0}(X,mD)$ (one for each $m\in\mathbb{N}$) such that $V_{m}\cdot V_{l}\subseteq V_{m+l}$ for all $m,l\in\mathbb{N}$. Its volume is defined as
(2.1) $\mathrm{vol}(V_{\bullet}):=\limsup_{m\to+\infty}\frac{\dim(V_{m})}{m^{n}/n!}$
where $n=\dim X$. We denote by $M(V_{\bullet})\subseteq\mathbb{N}^{+}$ the semigroup of those positive integers $m$ with $V_{m}\neq 0$. If $D$ is $\mathbb{R}$-Cartier, and $\Delta\geq 0$ is an $\mathbb{R}$-divisor such that $(X,\Delta)$ is klt, we also define the alpha invariant of $V_{\bullet}$ as
$\alpha(X,\Delta;V_{\bullet}):=\inf\left\{\,\mathrm{lct}(X,\Delta;\Gamma)\,\middle|\ m\in M(V_{\bullet}),\,\Gamma\in\frac{1}{m}|V_{m}|\,\right\}$
and simply write $\alpha(V_{\bullet})$ if the klt pair $(X,\Delta)$ is clear from the context. When $V_{\bullet}$ is the complete linear series of $D$, the corresponding alpha invariant is denoted by $\alpha(X,\Delta;D)$.
We say that a graded linear series $V_{\bullet}$ is eventually birational, if the rational map $X\dashrightarrow\mathbb{P}(V^{\vee}_{m})$ induced by the linear system $|V_{m}|$ is birational onto its image for all sufficiently large $m\in M(V_{\bullet})$. This condition is preserved under birational contractions; by the following lemma, it also ensures that the limsup in (2.1) is a limit. For this reason, in the rest of this paper we will exclusively consider graded linear series that are eventually birational.
Lemma 2.8.
Let $(X,\Delta)$ be a klt pair and let $V_{\bullet}$ be an eventually birational graded linear series such that $\alpha(V_{\bullet})>0$. Then $\mathrm{vol}(V_{\bullet})>0$ and we have
(2.2) $\mathrm{vol}(V_{\bullet})=\lim_{M(V_{\bullet})\ni m\to+\infty}\frac{\dim(V_{m})}{m^{n}/n!}.$
Proof.
We refer to [LM-Okounkov-body] for the construction of Okounkov bodies and their general properties. It is clear that $\mathrm{vol}(V_{\bullet})>0$ as $|V_{m}|$ defines a birational map for large enough $m$. To prove (2.2), let $\pi\colon X^{\prime}\to X$ be the blowup of some smooth point $x\in X$ and consider the Okounkov body $\Delta(V_{\bullet})$ constructed using a full flag of subvarieties of the form $X^{\prime}\supseteq E\supseteq\cdots\supseteq\{\mathrm{pt\}}$, where $E$ is the exceptional divisor of the blowup. Since $\alpha(V_{\bullet})>0$, there exists some constant $C>0$ such that $\mathrm{mult}_{E}(\pi^{*}\Gamma)=\mathrm{mult}_{x}(\Gamma)\leq Cm$ for all $m\in\mathbb{N}$ and $\Gamma\in|V_{m}|$. As $E\cong\mathbb{P}^{n-1}$ is projective, by [LM-Okounkov-body]*Lemma 1.11 we know that $\Delta(V_{\bullet})$ satisfies condition (A) from [LM-Okounkov-body]*Definition 2.4. Since $V_{\bullet}$ is eventually birational, it also satisfies condition (B) from [LM-Okounkov-body]*Definition 2.5. By [LM-Okounkov-body]*Theorem 2.13, we see that
$\lim_{M(V_{\bullet})\ni m\to+\infty}\frac{\dim(V_{m})}{m^{n}/n!}$
exists and equals $\mathrm{vol}(V_{\bullet})$. ∎
We also need the following Bertini type result.
Lemma 2.9.
Let $(X,\Delta)$ be a klt pair and let $H$ be an ample $\mathbb{Q}$-divisor on $X$. Then there exist some $\delta>0$ and some $m_{0}\in\mathbb{N}$ such that for any positive integer $m\geq m_{0}$, there exists some $D\in|mH|$ with
$\mathrm{lct}(X,\Delta;D)\geq\delta.$
Proof.
Let $m_{0}$ be a sufficiently large and divisible integer such that $|m_{0}H|$ is base point free and $|mH|\neq\emptyset$ for all $m\geq m_{0}$. For each $i=0,1,\dots,m_{0}-1$, choose some $D_{i}\in|(m_{0}+i)H|$ and let
$\delta:=\min_{0\leq i\leq m_{0}-1}\{\mathrm{lct}(X,\Delta;D_{i})\}>0.$
For any integer $m\geq m_{0}$ we may write $m=(\ell+1)m_{0}+i$ for some $\ell\in\mathbb{N}$ and some $10\leq i\leq m_{0}-$ 1. By our choice of $m_{0}$ and $\delta$, we know that $|\ell m_{0}H|$ is base point free and $(X,\Delta+\delta D_{i})$ is log canonical, thus by Bertini’s theorem, there exists some $D^{\prime}\in|\ell m_{0}H|$ such that $(X,\Delta+\delta D_{i}+D^{\prime})$ is still log canonical. Note that $\delta\leq 1$. Then $D=D_{i}+D^{\prime}\in|mH|$ satisfies $\mathrm{lct}(X,\Delta;D)\geq\delta$. ∎
2.6. Minimal Model Program (MMP)
We refer to e.g. [KM98, Paragraph 3.31] for a description of the MMP. In this subsection, we recall some results on the MMP that will be used in this paper. First we recall the two basic operations in the MMP.
Definition 2.10.
Let $(X,\Delta)$ be a $\mathbb{Q}$-factorial log canonical pair and let $f\colon X\longrightarrow Z$ be a fibration. Then $f$ is called a $(K_{X}+\Delta)$-flipping contraction if
  1. (1)
    $f$ is small birational,
  2. (2)
    the relative Picard number $\rho(X/Z)=1$, and
  3. (3)
    $-(K_{X}+\Delta)$ is $f$-ample.
The $(K_{X}+\Delta)$-flip $f^{+}\colon X^{+}\longrightarrow Z$ of the flipping contraction $f\colon X\longrightarrow Z$ is a small birational fibration $f^{+}\colon X^{+}\longrightarrow Z$ such that $K_{X^{+}}+\Delta^{+}$ is $f^{+}$-ample, where $\Delta^{+}$ is the strict transform of $\Delta$ on $X^{+}$. When the pair $(X,\Delta)$ is clear from the context, we shall refer to $f$ (resp. $f^{+}$) as simply the flipping contraction (resp. the flip). By abuse of terminology, we also call the composition $(f^{+})^{-1}\circ f\colon X\dashrightarrow X^{+}$ a flip (however, the fibers of a flip only refer to the fibers of $f^{+}$).
Definition 2.11.
Let $(X,\Delta)$ be a $\mathbb{Q}$-factorial log canonical pair and let $f\colon X\longrightarrow Z$ be a projective morphism of normal varieties. Then $f$ is called a divisorial contraction if
  1. (1)
    $f$ is birational with exceptional locus a divisor,
  2. (2)
    the relative Picard number $\rho(X/Z)=1$, and
  3. (3)
    $-(K_{X}+\Delta)$ is $f$-ample.
For some of our proofs, it will be convenient to work with a broader class of birational maps as follows.
Definition 2.12.
A birational map $\varphi\colon(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ of log pairs is called an MMP type contraction if $\varphi^{-1}$ has no exceptional divisor, and there exist proper birational morphisms $f\colon Y\to X$, $g\colon Y\to X^{\prime}$ such that
$f^{*}(K_{X}+\Delta)-g^{*}(K_{X^{\prime}}+\Delta^{\prime})\geq 0$
and $g=\varphi\circ f$. In particular, we have $\varphi_{*}\Delta\geq\Delta^{\prime}$.
Observe that if $\varphi\colon(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ is an MMP type contraction and $(X,\Delta)$ is klt (resp. log canonical), then so is $(X^{\prime},\Delta^{\prime})$. As the name suggests, if $\varphi\colon(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ is obtained by running a $(K_{X}+\Delta)$-MMP, then $\varphi\colon(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ is an MMP type contraction. If $\Delta^{\prime}=\varphi_{*}\Delta$, then the birational contraction $\varphi\colon(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ is an MMP type contraction if and only if $\varphi$ is $(K_{X}+\Delta)$-non-positive in the sense of [BCHM]*Definition 3.6.1. We will also frequently use the following simple observation.
Lemma 2.13.
Let $\varphi\colon(X,\Delta+D)\dashrightarrow(X^{\prime},\Delta^{\prime}+D^{\prime})$ be an MMP type contraction with $\Delta,\Delta^{\prime},D\geq 0$ and $D^{\prime}=\varphi_{*}D$. Let $\Gamma\geq 0$ be an $\mathbb{R}$-divisor on $X$ such that $K_{X}+\Gamma\sim_{\mathbb{R}}\mu(K_{X}+D)$ for some $\mu\geq 0$, and let $\Gamma^{\prime}=\varphi_{*}\Gamma$. Then $\varphi$ also induces an MMP type contraction $\varphi:(X,\mu\Delta+\Gamma)\dashrightarrow(X^{\prime},\mu\Delta^{\prime}+\Gamma^{\prime})$.
Proof.
Let $f:Y\rightarrow X,g:Y\rightarrow X^{\prime}$ be two proper birational morphisms such that
$E:=f^{*}(K_{X}+\Delta+D)-g^{*}(K_{X^{\prime}}+\Delta^{\prime}+D^{\prime})\geq 0$
and $g=\varphi\circ f$. So
(2.3) $\mu E=f^{*}(\mu K_{X}+\mu\Delta+\mu D)-g^{*}(\mu K_{X^{\prime}}+\mu\Delta^{\prime}+\mu D^{\prime})\geq 0.$
Since $K_{X}+\Gamma\sim_{\mathbb{R}}\mu(K_{X}+D)$ and $\varphi^{-1}$ does not contract any divisor,
$K_{X^{\prime}}+\Gamma^{\prime}-\mu(K_{X^{\prime}}+D^{\prime})=g_{*}f^{*}(K_{X}+\Gamma-\mu(K_{X}+D))\sim_{\mathbb{R}}0.$
Thus by the negativity lemma, we have
(2.4) $f^{*}(K_{X}+\Gamma-\mu(K_{X}+D))-g^{*}(K_{X^{\prime}}+\Gamma^{\prime}-\mu(K_{X^{\prime}}+D^{\prime}))=0,$
as the left hand side is $g$-exceptional and $\sim_{\mathbb{R}}0$. Combining (2.3) and (2.4), we get
$\mu E=f^{*}(K_{X}+\mu\Delta+\Gamma)-g^{*}(K_{X^{\prime}}+\mu\Delta^{\prime}+\Gamma^{\prime})\geq 0.$
Hence $\varphi:(X,\mu\Delta+\Gamma)\dashrightarrow(X^{\prime},\mu\Delta^{\prime}+\Gamma^{\prime})$ is an MMP type contraction. ∎
We also introduce a more flexible version of the MMP, where each step of the MMP is a contraction of an extremal face rather than an extremal ray. To avoid confusion, we call such MMP as “face-contracting MMP”. Face-contracting MMP naturally appears in the proofs of the special termination (cf. [Fuj07a]). The work [Kol21] also uses special face-contracting MMP in order to ensure that the scaling numbers strictly drop. We remark that projectivity and $\mathbb{Q}$-factoriality of the ambient variety are not automatically presumed in the face-contracting MMP.
Definition 2.14 (Face-contracting MMP).
Let $(X,\Delta)$ be a log pair. A (birational) face-contracting $(K_{X}+\Delta)$-MMP is a sequence of birational contractions
[…]
where $(X_{i},\Delta_{i})$ is a log pair and $\Delta_{i+1}:=(\varphi_{i})_{*}\Delta_{i}$ for each $i$, and each $\varphi_{i}$ fits into a diagram
[…]
satisfying the following:
  1. (1)
    $f_{i}$ and $f_{i}^{+}$ are birational fibrations,
  2. (2)
    $f_{i}^{+}$ is small,
  3. (3)
    $-(K_{X_{i}}+\Delta_{i})$ is $f_{i}$-ample, and
  4. (4)
    $K_{X_{i+1}}+\Delta_{i+1}$ is $f_{i}^{+}$-ample.
We show that every step of a face-contracting MMP is crepant equivalent to a sequence of divisorial contractions and flips.
Lemma 2.15.
Let $(X,\Delta)$ be a log canonical pair and let $\pi\colon(Y,D)\to(X,\Delta)$ be a birational morphism from a $\mathbb{Q}$-factorial dlt pair such that $K_{Y}+D=\pi^{*}(K_{X}+\Delta)+E$ for some $\pi$-exceptional $\mathbb{R}$-divisor $E\geq 0$. Then any face-contracting $(K_{X}+\Delta)$-MMP
[…]
can be lifted to a commutative diagram
[…]
where each $(Y_{i},D_{i})$ is a $\mathbb{Q}$-factorial dlt pair and each $\psi_{i}$ is a sequence of $(K_{Y_{i}}+D_{i})$-MMP. Moreover, for all $i\geq 1$ we have $K_{Y_{i}}+D_{i}=\pi_{i}^{*}(K_{X_{i}}+\Delta_{i})$.
Proof.
Let $f_{i}:X_{i}\rightarrow Z_{i}$ and $f_{i}^{+}:X_{i+1}\rightarrow Z_{i}$ be the birational fibrations associated to $\varphi_{i}$. We prove the lemma by induction on $i$. When $i=0$, we simply set $(Y_{0},D_{0}):=(Y,D)$ and $\pi_{0}:=\pi$. For $i\geq 1$, suppose that $(Y_{i-1},D_{i-1})$ has been constructed. By [KM98]*Lemma 3.38 and [BCHM]*Lemma 3.6.6(4), $(X_{i},\Delta_{i})$ is the ample model over $Z_{i-1}$ of $(Y_{i-1},D_{i-1})$. In particular, $(Y_{i-1},D_{i-1})$ has a minimal model over $Z_{i-1}$. Since the pair $(Y_{i-1},D_{i-1})$ is $\mathbb{Q}$-factorial dlt by induction hypothesis, by [Bir12lcflip]*Corollary 1.2 and Theorem 1.9 or [HX13, Corollaries 1.8 and 2.9], we may run a $(K_{Y_{i-1}}+D_{i-1})$-MMP over $Z_{i-1}$ with scaling of an ample divisor, and the MMP terminates with a $\mathbb{Q}$-factorial dlt minimal model over $Z_{i-1}$. This gives $\psi_{i-1}:(Y_{i-1},D_{i-1})\dashrightarrow(Y_{i},D_{i})$, where $K_{Y_{i}}+D_{i}$ is nef over $Z_{i-1}$. Then $(X_{i},\Delta_{i})$ is the ample model over $Z_{i-1}$ of $(Y_{i},D_{i})$, so there exists an induced birational morphism $\pi_{i}:Y_{i}\to X_{i}$ such that $K_{Y_{i}}+D_{i}=\pi_{i}^{*}(K_{X_{i}}+\Delta_{i})$. This finishes the induction. ∎
We can also lift an MMP to the terminalization at the price of possibly changing the terminal boundary during the process.
Definition 2.16.
A $\mathbb{Q}$-factorial terminalization of a klt pair $(X,\Delta)$ is a projective crepant birational morphism $\pi\colon(Y,\Delta_{Y})\rightarrow(X,\Delta)$ such that $(Y,\Delta_{Y})$ is a $\mathbb{Q}$-factorial terminal pair. We also say that $(Y,\Delta_{Y})$ is a $\mathbb{Q}$-factorial terminalization of $(X,\Delta)$.
Lemma 2.17.
Let $(X,\Delta)$ be a klt pair and let $(Y,D)$ be a $\mathbb{Q}$-factorial terminal pair with a projective birational morphism $\pi\colon Y\rightarrow X$ such that $K_{Y}+D=\pi^{*}(K_{X}+\Delta)+E$ for some $\pi$-exceptional $\mathbb{R}$-divisor $E\geq 0$. Then any face-contracting $(K_{X}+\Delta)$-MMP
[…]
can be lifted to a commutative diagram
[…]
where each $\psi_{i}$ is a sequence of $(K_{Y_{i}}+B_{i})$-MMP with $0\leq B_{i}\leq D_{i}$ defined by
(2.5) $\mathrm{mult}_{F}B_{i}=\max\{0,-a(F,X_{i+1},\Delta_{i+1})\}$
for all prime divisors $F$ on $Y_{i}$, the map $\psi_{i}$ does not contract any component of $B_{i}$, and $D_{i+1}=(\psi_{i})_{*}B_{i}$. Moreover, $\pi_{i}:(Y_{i},D_{i})\rightarrow(X_{i},\Delta_{i})$ is a $\mathbb{Q}$-factorial terminalization for every $i\geq 1$.
Proof.
The proof is similar to that of Lemma 2.15 and proceeds by induction on $i$. Let $f_{i}:X_{i}\rightarrow Z_{i}$ and $f_{i}^{+}:X_{i+1}\rightarrow Z_{i}$ be the associated birational fibrations of $\varphi_{i}$. When $i=0$, set $(Y_{0},D_{0}):=(Y,D)$ and $\pi_{0}:=\pi$. For $i\geq 1$, suppose that $(Y_{i-1},D_{i-1})$ has been constructed. Define $B_{i-1}$ by (2.5). Clearly $B_{i-1}\geq 0$. By the negativity lemma and the induction hypothesis that $K_{Y_{i-1}}+D_{i-1}\geq\pi_{i-1}^{*}(K_{i-1}+\Delta_{i-1})$, we have
$a(F,Y_{i-1},D_{i-1})\leq a(F,X_{i-1},\Delta_{i-1})\leq a(F,X_{i},\Delta_{i})$
for all divisors $F$ over $X$, hence $B_{i-1}\leq D_{i-1}$. By [KM98]*Lemma 3.38 and [BCHM]*Lemma 3.6.6(4), $(X_{i},\Delta_{i})$ is the ample model over $Z_{i-1}$ of $(Y_{i-1},B_{i-1})$. Thus as in the proof of Lemma 2.15, by [Bir12lcflip]*Theorem 1.9 we may run a $(K_{Y_{i-1}}+B_{i-1})$-MMP $\psi_{i-1}:(Y_{i-1},B_{i-1})\dashrightarrow(Y_{i},D_{i})$ over $Z_{i-1}$, where $D_{i}:=(\psi_{i-1})_{*}B_{i-1}$, such that $K_{Y_{i}}+D_{i}$ is nef over $Z_{i-1}$ and $(X_{i},\Delta_{i})$ is the ample model over $Z_{i-1}$ of $(Y_{i},D_{i})$, which gives the morphism $\pi_{i}$. If $F$ is a component of $B_{i-1}$, then we have
$\mathrm{mult}_{F}B_{i-1}=-a(F,X_{i},\Delta_{i})=-a(F,Y_{i},D_{i})\leq-a(F,Y_{i-1},B_{i-1})=\mathrm{mult}_{F}B_{i-1},$
thus the inequality in the middle is an equality; in particular, $F$ is not contracted by the MMP. By [KM98]*Corollary 3.43, $(Y_{i},D_{i})$ is terminal. We finish the induction. ∎
We will frequently use the following observation to reduce questions about general type MMPs to MMPs with a big boundary.
Lemma 2.18.
Let $(X,\Delta)$ be a klt pair and $X\rightarrow T$ a morphism.
  1. (1)
    If $K_{X}+\Delta$ is big over $T$, then there exists a klt pair $(X,\Delta^{\prime})$, such that $K_{X}+\Delta^{\prime}\sim_{\mathbb{R},T}\mu(K_{X}+\Delta)$ for some $\mu>0$ and $\Delta^{\prime}$ is big.
  2. (2)
    If $\Delta$ is big over $T$, then there exists a klt pair $(X,\Delta_{0})$ such that $K_{X}+\Delta\sim_{\mathbb{R},T}K_{X}+\Delta_{0}$ and $\Delta_{0}\geq A\geq 0$ for some $\mathbb{R}$-divisor $A$ that is ample over $T$.
Proof.
(1) We have $K_{X}+\Delta\sim_{\mathbb{R},T}D$ for some big $\mathbb{R}$-divisor $D\geq 0$. Then $(X,\Delta^{\prime}:=\Delta+\varepsilon D)$ is klt for some $\varepsilon>0$, and $K_{X}+\Delta+\varepsilon D\sim_{\mathbb{R},T}(1+\varepsilon)(K_{X}+\Delta)$. We may let $\mu=1+\varepsilon$.
(2) Write $\Delta\sim_{\mathbb{R},T}H+E$ for some ample $\mathbb{R}$-divisor $H\geq 0$ and $E\geq 0$. We may take $\Delta_{0}:=(1-t)\Delta+t(H+E)$ and $A:=tH$ for some $0<t\ll 1$. ∎
The following corollary of the negativity lemma is needed in the proof of special termination in Section 5.
Lemma 2.19.
Consider a commutative diagram
[…]
where $f,f^{+}$ are birational fibrations. Let $S$ be a prime divisor on $X$ such that $\varphi$ is an isomorphism at the generic point of $S$, and let $S^{+}$ be the strict transform of $S$ on $X^{+}$. Let $\Delta\geq 0$ be an $\mathbb{R}$-divisor on $X$ such that $(X,S+\Delta)$ is dlt, and let $\Delta^{+}\geq 0$ be an $\mathbb{R}$-divisor on $X^{+}$. Assume that
  1. (1)
    $f_{*}\Delta=f^{+}_{*}\Delta^{+}$,
  2. (2)
    $-(K_{X}+S+\Delta)$ is $f$-ample,
  3. (3)
    $K_{X^{+}}+S^{+}+\Delta^{+}$ is $f^{+}$-ample, and
  4. (4)
    the induced birational map $(S,\mathrm{Diff}_{S}(\Delta))\dashrightarrow(S^{+},\mathrm{Diff}_{S^{+}}(\Delta^{+}))$ extends to an isomorphism of pairs.
Then $\varphi$ extends to an isomorphism around $S$.
Proof.
Let $g\colon Y\rightarrow X$ and $h\colon Y\rightarrow X^{+}$ be a common resolution such that $h=\varphi\circ g$. Let $\psi\colon Y\to Z$ be the induced morphism, i.e. $\psi:=f\circ g=f^{+}\circ h$. Write
$g^{*}(K_{X}+S+\Delta)=h^{*}(K_{X^{+}}+S^{+}+\Delta^{+})+E$
for some $\mathbb{R}$-divisor $E$. Let $S_{Y}:=g^{-1}_{*}S$, then we also have $S_{Y}=h^{-1}_{*}S^{+}$, hence $S_{Y}\not\subseteq\mathrm{Supp}(E)$. By adjunction, we get
$g|_{S_{Y}}^{*}(K_{S}+\mathrm{Diff}_{S}(\Delta))=h|_{S_{Y}}^{*}(K_{S^{+}}+\mathrm{Diff}_{S^{+}}(\Delta^{+}))+E|_{S_{Y}}.$
By assumption (4), we also have
$g|_{S_{Y}}^{*}(K_{S}+\mathrm{Diff}_{S}(\Delta))=h|_{S_{Y}}^{*}(K_{S^{+}}+\mathrm{Diff}_{S^{+}}(\Delta^{+})).$
Combined with the previous equality we see that $E|_{S_{Y}}=0$. By assumption (1), the $\mathbb{R}$-divisor $E$ is exceptional over $Z$. By assumptions (2) and (3), $-E$ is nef over $Z$, hence by [KM98, Lemma 3.39], we have $E\geq 0$, and for any closed point $z\in Z$, either $\psi^{-1}(z)\cap\mathrm{Supp}(E)=\emptyset$ or $\psi^{-1}(z)\subseteq\mathrm{Supp}(E)$. Since $E|_{S_{Y}}=0$, this implies that
$\psi(\mathrm{Supp}(E))\cap\psi(S_{Y})=\emptyset.$
Therefore, for any prime divisor $D$ over $Z$ with $\mathrm{Center}_{Z}(D)\cap\psi(S_{Y})\neq\emptyset$ we have
(2.6) $a(D,X,S+\Delta)=a(D,X^{+},S^{+}+\Delta^{+}).$
If $f$ is not an isomorphism around $S$, then there exists a prime divisor $D$ over $X$ such that $\mathrm{Center}_{X}(D)\cap S\neq\emptyset$ and $\mathrm{Center}_{X}(D)\subseteq\mathrm{Ex}(f)$. In particular,
$\mathrm{Center}_{Z}(D)\cap\psi(S_{Y})=f(\mathrm{Center}_{X}(D))\cap f(S)\neq\emptyset.$
The negativity lemma [KM98, Lemma 3.38] and assumptions (2) and (3) then gives
$a(D,X,S+\Delta)<a(D,X^{+},S^{+}+\Delta^{+}),$
which contradicts (2.6). We get a similar contradiction if $f^{+}$ is not an isomorphism around $S^{+}$. Thus $f$ and $f^{+}$ are isomorphisms around $S$ and $S^{+}$ respectively, and hence $\varphi$ is an isomorphism near $S$. ∎
We also need the termination of $4$-fold terminal MMP.
Lemma 2.20.
Let $(X,\Delta)$ be a $\mathbb{Q}$-factorial terminal pair of dimension $\leq 4$. Then every $(K_{X}+\Delta)$-MMP
[…]
where $(X_{i},\Delta_{i})$ is terminal for all $i$, must terminate after finitely many steps.
Proof.
The same proof of [Fuj04, Fuj05] in the projective setting for $\mathbb{Q}$-divisors works here verbatim. ∎
2.7. $\mathbb{T}$-singularities
In this subsection, we collect some results concerning singularities with a torus action. Such singularities naturally appear in the local K-stability theory of klt singularities. We only need the results in this subsection when we prove the boundedness of the fibers of extremal contractions and flips in Section 4.
Let $X=\mathrm{Spec}(R)$ be an affine normal variety and let $\rho\colon\mathbb{G}_{m}\to\mathrm{Aut}(X)$ be a one-parameter subgroup. Then it induces a weight decomposition $R=\oplus_{m\in\mathbb{Z}}R_{m}$ and a $\mathbb{G}_{m}$-equivariant valuation $\mathrm{wt}_{\rho}$ on $X$ such that $\mathrm{wt}_{\rho}(f)=m$ when $0\neq f\in R_{m}$.
In general, let $\mathbb{T}\subseteq\mathrm{Aut}(X)$ be an algebraic torus (i.e. $\mathbb{T}\cong\mathbb{G}_{m}^{r}$ for some $r>0$). Let $N:=\mathrm{Hom}(\mathbb{G}_{m},\mathbb{T})$ be the co-weight lattice, and $M:=\mathrm{Hom}(\mathbb{T},\mathbb{G}_{m})$ the weight lattice. Then we have a weight decomposition
$R=\oplus_{\alpha\in M}R_{\alpha}.$
For any $f\in R$, we denote by $f_{\alpha}$ the component of weight $\alpha\in M$ in this weight decomposition. For any $0\neq\xi\in N_{\mathbb{R}}$, we can define a valuation $\mathrm{wt}_{\xi}$ on $X$ by
$\mathrm{wt}_{\xi}(f):=\min\{\langle\xi,\alpha\rangle\mid\alpha\in M,f_{\alpha}\neq 0\}.$
A one-parameter subgroup $\rho\colon\mathbb{G}_{m}\to\mathbb{T}$ can be viewed as a primitive element of $N$. The corresponding valuation $\mathrm{wt}_{\rho}$ is exactly the one defined in the previous paragraph.
We say that the $\mathbb{T}$-action on $X$ is good if there is a unique closed point $x\in X$ that is in the orbit closure of any $\mathbb{T}$-orbit. This is equivalent to the condition that $R_{0}=\mathbb{C}$. The weight monoid $S$ consists of all $\alpha\in M$ such that $R_{\alpha}\neq 0$, and the weight cone $\sigma\subseteq M_{\mathbb{R}}$ is defined as the (convex, closed, polyhedral) cone generated by $S$. Note that the valuation $\mathrm{wt}_{\xi}$ has a center in $X$ if and only if $\xi$ is contained in the dual $\sigma^{\vee}$ of the weight cone.
Lemma 2.21.
Let $(X,\Delta)$ be a log pair and $\mathbb{T}\subseteq\mathrm{Aut}(X,\Delta)$ a torus. Assume that $X$ is affine. Then there exists some $\alpha\in M_{\mathbb{R}}$ such that $A_{X,\Delta}(\mathrm{wt}_{\xi})=\langle\xi,\alpha\rangle$ for all $\xi\in\sigma^{\vee}$, where $\sigma\subseteq M_{\mathbb{R}}$ is the weight cone.
Proof.
(cf. [LX-stability-higher-rank]*Theorem 2.15(3).) By [AH-T-variety, Theorems 3.1 and 3.4], there exist a proper birational $\mathbb{T}$-equivariant morphism $\varphi\colon\widetilde{X}\to X$ and a morphism $\pi\colon\widetilde{X}\to Y=\widetilde{X}/\!\!/\mathbb{T}$ whose generic fiber $F$ is an affine toric variety. In particular, $X$ is $\mathbb{T}$-equivariantly birational to $\mathbb{T}\times Y$ (where $\mathbb{T}$ acts trivially on $Y$). Let $K_{\widetilde{X}}+\widetilde{\Delta}=\varphi^{*}(K_{X}+\Delta)$ be the crepant pullback. Then $\widetilde{\Delta}|_{F}$ is a torus invariant $\mathbb{R}$-divisor. Since $A_{X,\Delta}(\mathrm{wt}_{\xi})=A_{\widetilde{X},\widetilde{\Delta}}(\mathrm{wt}_{\xi})$ and the center of $\mathrm{wt}_{\xi}$ dominates $Y$, it suffices to compute the log discrepancy along the generic fiber $F$ of $\pi$. Thus we reduce to the toric case, where the desired statement is well known (see, for example, [Amb-mld]*Section 4). ∎
Given a morphism $\mathcal{X}\to\mathcal{B}$, we call a torus $\mathbb{T}$-action on $\mathcal{X}$ that commutes with the trivial $\mathbb{T}$-action on $\mathcal{B}$ a fiberwise $\mathbb{T}$-action. We next observe that the weight cone and the log discrepancy function are locally constant for any $\mathbb{R}$-Gorenstein family of klt singularities with a fiberwise torus action.
Definition 2.22.
We call $f\colon(\mathcal{X},\mathcal{D})\to\mathcal{B}$ an $\mathbb{R}$-Gorenstein family of pairs if
  1. (1)
    $f\colon\mathcal{X}\to\mathcal{B}$ is an affine, flat morphism,
  2. (2)
    for any closed point $b\in\mathcal{B}$, the fiber $\mathcal{X}_{b}$ is connected and normal,
  3. (3)
    $\mathcal{D}\geq 0$ is an $\mathbb{R}$-divisor on $\mathcal{X}$ whose support does not contain any fiber $\mathcal{X}_{b}$ of $f$,
  4. (4)
    $\mathcal{X}$ and $\mathcal{B}$ are normal, and $K_{\mathcal{X}/\mathcal{B}}+\mathcal{D}$ is $\mathbb{R}$-Cartier.
[…]
Lemma 2.23.
Let $\mathcal{X}\to\mathcal{B}$ be an affine, surjective morphism of varieties with a fiberwise torus $\mathbb{T}$-action. Then the weight monoid of the $\mathbb{T}$-action on the fiber $\mathcal{X}_{b}$ is locally constant as $b\in\mathcal{B}$ varies.
Proof.
We may assume that both $\mathcal{B}$ and $\mathcal{X}$ are affine, and $\mathcal{B}$ is connected. Let $\mathcal{X}=\mathrm{Spec}(R)$. It suffices to show that for any $\alpha\in M$ with $R_{\alpha}\neq 0$, we have $R_{\alpha}|_{\mathcal{X}_{b}}\neq 0$ for all $b\in\mathcal{B}$. Since we have a fiberwise $\mathbb{T}$-action, the morphism $\mathcal{X}\to\mathcal{B}$ factors through $X_{0}:=\mathcal{X}/\!\!/\mathbb{T}=\mathrm{Spec}(R_{0})$, thus we may replace $\mathcal{B}$ by $X_{0}$ to show that the restriction $R_{\alpha}|_{\mathcal{X}_{b}}$ is nonzero. Since $R$ is generated by finitely many homogeneous elements, it is not hard to see that each $R_{\alpha}$ is finitely generated over $R_{0}$. Alternatively, this also follows directly from [AH-T-variety, Theorems 3.1 and 3.4]: there exist a projective morphism $Y\to\mathrm{Spec}(R_{0})$ and a semiample line bundle $L_{\alpha}$ (one for each $\alpha$ in the weight monoid of $R$) on $Y$ such that $R_{\alpha}=H^{0}(Y,L_{\alpha})$. If $R_{\alpha}|_{X_{b}}=0$, then Nakayama’s lemma implies that $R_{\alpha}=0$ in a neighborhood of $b\in X_{0}$, thus $R_{\alpha}$ is a torsion $R_{0}$-module. Since $R$ is an integral domain, we deduce that $R_{\alpha}=0$. This proves the lemma. ∎
Lemma 2.24.
Let $\mathcal{B}\subseteq(\mathcal{X},\mathcal{D})\to\mathcal{B}$ be an $\mathbb{R}$-Gorenstein family of singularities with a fiberwise torus $\mathbb{T}$-action and let $\sigma$ be the weight cone of the this $\mathbb{T}$-action. Then for any $\xi\in\sigma^{\vee}$, the log discrepancy function $b\mapsto A_{\mathcal{X}_{b},\mathcal{D}_{b}}(\mathrm{wt}_{\xi})$ is locally constant on $\mathcal{B}$.
Note that the morphism $\mathcal{X}\to\mathcal{B}$ is surjective since it has a section. Thus by Lemma 2.23, the weight monoid $S$ and the weight cone $\sigma$ of the $\mathbb{T}$-action on the fiber $b\in(\mathcal{X}_{b},\mathcal{D}_{b})$ is locally constant in the family.
Proof.
We may assume that $\mathcal{B}$ is affine and connected. Since the log discrepancy function is linear on the weight cone by Lemma 2.21, it suffices to prove the statement when $\xi$ is a primitive element in the dual $S^{\vee}$ of the weight monoid. In this case $\xi$ corresponds to a one-parameter subgroup of $\mathbb{T}$. After replacing $\mathbb{T}$ by this one-parameter subgroup, we reduce to the case when $\mathbb{T}=\mathbb{G}_{m}$.
Let $\mathcal{X}=\mathrm{Spec}(\mathcal{R})$. Then we have a weight decomposition $\mathcal{R}=\oplus_{m\in\mathbb{N}}\mathcal{R}_{m}$. Let $\mathfrak{a}_{m}:=\sum_{k\geq m}\mathcal{R}_{k}\subseteq\mathcal{R}$ and $\mathcal{Y}:=\mathrm{Proj}\left(\bigoplus_{m\in\mathbb{N}}\mathfrak{a}_{m}\right)$. Then $\mathfrak{a}_{m}|_{\mathcal{X}_{b}}$ is the valuation ideal of $\mathrm{wt}_{\xi}$ (as a valuation on $\mathcal{X}_{b}$), and we have a proper birational morphism $\pi\colon\mathcal{Y}\to\mathcal{X}$ with a unique exceptional divisor $\mathcal{E}\cong\mathrm{Proj}(\mathcal{R})$. Note that $\mathcal{E}$ is $\mathbb{Q}$-Cartier, we have $\mathcal{E}_{b}\cong\mathrm{Proj}(\mathcal{R}_{b})$ (where $\mathcal{R}_{b}:=\mathcal{R}\otimes k(b)$) is a prime divisor on $\mathcal{Y}_{b}$ for all $b\in\mathcal{B}$ (since $\mathcal{X}_{b}\cong\mathrm{Spec}(\mathcal{R}_{b})$ is integral), and $\mathrm{wt}_{\xi}=\mathrm{ord}_{\mathcal{E}_{b}}$ as valuations on $\mathcal{X}_{b}$. Let $\mathcal{D}_{\mathcal{Y}}$ be the strict transform of $\mathcal{D}$ on $\mathcal{Y}$. Then
$K_{\mathcal{Y}}+D_{\mathcal{Y}}+\mathcal{E}\sim_{\mathbb{R}}\pi^{*}(K_{\mathcal{X}}+\mathcal{D})+a\mathcal{E}$
where $a=A_{\mathcal{X},\mathcal{D}}(\mathcal{E})$. As the family $(\mathcal{X},\mathcal{D})\to\mathcal{B}$ is $\mathbb{R}$-Gorenstein, by restricting to the fibers we get
$K_{\mathcal{Y}_{b}}+D_{\mathcal{Y},b}+\mathcal{E}_{b}\sim_{\mathbb{R}}\pi^{*}(K_{\mathcal{X}_{b}}+\mathcal{D}_{b})+a\mathcal{E}_{b}$
for all $b\in\mathcal{B}$; in other words, $A_{\mathcal{X}_{b},\mathcal{D}_{b}}(\mathrm{wt}_{\xi})=A_{\mathcal{X}_{b},\mathcal{D}_{b}}(\mathcal{E}_{b})=a$ is independent of $b\in\mathcal{B}$. ∎
We also need the special boundedness of klt singularities when the local volume is bounded away from zero. To state the precise result, we need some definitions.
Definition 2.25.
A test configuration of a singularity $x\in(X,\Delta)$ consists of a test configuration $(\mathcal{X},\mathcal{D})\to\mathbb{A}^{1}$ of the pair $(X,\Delta)$ and a $\mathbb{G}_{m}$-equivariant section $\sigma\colon\mathbb{A}^{1}\to\mathcal{X}$ such that $\sigma(t)=x$ for $t\neq 0$ under the isomorphism $(\mathcal{X}_{t},\mathcal{D}_{t})\cong(X,\Delta)$. We still denote the test configuration by $(\mathcal{X},\mathcal{D})\to\mathbb{A}^{1}$ as the section $\sigma$ is uniquely determined by this data. When $(X,\Delta)$ is klt, the test configuration $(\mathcal{X},\mathcal{D})\to\mathbb{A}^{1}$ is called special if $(\mathcal{X},\mathcal{X}_{0}+\mathcal{D})$ is plt. In this case $(\mathcal{X}_{0},\mathcal{D}_{0})$ is also klt (by adjunction) and we call $\sigma(0)\in(\mathcal{X}_{0},\mathcal{D}_{0})$ a special degeneration of $x\in(X,\Delta)$.
Example 2.26.
Let $x\in(X=\mathrm{Spec}(R),\Delta)$ be a singularity and let $E$ be a prime divisor over $x\in X$. Let
$\mathfrak{a}_{m}:=\{f\in R\,|\,\mathrm{ord}_{E}(f)\geq m\}.$
Assume that the graded algebra $\mathrm{gr}_{E}R:=\bigoplus_{m\in\mathbb{N}}\mathfrak{a}_{m}/\mathfrak{a}_{m+1}$ is finitely generated. Then so is $\bigoplus_{m\in\mathbb{N}}\mathfrak{a}_{m}$. By the Rees construction, we get a flat family $\pi\colon\mathcal{X}=\mathrm{Spec}(\mathcal{R})\to\mathbb{A}^{1}$ where
$\mathcal{R}=\bigoplus_{m\in\mathbb{Z}}t^{-m}\mathfrak{a}_{m}$
is the (extended) Rees algebra and $t$ is the affine coordinate on $\mathbb{A}^{1}$. The $\mathbb{Z}$-grading on $\mathcal{R}$ induces a $\mathbb{G}_{m}$-action on the family and this gives a test configuration of the singularity $x\in X$. Note that $\mathcal{X}_{0}\cong\mathrm{Spec}(\mathrm{gr}_{E}R)$, and $\mathrm{gr}_{E}R$ is an integral domain, hence $\mathcal{X}_{0}$ is integral.
Let $\mathcal{D}$ be the closure of $\Delta\times(\mathbb{A}^{1}\setminus\{0\})$ in $\mathcal{X}$ under the isomorphism $\mathcal{X}\setminus\mathcal{X}_{0}\cong X\times(\mathbb{A}^{1}\setminus\{0\})$. Then $(\mathcal{X},\mathcal{D})\to\mathbb{A}^{1}$ is a test configuration of $x\in(X,\Delta)$. Let us prove that if $K_{X}+\Delta$ is $\mathbb{R}$-Cartier, then so is $K_{\mathcal{X}}+\mathcal{D}$. Indeed, by shrinking $X$ around $x$ (which preserves $\mathrm{gr}_{E}R$ and thus does not change the central fiber $\mathcal{X}_{0}$) we may assume that $K_{X}+\Delta\sim_{\mathbb{R}}0$, thus $K_{\mathcal{X}\setminus\mathcal{X}_{0}}+\mathcal{D}|_{\mathcal{X}\setminus\mathcal{X}_{0}}\sim_{\mathbb{R}}0$. As $\mathcal{X}_{0}$ is irreducible, we deduce that $K_{\mathcal{X}}+\mathcal{D}\sim_{\mathbb{R}}c\mathcal{X}_{0}$ for some $c\in\mathbb{R}$. But $\mathcal{X}_{0}\sim_{\mathbb{R}}0$, hence $K_{\mathcal{X}}+\mathcal{D}\sim_{\mathbb{R}}0$. In particular, $K_{\mathcal{X}}+\mathcal{D}$ is $\mathbb{R}$-Cartier. It follows (by inversion of adjunction) that if $x\in(X,\Delta)$ is klt, then $(\mathcal{X},\mathcal{D})\to\mathbb{A}^{1}$ is a special test configuration if and only if $(\mathcal{X}_{0},\mathcal{D}_{0})$ is klt.
The following special boundedness statement is one of the main results in the local K-stability theory of klt singularities.
Theorem 2.27.
Let $n\in\mathbb{N}$, $\varepsilon>0$ and let $I\subseteq[0,1]$ be a finite set. Then there exists an $\mathbb{R}$-Gorenstein family $\mathcal{B}\subseteq(\mathcal{X},\mathcal{D})\to\mathcal{B}$ of klt singularities with the action of an algebraic torus $\mathbb{T}$ over $\mathcal{B}$, such that:
For any klt singularity $x\in(X,\Delta)$ of dimension $n$, local volume at least $\varepsilon$ and coefficients in $I$, and any torus $\mathbb{T}_{0}\subseteq\mathrm{Aut}(X,\Delta)$, there exists a $\mathbb{T}_{0}$-equivariant special degeneration of $x\in(X,\Delta)$ to $b\in(\mathcal{X}_{b},\mathcal{D}_{b})$ for some $b\in\mathcal{B}$ that identifies $\mathbb{T}_{0}$ with a subtorus of $\mathbb{T}_{b}$.
Proof.
This essentially follows from the stable degeneration theory of klt singularities together with the main results of [XZ-local-bdd]. We refer to [XZ-local-bdd, Section 2.2-2.3] for the relevant definitions that appear in this proof. Enlarging the finite set $I$ if necessarily, we may assume that for any $a,b\in I$ with $a+b\leq 1$ we have $a+b\in I$. By the Stable Degeneration Theorem [Blu-minimizer-exist, LX-stability-higher-rank, Xu-quasi-monomial, XZ-minimizer-unique, XZ-SDC] (see also [Z-SDC-survey] for a survey and [Z-mld^K-2] for the extension to real coefficients), for every klt singularity $x\in(X=\mathrm{Spec}(R),\Delta)$, up to rescaling there is a unique and hence $\mathbb{T}_{0}$-invariant quasi-monomial valuation $v$ minimizing the normalized volume function $\widehat{\rm vol}_{X,\Delta}$. It induces a $\mathbb{T}_{0}$-equivariant degeneration of $x\in(X,\Delta)$ to a K-semistable log Fano cone singularity $x_{0}\in(X_{0}=\mathrm{Spec}(\mathrm{gr}_{v}R),\Delta_{0};\xi_{v})$ with
$\widehat{\rm vol}(x,X,\Delta)=\widehat{\rm vol}(x_{0},X_{0},\Delta_{0}).$
Note that this implies that the torus $\langle\xi_{v}\rangle$ generated by the Reeb vector field $\xi_{v}$ commutes with $\mathbb{T}_{0}$. Here $\mathrm{gr}_{v}R$ is defined by $\bigoplus_{\lambda\in\mathbb{R}}\mathfrak{a}_{\lambda}/\mathfrak{a}_{>\lambda}$ where $\mathfrak{a}_{\lambda}=\{f\in R\,|\,v(f)\geq\lambda\}$ and $\mathfrak{a}_{>\lambda}=\{f\in R\,|\,v(f)>\lambda\}$.
We claim that $x_{0}\in(X_{0},\Delta_{0})$ can be realized as a $\mathbb{T}_{0}$-equivariant special degeneration of $x\in(X,\Delta)$. Indeed, by [LX-stability-higher-rank, Lemma 2.10], there exists some divisorial valuation $v_{0}$ over $x\in X$ (obtained from a small perturbation of the weights defining the quasi-monomial valuations $v$) such that $\mathrm{gr}_{v}R\cong\mathrm{gr}_{v_{0}}R$. Moreover, since $v$ is $\mathbb{T}_{0}$-invariant, the construction in the proof of loc. cit. can be made $\mathbb{T}_{0}$-equivariant (for example, we can choose $\mathbb{T}_{0}$-invariant generators of $\mathrm{gr}_{v}R$ as well as their lifts to $R$), thus we may assume that the perturbed valuation $v_{0}$ is also $\mathbb{T}_{0}$-invariant. Write $v_{0}=\lambda\cdot\mathrm{ord}_{E}$ where $E$ is a prime divisor over $x\in X$. Since $\mathrm{gr}_{E}R\cong\mathrm{gr}_{v_{0}}R\cong\mathrm{gr}_{v}R$, the divisor $E$ induces a $\mathbb{T}_{0}$-equivariant special degeneration of $x\in(X,\Delta)$ to $x_{0}\in(X_{0},\Delta_{0})$ by the construction in Example 2.26. This proves the claim.
By construction, $\widehat{\rm vol}(x_{0},X_{0},\Delta_{0})\geq\varepsilon$ and the coefficients of $\Delta_{0}$ belong to $I$. Since $x_{0}\in(X_{0},\Delta_{0};\xi_{v})$ is K-semistable, we also have $\Theta(X_{0},\Delta_{0};\xi_{v})=1$ where $\Theta(X_{0},\Delta_{0};\xi_{v})$ denotes the volume ratio (see [XZ-local-bdd, Definition 2.7]). By perturbing $\xi_{v}$ in the Reeb cone, we can choose some Reeb vector field $\xi_{0}$ such that the torus $\langle\xi_{0}\rangle$ it generates contains both $\langle\xi_{v}\rangle$ and $\mathbb{T}_{0}$ and we have $\Theta(X_{0},\Delta_{0};\xi_{0})\geq\frac{1}{2}$ (this uses the continuity of the normalized volume function on the Reeb cone, see [Z-mld^K-2, Lemma 2.11]). By [XZ-local-bdd, Theorem 1.4], such polarized log Fano cone singularities $x_{0}\in(X_{0},\Delta_{0};\xi_{0})$ are bounded. By definition (see [XZ-local-bdd, Definition 2.16]), this means that there exists an $\mathbb{R}$-Gorenstein family $\mathcal{B}\subseteq(\mathcal{X},\mathcal{D})\to\mathcal{B}$ of klt singularities with the action of an algebraic torus $\mathbb{T}$ over $\mathcal{B}$, such that $x_{0}\in(X_{0},\Delta_{0};\xi_{0})$ is isomorphic to some fiber $b\in(\mathcal{X}_{b},\mathcal{D}_{b})$ and $\xi_{0}$ is contained in the Reeb cone of $\mathbb{T}_{b}$. Thus $\mathbb{T}_{b}$ contains the torus generated by $\xi_{0}$ and in particular it contains $\mathbb{T}_{0}$. This proves the theorem. ∎
2.8. Log Fano and log Calabi-Yau fibrations
A key step in bounding the fibers of extremal contractions and flips (Section 4) is the special boundedness of certain log Fano and log Calabi-Yau fibrations. In this subsection, we discuss some properties of such fibrations and define their special boundedness.
Definition 2.28.
Let $(X,\Delta)$ be a pair and $X\to Z$ a projective morphism.
  1. (1)
    We say that $(X,\Delta)$ is log Fano over $Z$ if $(X,\Delta)$ is klt and $-(K_{X}+\Delta)$ is ample over $Z$.
  2. (2)
    We say that $(X,\Delta)$ is log Calabi-Yau over $Z$ if $(X,\Delta)$ is log canonical and $K_{X}+\Delta\sim_{\mathbb{R},Z}0$.
  3. (3)
    We say that $(X,\Delta)$ is of Fano type over $Z$ if there exists some $\mathbb{R}$-divisor $\Delta^{\prime}\geq\Delta$ such that $(X,\Delta^{\prime})$ is log Fano over $Z$. Equivalently, $-(K_{X}+\Delta)$ is big over $Z$ and there exists some $\mathbb{R}$-divisor $0\leq D\sim_{\mathbb{R},Z}-(K_{X}+\Delta)$ such that $(X,\Delta+D)$ is klt.
A fibration $(X,\Delta)\to Z$ (or fibration germ $(X,\Delta)\to Z\ni z$) is called a log Fano (resp. log Calabi-Yau, resp. Fano type) fibration (or fibration germ) if $(X,\Delta)$ is log Fano (resp. log Calabi-Yau, resp. of Fano type) over $Z$.
Lemma 2.29.
Let $f\colon(X^{\prime},\Delta^{\prime})\dashrightarrow(X,\Delta)$ be a birational map between pairs that are projective over $Z$. Assume that $K_{X}+\Delta$ is $\mathbb{R}$-Cartier, $f^{*}(K_{X}+\Delta)\geq K_{X^{\prime}}+\Delta^{\prime}$, and $(X,\Delta)$ is of Fano type over $Z$. Then $(X^{\prime},\Delta^{\prime})$ is also of Fano type over $Z$.
Proof.
Since $(X,\Delta)$ is of Fano type over $Z$, we know that $-(K_{X}+\Delta)$ is big over $Z$ and there exists some $\mathbb{R}$-divisor $0\leq D\sim_{\mathbb{R},Z}-(K_{X}+\Delta)$ such that $(X,\Delta+D)$ is klt. As $f^{*}(K_{X}+\Delta)\geq K_{X^{\prime}}+\Delta^{\prime}$, by taking crepant pullback we see that both conditions are also satisfied by $(X^{\prime},\Delta^{\prime})$, hence $(X^{\prime},\Delta^{\prime})$ is of Fano type over $Z$. ∎
Definition 2.30.
A test configuration of a fibration $f\colon(X,\Delta)\to Z$ consists of a test configuration $(\mathcal{X},\mathcal{D})\to\mathbb{A}^{1}$ of the pair $(X,\Delta)$ and a $\mathbb{G}_{m}$-equivariant fibration $\widetilde{f}\colon(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ over $\mathbb{A}^{1}$, where $\mathcal{Z}$ is a test configuration of $Z$, such that $\widetilde{f}|_{\mathcal{X}_{t}}=f$ for all $0\neq t\in\mathbb{A}^{1}$. For simplicity, we denote the test configuration by $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$. Test configuration of a fibration germ $(X,\Delta)\to Z\ni z$ is defined as a test configuration $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ of the fibration $(X,\Delta)\to Z$ such that $\mathcal{Z}$ is also a test configuration of the singularity $z\in Z$.
A test configuration $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ of a log Fano fibration $(X,\Delta)\to Z$ is called special if $(\mathcal{X},\mathcal{X}_{0}+\mathcal{D})$ is plt, $\mathcal{Z}_{0}$ is normal and $-(K_{\mathcal{X}}+\mathcal{D})$ is ample over $\mathcal{Z}$. Note that under these conditions $\mathcal{X}_{0}$ is normal by [Kol13]*Theorem 4.16. In this case $(\mathcal{X}_{0},\mathcal{D}_{0})\to\mathcal{Z}_{0}$ is also a log Fano fibration and we call it a special degeneration of $(X,\Delta)\to Z$. Special test configurations and special degenerations of a log Fano fibration germ are defined analogously.
Definition 2.31.
We say that $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ is an $\mathbb{R}$-Gorenstein family of log Fano fibrations over a finite type base $\mathcal{B}$ if:
  1. (1)
    Both $\mathcal{X}$ and $\mathcal{Z}$ are flat over $\mathcal{B}$, and $\mathcal{X}\to\mathcal{Z}$ is a fibration,
  2. (2)
    for any $b\in\mathcal{B}$, the fibers $\mathcal{X}_{b}$ and $\mathcal{Z}_{b}$ are connected and normal,
  3. (3)
    $\mathcal{D}\geq 0$ is an $\mathbb{R}$-divisor on $\mathcal{X}$ whose support does not contain any fiber $\mathcal{X}_{b}$ of $\mathcal{X}\to\mathcal{B}$,
  4. (4)
    $\mathcal{X}$ and $\mathcal{B}$ are normal, $-(K_{\mathcal{X}/\mathcal{B}}+\mathcal{D})$ is $\mathbb{R}$-Cartier and ample over $\mathcal{Z}$, and
  5. (5)
    the pair $(\mathcal{X}_{b},\mathcal{D}_{b})$ is klt for any closed point $b\in\mathcal{B}$.
In particular, $(\mathcal{X}_{b},\mathcal{D}_{b})\to\mathcal{Z}_{b}$ is a log Fano fibration for every $b\in\mathcal{B}$.
Definition 2.32.
We say that a set $\mathcal{S}$ of Fano type fibrations is bounded if there exist an $\mathbb{R}$-Gorenstein family of log Fano fibration $(\mathcal{X},\widetilde{\mathcal{D}})\to\mathcal{Z}$ over a finite type base $\mathcal{B}$ and an $\mathbb{R}$-divisor $0\leq\mathcal{D}\leq\widetilde{\mathcal{D}}$, such that any Fano type fibration $(X,\Delta)\to Z$ in $\mathcal{S}$ is isomorphic to $(\mathcal{X}_{b},\mathcal{D}_{b})\to\mathcal{Z}_{b}$ for some $b\in\mathcal{B}$.
For log Fano fibrations, it is tempting to define their boundedness by requiring $\mathcal{D}=\widetilde{\mathcal{D}}$ in the above definition. The next lemma shows that these two definitions are equivalent.
Lemma 2.33.
Let $\mathcal{S}$ be a set of log Fano fibrations. Assume that $\mathcal{S}$ is bounded as a set of Fano type fibrations. Then there exists a $\mathbb{R}$-Gorenstein family $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ of log Fano fibrations over a finite type base $\mathcal{B}$ such that any log Fano fibration $(X,\Delta)\to Z$ in $\mathcal{S}$ is isomorphic to $(\mathcal{X}_{b},\mathcal{D}_{b})\to\mathcal{Z}_{b}$ for some $b\in\mathcal{B}$.
Proof.
Let $(\mathcal{X},\widetilde{\mathcal{D}})\to\mathcal{Z}$ and $0\leq\mathcal{D}\leq\widetilde{\mathcal{D}}$ be as in Definition 2.32. Then there exists some constant $\varepsilon>0$ such that the all fibers of $(\mathcal{X},\widetilde{\mathcal{D}})\to\mathcal{B}$ are $\varepsilon$-klt (i.e. their total minimal log discrepancies are larger than $\varepsilon$). It follows that for any $(X,\Delta)\to Z$ in $\mathcal{S}$ we also have $(X,\Delta)$ is $\varepsilon$-klt. By [HX-CY-bdd, Proposition 2.4] and Noetherian induction, we deduce that possibly after stratifying $\mathcal{B}$, we may assume that $K_{\mathcal{X}/\mathcal{B}}+\mathcal{D}$ is $\mathbb{R}$-Cartier. Thus $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ is an $\mathbb{R}$-Gorenstein family of log Fano fibrations and we are done. ∎
Definition 2.34.
Let $(X,\Delta)\to Z$ be a Fano type fibration (or fibration germ). A test configuration $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ of $(X,\Delta)\to Z$ is said to be of Fano type if there exists an $\mathbb{R}$-divisor $\widetilde{\mathcal{D}}\geq\mathcal{D}\geq 0$ on $\mathcal{X}$ such that $(\mathcal{X},\widetilde{\mathcal{D}})\to\mathcal{Z}$ is an $\mathbb{R}$-Gorenstein family of log Fano fibrations over $\mathbb{A}^{1}$. In particular, $(\mathcal{X}_{0},\mathcal{D}_{0})\to\mathcal{Z}_{0}$ is a Fano type fibration. In this case we also say that $(X,\Delta)\to Z$ degenerates to $(\mathcal{X}_{0},\mathcal{D}_{0})\to\mathcal{Z}_{0}$ by a Fano type test configuration.
Lemma 2.35.
Let $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ be a Fano type test configuration (of some fibration), let $(\mathcal{X}^{\prime},\mathcal{D}^{\prime})\to\mathcal{Z}$ be another test configuration (of a possibly different fibration), and let $f\colon\mathcal{X}^{\prime}\dashrightarrow\mathcal{X}$ be a $\mathbb{G}_{m}$-equivariant birational map over $\mathcal{Z}$. Assume that $\mathcal{X}^{\prime}_{0}$ is irreducible, $f$ is an isomorphism at the generic point of $\mathcal{X}_{0}$, and $f^{*}(K_{\mathcal{X}}+{\mathcal{D}})\geq K_{\mathcal{X}^{\prime}}+\mathcal{D}^{\prime}$. Then $(\mathcal{X}^{\prime},\mathcal{D}^{\prime})\to\mathcal{Z}$ is also a Fano type test configuration.
Proof.
The condition that $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ is a Fano type test configuration is equivalent to saying that $-(K_{\mathcal{X}}+\mathcal{D})$ is big over $\mathcal{Z}$ and there exists some $\mathbb{R}$-divisor $0\leq\mathcal{G}\sim_{\mathbb{R},\mathcal{Z}}-(K_{\mathcal{X}}+\mathcal{D})$ on $\mathcal{X}$ such that $(\mathcal{X},\mathcal{X}_{t}+\mathcal{D}+\mathcal{G})$ is plt for every $t\in\mathbb{A}^{1}$. These conditions are preserved by the given birational map $f$. ∎
Definition 2.36.
We say that a set $\mathcal{S}$ of log Fano fibrations germs is specially bounded if there exists a bounded set $\mathcal{T}$ of log Fano fibrations such that any log Fano fibration germ $(X,\Delta)\to Z\ni z$ in $\mathcal{S}$ specially degenerates to some $(X_{0},\Delta_{0})\to Z_{0}\ni z_{0}$ in $\mathcal{T}$.
We say that a set $\mathcal{S}$ of Fano type fibrations germs is bounded up to Fano type degenerations if there exists a bounded set $\mathcal{T}$ of Fano type fibrations such that any Fano type fibration germ $(X,\Delta)\to Z\ni z$ in $\mathcal{S}$ degenerates to some $(X_{0},\Delta_{0})\to Z_{0}\ni z_{0}$ in $\mathcal{T}$ by a Fano type test configuration.
Finally, we say that a set $\mathcal{S}$ of fibration germs is log specially bounded if there exists a bounded set $\mathcal{T}$ of Fano type fibrations, such that for any fibration germ $(X,\Delta)\to Z\ni z$ in $\mathcal{S}$, we can find an $\mathbb{R}$-divisor $D\geq 0$ on $X$ with $\mathrm{Supp}(\Delta)\subseteq\mathrm{Supp}(D)$, a Fano type fibration $(X_{0},\Delta_{0})\to Z_{0}\ni z_{0}$ in $\mathcal{T}$, and some $\mathbb{R}$-divisor $D_{0}\geq 0$ on $X_{0}$ with $\mathrm{Supp}(D_{0})\subseteq\mathrm{Supp}(\Delta_{0})$, such that $(X,D)\to Z\ni z$ is a Fano type fibration, and it degenerates to $(X_{0},D_{0})\to Z_{0}\ni z_{0}$ by a Fano type test configuration. Intuitively, the conditions mean that $\mathrm{Supp}(\Delta)$ degenerates into $\mathrm{Supp}(\Delta_{0})$ by some Fano type test configuration.
3. Boundedness of invariants in MMP
The major goal of this section is to prove the following theorem, which is a relative version of Theorem 1.1.
Theorem 3.1.
Let $(X,\Delta)$ be a $\mathbb{Q}$-factorial klt pair and $X\rightarrow T$ a projective morphism. Assume that either $K_{X}+\Delta$ or $\Delta$ is big over $T$. Then there exist a positive integer $r$ and a finite set $S\subseteq\mathbb{R}_{+}$, depending only on $(X,\Delta)$ and $\pi$, such that for any sequence of steps of a $(K_{X}+\Delta)$-MMP $(X,\Delta)\dashrightarrow(Y,\Delta_{Y})$ over $T$,
  1. (1)
    the Cartier index of any Weil divisor on $Y$ is at most $r$, and
  2. (2)
    for any (not necessarily closed) point $y\in Y$, we have $\mathrm{mld}(y,Y,\Delta_{Y})\in S$.
3.1. Local volumes in MMP type contractions
We shall deduce Theorem 3.1 from the following more general result on the boundedness of invariants in MMP type contractions (Definition 2.12).
Theorem 3.2.
Let $(X,\Delta+D)$ be a klt pair where $\Delta,D\geq 0$ and $D$ is big. Then there exist a positive integer $r$ and two real number $\varepsilon,A>0$, depending only on $X$, $\Delta$ and $D$, such that for any MMP type contraction $\varphi\colon(X,\Delta+D)\dashrightarrow(X^{\prime},\Delta^{\prime}+D^{\prime})$ with $\Delta^{\prime}\geq 0$ and $D^{\prime}=\varphi_{*}D$, we have:
  1. (1)
    The Cartier index of any $\mathbb{Q}$-Cartier Weil divisor on $X^{\prime}$ is at most $r$,
  2. (2)
    for any (not necessarily closed) point $x^{\prime}\in X^{\prime}$, we have $\mathrm{mld}(x^{\prime},X^{\prime},\Delta^{\prime}+D^{\prime})\leq A$, and
  3. (3)
    if $K_{X^{\prime}}+\Delta^{\prime}$ is $\mathbb{R}$-Cartier, then $\widehat{\rm vol}(x^{\prime},X^{\prime},\Delta^{\prime})\geq\varepsilon$ for all closed points $x^{\prime}\in X^{\prime}$.
Before we prove this more general result, let us explain how it implies Theorem 3.1.
Proof of Theorem 3.1 assuming Theorem 3.2.
By Lemma 2.18, there exists some $\mathbb{R}$-divisor $\Delta_{0}\geq 0$ on $X$ such that $(X,\Delta_{0})$ is klt, $\Delta_{0}$ is big, and any $(K_{X}+\Delta)$-MMP over $T$ is also a $(K_{X}+\Delta_{0})$-MMP. Part (1) then follows directly from Theorem 3.2(1). By Lemma 2.2, this implies that the log discrepancies $A_{Y,\Delta_{Y}}(E)$ of prime divisors $E$ over $Y$ belong to a discrete set $S_{0}$ that depends only on the coefficients of $\Delta$ and the index bound $r$. Note that $Y$ is $\mathbb{Q}$-factorial. Thus by Theorem 3.2(3), we also know that there exists some positive real number $\varepsilon$ (depending only on $X$ and $\Delta_{0}$) such that $\widehat{\rm vol}(y,Y)\geq\varepsilon$ for any closed point $y\in Y$. Combined with Theorem 2.5, we see that there exists some constant $A>0$ (depending only on $\dim X$ and $\varepsilon$) such that $\mathrm{mld}(y,Y)\leq A$ for all points $y\in Y$, hence $\mathrm{mld}(y,Y,\Delta_{Y})\leq A$ as well. It follows that $\mathrm{mld}(y,Y,\Delta_{Y})$ belongs to the finite set $S:=S_{0}\cap[0,A]$. ∎
In the rest of this subsection, we prove Theorem 3.2. Since local volume controls most other invariants such as index and mld, we will primarily focus on bounding the local volume away from zero. A key ingredient is the following generalization of [XZ-local-bdd]*Lemma 2.13 (cf. [BJ-delta]*Theorem D).
Proposition 3.3.
Let $(X,\Delta)$ be a klt pair of dimension $n$, and let $V_{\bullet}$ be an eventually birational graded linear series (see Section 2.5) of some $\mathbb{R}$-Cartier $\mathbb{R}$-divisor $L$ on $X$. Then for any closed point $x\in X$ we have
$\widehat{\rm vol}(x,X,\Delta)\geq\alpha(X,\Delta;V_{\bullet})^{n}\cdot\mathrm{vol}(V_{\bullet}).$
The proof is in turn based on the following auxiliary lemma that we will apply again in Section 4.
Lemma 3.4.
Let $(X,\Delta)$ be a klt pair of dimension $n$, let $x\in X$ be a closed point, and let $c_{0}>c_{1}$ be positive constants. Then for any sufficiently large integer $m$ and any linear subspace $V\subseteq\mathcal{O}_{X,x}$ of dimension at least $c_{0}\frac{m^{n}}{n!}$, there exists some nonzero element $s\in V$ such that
$m\cdot\mathrm{lct}_{x}(X,\Delta;\{s=0\})\leq\left(\frac{\widehat{\rm vol}(x,X,\Delta)}{c_{1}}\right)^{1/n}.$
Proof.
This is proved by the same argument as in [BJ-delta]*Theorem D or [XZ-local-bdd]*Lemma 2.13. Let $c:=\left(\frac{\widehat{\rm vol}(x,X,\Delta)}{c_{1}}\right)^{-1/n}$ and let $v$ be a valuation over $x\in X$ such that $\widehat{\rm vol}_{X,\Delta}(v)=\widehat{\rm vol}(x,X,\Delta)$. Since the normalized volume is rescaling invariant (i.e. $\widehat{\rm vol}_{X,\Delta}(\lambda v)=\widehat{\rm vol}_{X,\Delta}(v)$ for all $v\in\mathrm{Val}_{X,x}$ and $\lambda>0$), we may rescale the minimizing valuation $v$ and assume that $A_{X,\Delta}(v)=1$. It follows that $\mathrm{vol}(v)=\widehat{\rm vol}(x,X,\Delta)$. Let $\mathfrak{a}_{\lambda}:=\mathfrak{a}_{\lambda}(v)$ ($\lambda\in\mathbb{R}$) be the valuation ideals of $v$. Since $c_{0}>c_{1}>0$, for sufficiently large $m$ we have
$\dim(\mathcal{O}_{X,x}/\mathfrak{a}_{cm})=\widehat{\rm vol}(x,X,\Delta)\cdot\frac{(cm)^{n}}{n!}+o(m^{n})=\frac{c_{1}m^{n}}{n!}+o(m^{n})<\frac{c_{0}m^{n}}{n!}.$
Thus for any linear subspace $V\subseteq\mathcal{O}_{X,x}$ of dimension at least $c_{0}\frac{m^{n}}{n!}$, the natural map $V\to\mathcal{O}_{X,x}/\mathfrak{a}_{cm}$ cannot be injective. It follows that there exists some nonzero element $s\in V\cap\mathfrak{a}_{cm}(v)$; in particular $v(s)\geq cm$ and
$m\cdot\mathrm{lct}_{x}(X,\Delta;\{s=0\})\leq m\cdot\frac{A_{X,\Delta}(v)}{v(s)}\leq\frac{1}{c}=\left(\frac{\widehat{\rm vol}(x,X,\Delta)}{c_{1}}\right)^{1/n}.$
This completes the proof. ∎
Proof of Proposition 3.3.
First assume that $L$ is $\mathbb{Q}$-Cartier. For each sufficiently divisible positive integer $m$, after choosing a local trivialization of $mL$ around $x\in X$, we may view each $V_{m}$ as a subspace of $\mathcal{O}_{X,x}$. Let $0<\varepsilon\ll 1$ and set $c_{0}=\mathrm{vol}(V_{\bullet})-\varepsilon$ and $c_{1}=\mathrm{vol}(V_{\bullet})-2\varepsilon$. Then $\dim V_{m}\geq c_{0}\frac{m^{n}}{n!}$ when $m\gg 0$. Thus Lemma 3.4 gives $\widehat{\rm vol}(x,X,\Delta)\geq c_{1}\cdot\alpha(X,\Delta;V_{\bullet})^{n}$. Letting $\varepsilon\to 0$ and hence $c_{1}\to\mathrm{vol}(V_{\bullet})$, we get the desired result.
In general, choose some Cartier divisor $D\geq 0$ on $X$ such that $\mathrm{Supp}(L)\subseteq\mathrm{Supp}(D)$. For any $\varepsilon>0$, there exists some $\mathbb{Q}$-Cartier $\mathbb{Q}$-divisor $L^{\prime}$ such that $L\leq L^{\prime}\leq L+\varepsilon D$. Let $V^{\prime}_{\bullet}$ be the image of $V_{\bullet}$ under the natural embedding $H^{0}(mL)\hookrightarrow H^{0}(mL^{\prime})$. Since $L^{\prime}$ is $\mathbb{Q}$-Cartier, the previous case implies that
$\widehat{\rm vol}(x,X,\Delta)\geq\alpha(X,\Delta;V^{\prime}_{\bullet})^{n}\cdot\mathrm{vol}(V^{\prime}_{\bullet})=\alpha(X,\Delta;V^{\prime}_{\bullet})^{n}\cdot\mathrm{vol}(V_{\bullet}).$
On the other hand, as $|V^{\prime}_{m}|=|V_{m}|+m(L^{\prime}-L)$ and $0\leq L^{\prime}-L\leq\varepsilon D$, by Lemma 2.6, we deduce that
$\left(\frac{1}{\alpha(X,\Delta;V_{\bullet})}+\frac{\varepsilon}{\mathrm{lct}(X,\Delta;D)}\right)^{-1}\leq\alpha(X,\Delta;V^{\prime}_{\bullet})\leq\alpha(X,\Delta;V_{\bullet}),$
hence $\alpha(X,\Delta;V^{\prime}_{\bullet})$ can be arbitrarily close to $\alpha(X,\Delta;V_{\bullet})$ when $\varepsilon\to 0$. The proposition thus follows. ∎
In the context of Theorem 3.2, the linear systems we shall use come from the following construction.
Lemma 3.5.
Let $(X,\Delta+D)$ be a klt pair such that $\Delta,D\geq 0$ and $D$ is big. Then there exists an eventually birational graded linear series of $D$ such that $(X,\Delta+\Gamma)$ is log canonical for all $m\in M(V_{\bullet})$ and all $\Gamma\in\frac{1}{m}|V_{m}|$.
Proof.
Since $D$ is big, we may write $D=A+B$ where $A$ is an ample $\mathbb{Q}$-divisor and $B\geq 0$. Let $\overline{X}$ be a normal projective compactification of $X$ such that $A$ extends to an ample $\mathbb{Q}$-divisor $\overline{A}$ (for example, consider the morphism from $X$ into some projective space induced by the linear system $|qA|$ for some sufficiently positive and divisible integer $q$, and take $\overline{X}$ to be the normalization of the closure of the image). Then $D$ also extends to a big $\mathbb{R}$-divisor $\overline{D}=\overline{A}+\overline{B}$ on $\overline{X}$, where $\overline{B}$ is the closure of $B$. Let $V_{m}\subseteq H^{0}(X,mD)$ be the image of the restriction map $H^{0}(\overline{X},m\overline{D})\to H^{0}(X,mD)$. Since $(X,\Delta+D)$ is klt, by Lemma 2.7 applied to some log resolution of the pair $(X,\Delta+D)$, we know that there exists some sufficiently large integer $\ell>0$ such that $(X,\Delta+\left(1-\frac{1}{\ell}\right)D+\frac{1}{\ell}G)$ is log canonical for all $m\in\mathbb{N}^{+}$ and all $G\in\frac{1}{m}|V_{m}|$. Define a new graded linear series $V^{\prime}_{\bullet}$ of $D$ by
$|V^{\prime}_{m}|=m\left(1-\frac{1}{\ell}\right)D+|V_{m/\ell}|$
when $\ell$ divides $m$, and $V^{\prime}_{m}=0$ otherwise. Since $\overline{D}$ is big, both $V_{\bullet}$ and $V^{\prime}_{\bullet}$ are eventually birational. By construction, $(X,\Delta+\Gamma)$ is log canonical for all $m\in\ell\cdot\mathbb{N}^{+}$ and $\Gamma\in\frac{1}{m}|V^{\prime}_{m}|$. Thus $V^{\prime}_{\bullet}$ satisfies the desired conditions. ∎
We are now ready to prove the theorem.
Proof of Theorem 3.2.
We first prove (3). Since $K_{X^{\prime}}+\Delta^{\prime}$ is $\mathbb{R}$-Cartier, so is $D^{\prime}$. By Lemma 3.5, we can find an eventually birational graded linear series $V_{\bullet}$ of $D$ such that $(X,\Delta+\Gamma)$ is log canonical for all $m\in M(V_{\bullet})$ and all $\Gamma\in\frac{1}{m}|V_{m}|$. Since $\Gamma\sim_{\mathbb{R}}D$ and $\varphi$ is an MMP type contraction between $(X,\Delta+D)$ and $(X^{\prime},\Delta^{\prime}+D^{\prime})$, by Lemma 2.13 it is also an MMP type contraction between $(X,\Delta+\Gamma)$ and $(X^{\prime},\Delta^{\prime}+\Gamma^{\prime})$ where $\Gamma^{\prime}:=\varphi_{*}\Gamma$. It follows that $(X^{\prime},\Delta^{\prime}+\Gamma^{\prime})$ is log canonical and thus $\alpha(X^{\prime},\Delta^{\prime};\varphi_{*}V_{\bullet})\geq 1$. Note that $\varphi_{*}V_{\bullet}$ is a graded linear series of $D^{\prime}$. By Proposition 3.3, we obtain
$\widehat{\rm vol}(x^{\prime},X^{\prime},\Delta^{\prime})\geq\mathrm{vol}(\varphi_{*}V_{\bullet})=\mathrm{vol}(V_{\bullet})$
for all closed points $x^{\prime}\in X^{\prime}$. Hence (3) holds with $\varepsilon:=\mathrm{vol}(V_{\bullet})>0$.
Next we prove (1). Let $n=\dim X$. Take a small $\mathbb{Q}$-factorial modification $\pi\colon X^{\prime\prime}\to X^{\prime}$ (which exists by [BCHM]*Corollary 1.4.3) and let $\Delta^{\prime\prime},D^{\prime\prime}$ be the strict transforms of $\Delta^{\prime},D^{\prime}$ respectively. Then $K_{X^{\prime\prime}}+\Delta^{\prime\prime}$ is $\mathbb{R}$-Cartier and the induced map $(X,\Delta+D)\dashrightarrow(X^{\prime\prime},\Delta^{\prime\prime}+D^{\prime\prime})$ is also an MMP type contraction. By the base point free theorem [KM98]*Theorem 3.24, the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor $L^{\prime}$ on $X^{\prime}$ is the same as the Cartier index of $\pi^{*}L^{\prime}$ on $X^{\prime\prime}$. Now for any closed point $x^{\prime\prime}\in X^{\prime\prime}$ we have $\widehat{\rm vol}(x^{\prime\prime},X^{\prime\prime},\Delta^{\prime\prime})\geq\varepsilon$ by (3) and hence the Cartier index of $\pi^{*}L^{\prime}$ is at most $\left\lfloor\frac{n^{n}}{\varepsilon}\right\rfloor$ at $x^{\prime\prime}$ by Theorem 2.4. Thus (1) holds with $r=\left\lfloor\frac{n^{n}}{\varepsilon}\right\rfloor!$.
Finally, by Theorem 2.5, there exists some constant $A>0$ (depending only on $n,\varepsilon$) such that $\mathrm{mld}(x^{\prime\prime},X^{\prime\prime},\Delta^{\prime\prime})\leq A$ for all points $x^{\prime\prime}\in X^{\prime\prime}$. As $\pi$ is small, we have $\pi^{*}(K_{X^{\prime}}+\Delta^{\prime}+D^{\prime})=K_{X^{\prime\prime}}+\Delta^{\prime\prime}+D^{\prime\prime}$, hence
$\mathrm{mld}(x^{\prime},X^{\prime},\Delta^{\prime}+D^{\prime})\leq\mathrm{mld}(x^{\prime\prime},X^{\prime\prime},\Delta^{\prime\prime}+D^{\prime\prime})\leq\mathrm{mld}(x^{\prime\prime},X^{\prime\prime},\Delta^{\prime\prime})$
for any $x^{\prime\prime}\in\pi^{-1}(x)$. Thus it also holds that $\mathrm{mld}(x^{\prime},X^{\prime},\Delta^{\prime}+D^{\prime})\leq A$, which gives (2). ∎
Remark 3.6.
It is not clear if the local volume satisfies any monotonicity in the MMP; at least the naïve versions fail. For example, there exist flops $X_{1}\dashrightarrow X_{2}$ and valuations $v$ whose center is a closed point on both $X_{1}$ and $X_{2}$ but $\widehat{\rm vol}_{X_{1}}(v)\neq\widehat{\rm vol}_{X_{2}}(v)$. To get an explicit example, let $Z$ be the cone over the smooth quadric surface $Q\cong\mathbb{P}^{1}\times\mathbb{P}^{1}\subseteq\mathbb{P}^{3}$ and consider the Atiyah flop $X_{1}\dashrightarrow X_{2}$ between its two small resolutions. Then $Z$ is a toric variety and most toric valuations $v$ satisfy $\widehat{\rm vol}_{X_{1}}(v)\neq\widehat{\rm vol}_{X_{2}}(v)$. By adding appropriate boundaries to $X_{i}$ we can turn the flop $X_{1}\dashrightarrow X_{2}$ into either a flip or an anti-flip. Thus there exist flips $(X,\Delta)\dashrightarrow(X^{+},\Delta^{+})$ where the normalized volume of valuations increases (resp. decreases).
Alternatively, one may ask whether the smallest local volume $\inf_{x\in X}\widehat{\rm vol}(x,X,\Delta)$ of the klt pair $(X,\Delta)$ is monotone in the MMP. The answer is again negative. Indeed, by [LX-cubic-3fold]*Theorem A.4, we have $\inf_{x\in X}\widehat{\rm vol}(x,X,\Delta)\leq n^{n}$ with equality if and only if $X$ is smooth and $\Delta=0$. There are flips $X\dashrightarrow X^{+}$ such that $X$ is singular (resp. smooth) whereas $X^{+}$ is smooth (resp. singular), in which case the smallest local volume increases (resp. decreases) after the flips. For some explicit examples, consider the total space $\widetilde{X}$ of the line bundle $L=\mathcal{O}_{\mathbb{P}^{m}\times\mathbb{P}^{n}}(-1,-2)$ on $\mathbb{P}^{m}\times\mathbb{P}^{n}$. Let $X_{1}$ (resp. $X_{2}$) be obtained from $\widetilde{X}$ by contracting the zero section $E\cong\mathbb{P}^{m}\times\mathbb{P}^{n}$ using the first (resp. second) projection of $\mathbb{P}^{m}\times\mathbb{P}^{n}$. Then $X_{1}$ is smooth and $X_{2}$ is singular. On the other hand, a direct calculation shows that the induced birational map $X_{1}\dashrightarrow X_{2}$ (resp. $X_{2}\dashrightarrow X_{1}$) is a flip if and only if $2m+1<n$ (resp. $2m+1>n$).
3.2. Log canonical volume
Although the local volume does not seem to satisfy monotonicity in the MMP, we can nonetheless introduce a global variant that does. The definition is motivated by the argument in the previous subsection. We begin with an auxiliary definition.
Definition 3.7.
Let $(X,\Delta)$ be a pair and let $L$ be a big $\mathbb{R}$-divisor such that $K_{X}+\Delta+L$ is $\mathbb{R}$-Cartier. A graded linear series $V_{\bullet}$ of $L$ is called admissible (with respect to $(X,\Delta)$) if it is eventually birational and $(X,\Delta+\Gamma)$ is log canonical for all $m\in M(V_{\bullet})$ and all $\Gamma\in\frac{1}{m}|V_{m}|$. Note that the latter condition is equivalent to $\alpha(X,\Delta;V_{\bullet})\geq 1$ when $K_{X}+\Delta$ is $\mathbb{R}$-Cartier.
Lemma 3.8.
Let $(X,\Delta)$ and $L$ be as in Definition 3.7. Then an admissible graded linear series of $L$ exists if and only if there exists some $0\leq D\sim_{\mathbb{R}}L$ such that $(X,\Delta+D)$ is klt.
Proof.
If there exists some $0\leq D\sim_{\mathbb{R}}L$ such that $(X,\Delta+D)$ is klt, then we can choose such $D$ with $D\sim_{\mathbb{Q}}L$. Thus one direction is implied by Lemma 3.5. For the other direction, let $V_{\bullet}$ be an admissible graded linear series of $L$. By assumption, there exists some $m\in\mathbb{N}_{+}$ such that the rational map $f\colon X\dashrightarrow\mathbb{P}^{N}$ induced by the linear system $|V_{m}|$ is birational onto its image. Let $\pi\colon X^{\prime}\to X$ be a log resolution of $(X,\Delta)$ that resolve the indeterminacy of $f$ and let $g\colon X^{\prime}\to\mathbb{P}^{N}$ the induced morphism. For any prime divisor $F$ on $X^{\prime}$, let
$a_{F}:=-\sup\left\{a(F,X,\Delta+D)\middle|\ D\in\frac{1}{m}|V_{m}|\right\},$
and let $\Delta^{\prime}:=\sum a_{F}F$. Note that this is a finite sum since $a_{F}=0$ unless $F$ is $\pi$-exceptional or the strict transform of some component of $\Delta+D$ for any $D\in|V_{m}|$. We also note that $\pi_{*}\Delta^{\prime}-\Delta$ is the fixed part of $|V_{m}|$.
Let $L^{\prime}:=\pi^{*}(K_{X}+\Delta+L)-(K_{X^{\prime}}+\Delta^{\prime})$. Then the linear series $V_{m}$ of $mL$ induces a linear series $V^{\prime}_{m}$ of $mL^{\prime}$ defined by
$|V^{\prime}_{m}|=\{\pi^{*}(mK_{X}+m\Delta+D)-mK_{X^{\prime}}-m\Delta^{\prime}\,|\,D\in|V_{m}|\},$
and the morphism $g$ is induced by $|V^{\prime}_{m}|$. In particular, the linear system $|V^{\prime}_{m}|$ is base point free. We claim that the sub-pair $(X^{\prime},\Delta^{\prime})$ is klt. In fact, since $V_{\bullet}$ is admissible, we know that $(X,\Delta+D)$ is log canonical for any positive integer $\ell$ and all $D\in\frac{1}{\ell m}|V_{\ell m}|$. This implies that for any hypersurface $H\subseteq\mathbb{P}^{N}$ of degree $\ell$ whose support does not contain $g(X^{\prime})$, the sub-pair $(X^{\prime},\Delta^{\prime}+\frac{1}{\ell m}g^{*}H)$ is log canonical, as $\frac{1}{\ell m}g^{*}H\sim L^{\prime}$ and thus the sub-pair is the crepant pullback of $(X,\Delta+D)$ where $D=\pi_{*}\Delta^{\prime}-\Delta+\frac{1}{\ell m}\pi_{*}g^{*}H\in\frac{1}{\ell m}|V_{\ell m}|$. If $(X^{\prime},\Delta^{\prime})$ is not klt, then there exists some prime divisor $E$ over $X^{\prime}$ such that $A_{X^{\prime},\Delta^{\prime}}(E)\leq 0$. As $g$ is a birational morphism, we can choose some hypersurface $H$ as above of sufficiently large degree such that $g^{*}H$ contains the center of $E$. But then $A_{X^{\prime},\Delta^{\prime}+\frac{1}{\ell m}g^{*}H}(E)<0$ and hence $(X^{\prime},\Delta^{\prime}+\frac{1}{\ell m}g^{*}H)$ cannot be log canonical, a contradiction. This proves the lemma. ∎
We now introduce the global invariant that behaves well in the MMP.
Definition 3.9.
Let $(X,\Delta)$ be a pair and let $L$ be a big $\mathbb{R}$-divisor such that $K_{X}+\Delta+L$ is $\mathbb{R}$-Cartier. We define the log canonical volume of $L$ with respect to $(X,\Delta)$, denoted by $\widehat{\rm Vol}_{X,\Delta}(L)$, as follows:
  1. (1)
    If $(X,\Delta+D)$ is klt for some $0\leq D\sim_{\mathbb{R}}L$, we set
    $\widehat{\rm Vol}_{X,\Delta}(L):=\sup\{\mathrm{vol}(V_{\bullet})\,|\,V_{\bullet}\mbox{ is an admissible graded linear series of }L\}.$
    In this case $\widehat{\rm Vol}_{X,\Delta}(L)>0$ by Lemma 3.5.
  2. (2)
    Otherwise, by Lemma 3.8, there are no admissible graded linear series of $L$ and we set $\widehat{\rm Vol}_{X,\Delta}(L):=0$.
While we do not need the next statement in this paper, we note that the log canonical volume defined above only depends on the numerical equivalence class of $L$. Throughout the rest of this subsection, let $n=\dim X$.
Lemma 3.10.
Let $(X,\Delta)$ and $L$ be as in Definition 3.9, and let $L^{\prime}$ be another $\mathbb{R}$-divisor on $X$ such that $L^{\prime}-L$ extends to an $\mathbb{R}$-divisor in $\mathrm{Pic}^{0}(\overline{X})\otimes\mathbb{R}$ for some projective compactification $\overline{X}$ of $X$. Then
$\widehat{\rm Vol}_{X,\Delta}(L)=\widehat{\rm Vol}_{X,\Delta}(L^{\prime}).$
Proof.
It suffices to show that $\widehat{\rm Vol}_{X,\Delta}(L^{\prime})\geq\widehat{\rm Vol}_{X,\Delta}(L)$, as by symmetry we also get the reverse inequality. To prove this, we may assume that there exists some big $\mathbb{R}$-divisor $0\leq D\sim_{\mathbb{R}}L$ such that $(X,\Delta+D)$ is klt, and it is enough to show that $\widehat{\rm Vol}_{X,\Delta}(L^{\prime})\geq\left(1-\frac{1}{\ell}\right)^{n}\mathrm{vol}(V_{\bullet})$ for any sufficiently large integer $\ell$ and any admissible graded linear series $V_{\bullet}$ of $L$. Since $L^{\prime}-L\in\mathrm{Pic}^{0}(\overline{X})\otimes\mathbb{R}$, we can find some $\mathbb{R}$-divisor $0\leq D^{\prime}\sim_{\mathbb{Q}}\ell(L^{\prime}-L)+L$ such that $(X,\Delta+D^{\prime})$ is also klt. To see this, write $D=A+B$ where $A$ is ample and $B\geq 0$ such that the pair $(X,\Delta+B)$ is also klt, then $A+\ell(L^{\prime}-L)$ is still ample and we can take $D^{\prime}=A^{\prime}+B$ for some general $0\leq A^{\prime}\sim_{\mathbb{Q}}A+\ell(L^{\prime}-L)$. By construction, $(1-\frac{1}{\ell})L+\frac{1}{\ell}D^{\prime}\sim_{\mathbb{Q}}L^{\prime}$. For any sufficiently divisible integer $m$, let $V^{\prime}_{m}$ be the linear series of $mL^{\prime}$ defined by $|V^{\prime}_{m}|=|V_{(1-\frac{1}{\ell})m}|+\frac{m}{\ell}D^{\prime}$. Then $V^{\prime}_{\bullet}$ is admissible by the convexity of log canonicity. It follows that $\widehat{\rm Vol}_{X,\Delta}(L^{\prime})\geq\mathrm{vol}(V^{\prime}_{\bullet})=\left(1-\frac{1}{\ell}\right)^{n}\mathrm{vol}(V_{\bullet})$ as desired. ∎
The proof of Theorem 3.2 can be reformulated in terms of some properties of the log canonical volume $\widehat{\rm Vol}_{X,\Delta}(L)$. These include its monotonicity in the MMP, and its comparison with the local volume.
Lemma 3.11.
Let $\varphi\colon(X,\Delta+D)\dashrightarrow(X^{\prime},\Delta^{\prime}+D^{\prime})$ be an MMP type contraction between klt pairs. Assume that $\Delta,\Delta^{\prime}\geq 0$, $D$ is big, and $D^{\prime}=\varphi_{*}D$. Then $\widehat{\rm Vol}_{X^{\prime},\Delta^{\prime}}(D^{\prime})\geq\widehat{\rm Vol}_{X,\Delta}(D)$.
Proof.
If $V_{\bullet}$ is an admissible graded linear series of $D$, then $\varphi_{*}V_{\bullet}$ is an admissible graded linear series of $D^{\prime}$ as in the proof of Theorem 3.2. ∎
Lemma 3.12.
Let $(X,\Delta)$ be a klt pair and let $D\geq 0$ be a big $\mathbb{R}$-Cartier $\mathbb{R}$-divisor on $X$. Then for any closed point $x\in X$ we have $\widehat{\rm vol}(x,X,\Delta)\geq\widehat{\rm Vol}_{X,\Delta}(D)$.
Proof.
This follows immediately from Proposition 3.3. ∎
We make a few more elementary observations on the log canonical volume $\widehat{\rm Vol}_{X,\Delta}(L)$ that will be useful in the next several sections.
Lemma 3.13.
Let $(X,\Delta)$ be a klt pair such that $\Delta$ is big. Then the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor on $X$ is at most $\left\lfloor\frac{n^{n}}{\widehat{\rm Vol}_{X}(\Delta)}\right\rfloor!$.
Proof.
Let $X^{\prime}\to X$ be a small $\mathbb{Q}$-factorial modification and let $\Delta^{\prime}$ be the strict transform of $\Delta$. By Lemma 3.12, we have $\widehat{\rm vol}(x^{\prime},X^{\prime})\geq\widehat{\rm Vol}_{X^{\prime}}(\Delta^{\prime})$ for any closed point $x^{\prime}\in X^{\prime}$, hence by Theorem 2.4 we know that the Cartier index of any Weil divisor on $X^{\prime}$ is at most $\left\lfloor\frac{n^{n}}{\widehat{\rm Vol}_{X^{\prime}}(\Delta^{\prime})}\right\rfloor!$. Note that $\widehat{\rm Vol}_{X^{\prime}}(\Delta^{\prime})=\widehat{\rm Vol}_{X}(\Delta)$ by Lemma 3.11. The result then follows from the base point free theorem as in the proof of Theorem 3.2. ∎
Lemma 3.14.
Let $(X,\Delta)$ be a klt pair such that $\Delta$ is big, let $\lambda\in[0,1)$, and let $0\leq\Delta_{0}\leq\lambda\Delta$ be an $\mathbb{R}$-divisor. Then
$\widehat{\rm Vol}_{X}(\Delta)\geq\widehat{\rm Vol}_{X,\Delta_{0}}(\Delta-\Delta_{0})\geq(1-\lambda)^{n}\widehat{\rm Vol}_{X}(\Delta).$
Proof.
If $V_{\bullet}$ is an admissible graded linear series of $\Delta-\Delta_{0}$ with respect to $(X,\Delta_{0})$, then its image $V^{\prime}_{\bullet}$ under the natural inclusion $H^{0}(X,m(\Delta-\Delta_{0}))\hookrightarrow H^{0}(X,m\Delta)$ is an admissible graded linear series of $\Delta$ with respect to $X$. This gives the first inequality. To prove the second inequality we may assume that $\lambda\in\mathbb{Q}$. For any admissible graded linear series $V_{\bullet}$ of $\Delta$, let $V^{\prime}_{\bullet}$ be the graded linear series of $\Delta-\Delta_{0}$ defined by $|V^{\prime}_{m}|=|V_{(1-\lambda)m}|+m(\lambda\Delta-\Delta_{0})$ for sufficiently divisible $m\in\mathbb{N}$. By convexity of log canonicity, we know that $(X,\lambda\Delta+(1-\lambda)D)$ is log canonical for all $m\in M(V_{\bullet})$ and $D\in\frac{1}{m}|V_{m}|$. This implies that $V^{\prime}_{\bullet}$ is admissible with respect to $(X,\Delta_{0})$. Since $V_{\bullet}$ is arbitrary and $\mathrm{vol}(V^{\prime}_{\bullet})=(1-\lambda)^{n}\mathrm{vol}(V_{\bullet})$, the lemma follows. ∎
Lemma 3.15.
Let $(X,\Delta+D)$ be a klt pair where $\Delta\geq 0$ and $D$ is big. Let $H$ be an ample $\mathbb{R}$-divisor on $X$. Then
$\widehat{\rm Vol}_{X,\Delta}(D+H)\geq\widehat{\rm Vol}_{X,\Delta}(D).$
Proof.
Let $\ell$ be a sufficiently large and divisible integer. It is enough to show that $\widehat{\rm Vol}_{X,\Delta}(D+H)\geq(1-\frac{1}{\ell})^{n}\mathrm{vol}(V_{\bullet})$ for any admissible graded linear series $V_{\bullet}$ of $D$. Since $(X,\Delta+D)$ is klt, $H$ is ample, and $\ell$ is sufficiently large and divisible, by Bertini’s theorem we can find some $H_{0}\in|\ell H|$ such that $(X,\Delta+D+H_{0})$ is log canonical. Define a graded linear series $V^{\prime}_{\bullet}$ of $D+H$ by $|V^{\prime}_{m}|=|V_{\left(1-\frac{1}{\ell}\right)m}|+\frac{m}{\ell}(D+H_{0})$ for all $m\in\ell\cdot\mathbb{N}$. Then $V^{\prime}_{\bullet}$ is admissible by convexity of log canonicity as before and this proves the result. ∎
We also prove a semi-continuity result for the log canonical volume, which allows us to reduce many questions to the case of rational coefficients. Let $\mathrm{Div}(X)$ be the free group generated by the prime divisors on $X$.
Lemma 3.16.
Let $(X,\Delta)$ be a pair, let $\Lambda\subseteq\mathrm{Div}(X)_{\mathbb{R}}$ be a finite dimensional subspace, and let $\mathcal{L}(\Lambda)\subseteq\Lambda$ be the subset of $\mathbb{R}$-divisors $D\geq 0$ such that $(X,\Delta+D)$ is klt and $D$ is big. Then the function $D\mapsto\widehat{\rm Vol}_{X,\Delta}(D)$ is lower semi-continuous in $\mathcal{L}(\Lambda)$.
Proof.
Let $\ell\in\mathbb{N}$ be a sufficiently large integer. It suffices to show that $\widehat{\rm Vol}_{X,\Delta}(D^{\prime})\geq\left(1-\frac{1}{\ell}\right)^{n}\widehat{\rm Vol}_{X,\Delta}(D)$ whenever $D^{\prime}\in\mathcal{L}(\Lambda)$ is sufficiently close to $D$. To see this, we may assume that $D^{\prime}\geq\left(1-\frac{1}{\ell}\right)D$ and
$\left(X,\Delta+\ell\left(D^{\prime}-\left(1-\frac{1}{\ell}\right)D\right)\right)=(X,\Delta+D+\ell(D^{\prime}-D))$
is klt. If $V_{\bullet}$ is an admissible graded linear series of $D$, and define
$\left|V^{\prime}_{m}\right|=\left|V_{\left(1-\frac{1}{\ell}\right)m}\right|+mD^{\prime}-m\left(1-\frac{1}{\ell}\right)D$
when $\ell$ divides $m$ (otherwise set $V^{\prime}_{m}=0$), then by convexity of log canonicity as above we see that $V^{\prime}_{\bullet}$ is an admissible graded linear series of $D^{\prime}$ and by construction $\mathrm{vol}(V^{\prime}_{\bullet})\geq(1-\frac{1}{\ell})^{n}\mathrm{vol}(V_{\bullet})$. This proves the result. ∎
3.3. Explicit bounds
In the remaining part of this section, we prove some effective versions of Theorem 3.2, giving explicit bounds on the invariants at least when the pair is log smooth and projective (in general, one can reduce to this case by passing to a log resolution of some projective compactification). We also prove an explicit index bound for $3$-fold MMP. These will be useful in Section 6 where we bound the number of steps in certain MMP.
We start with an explicit bound of the log canonical volume.
Lemma 3.17.
Let $(X,\Delta)$ be a projective log smooth klt pair of dimension $n$ such that $\Delta$ is big, and let $H$ be a very ample divisor on $X$. Then
$\widehat{\rm Vol}_{X}(\Delta)\geq\left\lceil\frac{(\Delta\cdot H^{n-1})}{1-\max\mathrm{Coeff}(\Delta)}\right\rceil^{-n}\cdot\mathrm{vol}(\Delta)$
Proof.
Since $H$ is very ample, we have $\mathrm{mult}_{x}D\leq(\Delta\cdot H^{n-1})$ for any $x\in X$ and any $0\leq D\sim_{\mathbb{R}}\Delta$. By Lemma 2.7, this gives
$\mathrm{lct}(X,\Delta;D)\geq\frac{1-\max\mathrm{Coeff}(\Delta)}{(\Delta\cdot H^{n-1})}.$
Thus in the proof of Lemma 3.5 we can take $\ell$ to be $\left\lceil\frac{(\Delta\cdot H^{n-1})}{1-\max\mathrm{Coeff}(\Delta)}\right\rceil$. It then follows from the argument in loc. cit. that there exists an admissible linear series $V_{\bullet}$ of $\Delta$ with $\mathrm{vol}(V_{\bullet})\geq\ell^{-n}\mathrm{vol}(\Delta)$, hence $\widehat{\rm Vol}_{X}(\Delta)\geq\ell^{-n}\mathrm{vol}(\Delta)$ as desired. ∎
Corollary 3.18.
Let $(X,\Delta)$ be a projective log smooth klt pair of dimension $n$, and let $N$ be a positive integer. Assume that $N\Delta$ has integer coefficients, $\mathrm{vol}(\Delta)\geq\frac{1}{N}$, and $(\Delta\cdot H^{n-1})\leq N$ for some very ample divisor $H$ on $X$. Then in any $(K_{X}+\Delta)$-MMP, the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor is at most $(n^{n}N^{2n+1})!$.
Proof.
By assumption, we have $1-\max\mathrm{Coeff}(\Delta)\geq\frac{1}{N}$. Hence by Lemma 3.17 and our assumptions, we obtain $\widehat{\rm Vol}_{X}(\Delta)\geq N^{-2n-1}$. The result then follows from Lemmas 3.11 and 3.13. ∎
By a similar argument, we can also bound the index explicitly in a general type MMP.
Lemma 3.19.
Let $(X,\Delta)$ be a projective klt pair of dimension $n$ such that $K_{X}+\Delta$ is big. Then for any sequence of steps of a $(K_{X}+\Delta)$-MMP $\varphi\colon(X,\Delta)\dashrightarrow(Y,\Delta_{Y})$ and any closed point $y\in Y$, we have
$\widehat{\rm vol}(y,Y,\Delta_{Y})\geq\alpha(X,\Delta;K_{X}+\Delta)^{n}\cdot\mathrm{vol}(K_{X}+\Delta).$
Proof.
Let $\alpha:=\alpha(X,\Delta;K_{X}+\Delta)$ and let $V_{\bullet}$ be the complete linear series of $K_{X}+\Delta$, i.e. $V_{m}=H^{0}(X,m(K_{X}+\Delta))$. Then for any positive integer $m$ and any $D\in\frac{1}{m}|V_{m}|$, the pair $(X,\Delta+\alpha D)$ is log canonical, and the $(K_{X}+\Delta)$-MMP is also a $(K_{X}+\Delta+\alpha D)$-MMP. It follows that $(Y,\Delta_{Y}+\alpha D_{Y})$ is also log canonical, hence $\alpha(Y,\Delta;\varphi_{*}V_{\bullet})\geq\alpha$. Note that $\mathrm{vol}(\varphi_{*}V_{\bullet})=\mathrm{vol}(K_{X}+\Delta)$. The result then follows from Proposition 3.3. ∎
Corollary 3.20.
Let $(X,\Delta)$ be a projective log smooth klt pair of dimension $n$, and let $N$ be a positive integer. Assume that $N\Delta$ has integer coefficients, $\mathrm{vol}(K_{X}+\Delta)\geq\frac{1}{N}$, and $((K_{X}+\Delta)\cdot H^{n-1})\leq N$ for some very ample divisor $H$ on $X$. Then in any $(K_{X}+\Delta)$-MMP, the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor is at most $(n^{n}N^{2n+1})!$.
Proof.
By Lemma 2.7, we have
$\alpha(X,\Delta;K_{X}+\Delta)\geq\frac{1-\max\mathrm{Coeff}(\Delta)}{((K_{X}+\Delta)\cdot H^{n-1})}\geq\frac{1}{N^{2}}$
as in the proof of Lemma 3.17. We then conclude by Lemma 3.19 and Theorem 2.4. ∎
Finally, we show that the Cartier index is uniformly bounded in $3$-dimensional MMP (without any general type or bigness assumption). For any klt pair $(X,\Delta)$, we set
(3.1) $e(X,\Delta):=\#\left\{\mbox{Exceptional prime divisors $E$ over $X$ with $a(E,X,\Delta)\leq 0$}\right\}.$
We also define $e_{+}(X,\Delta)$ as the sum of $e(X,\Delta)$ and the number of irreducible components of $\mathrm{Supp}(\Delta)$. By [KM98, Proposition 2.36], $e(X,\Delta)<+\infty$.
Lemma 3.21.
Let $(X,\Delta)$ be a klt pair of dimension $3$. Let $m\in\mathbb{N}$ and $\varepsilon>0$. Assume that there are at most $m$ divisors $E$ over $X$ such that $a(E,X,\Delta)\in(0,\varepsilon)$. Then the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor on $X$ is at most
$\left(\left\lceil\frac{m+2}{\varepsilon}\right\rceil!\right)^{2^{e(X,\Delta)}}\left(\frac{3}{\mathrm{mld}(X,\Delta)}\right)^{2^{e(X,\Delta)}-1}.$
Proof.
This follows from the same proof of [K+-flip-abundance]*Theorem (6.11.5), replacing the base case (i.e. when $e(X,\Delta)=0$ and hence $(X,\Delta)$ is terminal) [K+-flip-abundance]*Lemma 6.9.7 of the induction by [Sho96]*Corollary 3.4. ∎
Corollary 3.22.
Let $(X,\Delta)$ be a klt pair of dimension $3$. Let $\varepsilon$ be the smallest positive discrepancy of $(X,\Delta)$ (i.e. for any prime divisor $E$ over $X$ we have $a(E,X,\Delta)>0$ if and only if $a(E,X,\Delta)\geq\varepsilon$). Then in any $(K_{X}+\Delta)$-MMP, the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor is at most
$\left(\left\lceil\frac{e_{+}(X,\Delta)+3}{\varepsilon}\right\rceil!\right)^{2^{1+e_{+}(X,\Delta)}}.$
Proof.
Let $(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ be a sequence of steps of a $(K_{X}+\Delta)$-MMP. We first prove that $\mathrm{mld}(X,\Delta)\geq\varepsilon$. In fact, by passing to the terminalization, it suffices to prove this when $(X,\Delta)$ is terminal. Write $\Delta=\sum b_{i}B_{i}$ where the $B_{i}$ are irreducible components of $\mathrm{Supp}(\Delta)$, then $\mathrm{mld}(X,\Delta)=1-\max_{i}\{b_{i}\}$. On the other hand, if we blow up a prime divisor in the smooth locus of $B_{i}$, we get an exceptional divisor $E$ with $a(E,X,\Delta)=1-b_{i}>0$, thus as $\varepsilon$ is the smallest positive discrepancy we have $1-b_{i}\geq\varepsilon$. This gives $\mathrm{mld}(X,\Delta)\geq\varepsilon$ and hence we also have $\mathrm{mld}(X^{\prime},\Delta^{\prime})\geq\varepsilon$.
Let $e:=e_{+}(X,\Delta)$. By definition, every divisor $E$ over $X$ with $a(E,X,\Delta)<\varepsilon$ also satisfies $a(E,X,\Delta)\leq 0$. As $(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ is a $(K_{X}+\Delta)$-MMP, we have $a(E,X,\Delta)\leq a(E,X^{\prime},\Delta^{\prime})$, hence any divisor $E$ with $a(E,X^{\prime},\Delta^{\prime})<\varepsilon$ also satisfies $a(E,X,\Delta)\leq 0$. If $E$ is exceptional over $X^{\prime}$, then either it is also exceptional over $X$, or it is a component of $\mathrm{Supp}(\Delta)$. It follows that there are at most $e$ exceptional divisors $E$ over $X^{\prime}$ with $a(E,X^{\prime},\Delta^{\prime})<\varepsilon$; in particular, there are at most $e$ divisors $E$ with $0<a(E,X^{\prime},\Delta^{\prime})<\varepsilon$ (all such divisors are necessarily exceptional over $X^{\prime}$). Since we also have $e(X^{\prime},\Delta^{\prime})\leq e_{+}(X^{\prime},\Delta^{\prime})\leq e_{+}(X,\Delta)=e$ and $\mathrm{mld}(X^{\prime},\Delta^{\prime})\geq\varepsilon$, Lemma 3.21 implies that the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor on $X^{\prime}$ is at most
$\left(\left\lceil\frac{e+2}{\varepsilon}\right\rceil!\right)^{2^{e}}\cdot\left(\left\lceil\frac{3}{\varepsilon}\right\rceil!\right)^{2^{e-1}}\leq\left(\left\lceil\frac{e+3}{\varepsilon}\right\rceil!\right)^{2^{e+1}}.$
4. Boundedness of fibers
In this section, we improve the argument from the previous section to prove the boundedness of the fibers of the extremal contractions and flips in the MMP.
Theorem 4.1.
Let $(X,\Delta)$ be a $\mathbb{Q}$-factorial klt pair and $X\rightarrow T$ a projective morphism. Assume that $K_{X}+\Delta$ or $\Delta$ is big over $T$. Then there exists a projective family $\mathcal{W}\to\mathcal{B}$ over a finite type base $\mathcal{B}$, such that in any sequence of a $(K_{X}+\Delta)$-MMP over $T$, every fiber of the extremal contractions or the flips is isomorphic to $\mathcal{W}_{b}$ for some $b\in\mathcal{B}$.
Recall that a projective morphism $f\colon(X,\Delta)\to Z$ is called an extremal contraction if it is the contraction of a $(K_{X}+\Delta)$-negative extremal ray. There are three types of extremal contractions: divisorial contractions, flipping contractions, and Mori fiber spaces.
4.1. Strategy
Our strategy towards Theorem 4.1 is to first prove that the germs of the extremal contractions and the flips are bounded up to special degenerations, and then analyze the possible deformations to show the boundedness of their fibers. For the special boundedness step, we shall prove the following statement.
Theorem 4.2.
Let $n\in\mathbb{N}$ and $\varepsilon>0$. Let $\mathcal{S}$ be the set of Fano type fibration germs $X\to Z\ni z$ with $\dim X=n$ such that $\widehat{\rm Vol}_{X}(\Delta)\geq\varepsilon$ for some $\mathbb{R}$-divisor $0\leq\Delta\sim_{\mathbb{R},Z}-K_{X}$. Then $\mathcal{S}$ is bounded up to Fano type degenerations (Definition 2.36).
Observe that the Fano type assumption implies that $\Delta\sim_{\mathbb{R},Z}-K_{X}$ is big over $Z$. As $Z$ is affine (by the definition of a fibration germ), the $\mathbb{R}$-divisor $\Delta$ is in fact big, thus its log canonical volume $\widehat{\rm Vol}_{X}(\Delta)$ is well-defined.
Using similar arguments, we also prove the following more general result on log Calabi-Yau fibration germs, which is in turn a variant of [HLQ-vol-ACC]*Conjecture 8.9 and [HX-CY-bdd]*Theorem 1.3.
Theorem 4.3.
Let $n\in\mathbb{N}$ and $\varepsilon,\delta>0$. Let $\mathcal{S}$ be the set of log Calabi-Yau fibration germs $(X,\Delta)\to Z\ni z$ such that
  1. (1)
    $\dim X=n$, and all the nonzero coefficients of $\Delta$ are at least $\delta$,
  2. (2)
    $X$ is of Fano type over $Z$, and
  3. (3)
    $\widehat{\rm Vol}_{X}(\Delta)\geq\varepsilon$ (Definition 3.9).
Then $\mathcal{S}$ is log specially bounded (Definition 2.36).
Section 4.24.5 are devoted to the proof of Theorems 4.2 and 4.3. Before we delve into this somewhat lengthy proof, we first explain the deformation step in the proof of Theorem 4.1 (assuming Theorem 4.2). This is done through the following auxiliary result.
Lemma 4.4.
Let $\mathcal{S}$ be a set of Fano type fibration germs that is bounded up to Fano type degenerations. Then the set
$\mathcal{T}=\{f^{-1}(z)\,|\,f\colon X\to Z\ni z\mbox{ belongs to }\mathcal{S}\}$
is bounded.
Proof.
By assumption, there exists a bounded family $\mathcal{S}_{0}$ of Fano type fibrations such that any fibration germ $f\colon X\to Z\ni z$ in $\mathcal{S}$ degenerates to some $f_{0}\colon X_{0}\to Z_{0}\ni z_{0}$ in $\mathcal{S}_{0}$ by a Fano type test configuration $\widetilde{f}\colon\mathcal{X}\to\mathcal{Z}$. Since $\mathcal{S}_{0}$ is bounded, there exist some positive integers $r$ and $N$ depending only on $\mathcal{S}_{0}$, some very ample line bundle $L_{0}$ on $X_{0}$, and a surjective homomorphism
$\mathcal{E}_{0}:=\mathcal{O}_{Z_{0}}^{\oplus N}\to f_{0*}\mathcal{O}_{X_{0}}(L_{0}),$
such that the induced map
$\mu\colon\mathrm{Sym}(\mathcal{E}_{0})\to\mathcal{R}_{0}:=\oplus_{m\in\mathbb{N}}f_{0*}\mathcal{O}_{X_{0}}(mL_{0})$
of graded $\mathcal{O}_{Z_{0}}$-algebras is also surjective, and its kernel $\mathcal{G}$ is generated by homogeneous elements of degree at most $r$. In particular, the linear system $|L_{0}|$ induces a projectively normal embedding $X_{0}\hookrightarrow Z_{0}\times\mathbb{P}^{N-1}$ and $X_{0}$ is scheme-theoretically cut out by equations of degree at most $r$ in $Z_{0}\times\mathbb{P}^{N}$.
Since $\mathcal{X}$ is of Fano type over $\mathcal{Z}$, we have $R^{1}\widetilde{f}_{*}\mathcal{O}_{\mathcal{X}}=R^{2}\widetilde{f}_{*}\mathcal{O}_{\mathcal{X}}=0$ by Kawamata-Viehweg vanishing, thus from the long exact sequence associated to the exponential exact sequence $0\to\mathbb{Z}\to\mathcal{O}_{\mathcal{X}}\xrightarrow{\mathrm{exp}}\mathcal{O}^{*}_{\mathcal{X}}\to 0$ we see that over some sufficiently small analytic neighborhood $U\subseteq\mathcal{Z}$ of $z_{0}$ we have
$\mathrm{Pic}(\widetilde{f}^{-1}(U))\cong H^{2}(\widetilde{f}^{-1}(U),\mathbb{Z})\cong H^{2}(\widetilde{f}^{-1}(z_{0}),\mathbb{Z})$
where the last isomorphism comes from a deformation retraction of $\widetilde{f}^{-1}(U)$ to $\widetilde{f}^{-1}(z_{0})$. Similarly, as $\mathcal{X}_{0}$ is of Fano type over $\mathcal{Z}_{0}$, there exists some sufficiently small analytic neighborhood $V\subseteq Z_{0}$ of $z_{0}$ such that
$\mathrm{Pic}(f_{0}^{-1}(V))\cong H^{2}(f_{0}^{-1}(z_{0}),\mathbb{Z})=H^{2}(\widetilde{f}^{-1}(z_{0}),\mathbb{Z}).$
It follows that $L_{0}$ extends to a line bundle $\mathcal{L}$ over some small analytic neighborhood $U$ of $z_{0}\in\mathcal{Z}$. Possibly shrinking $U$, we may assume that $\mathcal{L}$ is ample over $U$.
By Kawamata-Viehweg vanishing, we have $R^{1}\widetilde{f}_{*}\mathcal{L}^{m}=0$ for all $m\in\mathbb{N}$. Thus the long exact sequence associated to $0\to\mathcal{L}^{m}\stackrel{{\scriptstyle t}}{{\to}}\mathcal{L}^{m}\to L_{0}^{m}\to 0$ (where $t$ is the linear coordinate on $\mathbb{A}^{1}$) gives
(4.1) $0\to\widetilde{f}_{*}\mathcal{L}^{m}|_{U}\stackrel{{\scriptstyle t}}{{\to}}\widetilde{f}_{*}\mathcal{L}^{m}|_{U}\to f_{0*}\mathcal{O}_{X_{0}}(mL_{0})|_{U}\to 0$
for all $m\in\mathbb{N}$. Since $\mathcal{L}$ is ample over $U$, the $\mathcal{O}_{U}$-algebra $\mathcal{R}:=\oplus_{m\in\mathbb{N}}\widetilde{f}_{*}\mathcal{L}^{m}|_{U}$ is finitely generated. Recall that $\mathcal{R}_{0}$ is generated by at most $N$ homogeneous elements of degree $1$ by our choice of $L_{0}$. Combined with (4.1) and Nakayama’s lemma we deduce that $\mathcal{R}$ is also generated (as an $\mathcal{O}_{U}$-algebra) by at most $N$ homogeneous elements of degree $1$. More precisely, on the open set $U$ by (4.1) we may lift the homomorphism $\mathcal{E}_{0}|_{U}\to f_{0*}\mathcal{O}_{X_{0}}(L_{0})|_{U}$ to a homomorphism
$\mathcal{E}:=\mathcal{O}_{U}^{\oplus N}\to\widetilde{f}_{*}\mathcal{L}|_{U},$
and for any positive integer $m$, after possibly shrinking $U$, the induced homomorphism
$\widetilde{\mu}\colon\mathrm{Sym}(\mathcal{E})\to\mathcal{R}$
is surjective in degree $\leq m$ by Nakayama’s lemma. Since $\mathcal{R}$ is generated by finitely many homogeneous elements, we see that $\widetilde{\mu}$ is surjective after shrinking $U$. In particular, we have a closed embedding $\widetilde{f}^{-1}(U)\hookrightarrow U\times\mathbb{P}^{N-1}$. Let $\mathcal{F}=\ker(\widetilde{\mu})$ and recall $\mathcal{G}=\ker(\mu)$. By construction, $\mu$ is the reduction of $\widetilde{\mu}$ modulo $t$. Note that $\mathcal{R}$ is torsion free and hence flat over $\mathbb{A}^{1}$. It follows that the exact sequence $0\to\mathcal{F}\to\mathrm{Sym}(\mathcal{E})\to\mathcal{R}\to 0$ remains exact modulo $t$ and hence $\mathcal{G}=\mathcal{F}/t\mathcal{F}$. Therefore, as $\mathcal{G}$ is generated by homogeneous elements of degree at most $r$, the same holds for $\mathcal{F}$ by another application of Nakayama’s lemma as before. Thus $\widetilde{f}^{-1}(U)$ is cut out by equations of degree at most $r$ in $U\times\mathbb{P}^{N-1}$ and hence every fiber $\widetilde{f}^{-1}(u)$, where $u\in U$, is cut out by equations of degree at most $r$ in $\mathbb{P}^{N-1}$.
Since $\widetilde{f}\colon\mathcal{X}\to\mathcal{Z}$ is a test configuration of $f\colon X\to Z\ni z$, for any open neighborhood $U$ of $z_{0}\in\mathcal{Z}$ there exists some $u\in U$ such that $\widetilde{f}^{-1}(u)$ is isomorphic to $f^{-1}(z)$. Therefore, every member of $\mathcal{T}$ can be cut out by equations of degree at most $r$ in $\mathbb{P}^{N-1}$ and hence $\mathcal{T}$ is bounded. ∎
Proof of Theorem 4.1 assuming Theorem 4.2.
By Lemma 2.18, there exists some big $\mathbb{R}$-divisor $\Delta_{0}\geq 0$ on $X$ such that $(X,\Delta_{0})$ is klt and every $(K_{X}+\Delta)$-MMP over $T$ is also a $(K_{X}+\Delta_{0})$-MMP. Replacing $\Delta$ by $\Delta_{0}$, we see that it suffices to prove the theorem when $\Delta$ is big.
Suppose that $(X,\Delta)\dashrightarrow(Y,\Delta_{Y})$ is a sequence of steps of a $(K_{X}+\Delta)$-MMP and $f\colon(Y,\Delta_{Y})\to Z$ is an extremal contraction over $T$. By Lemma 3.11, we have $\widehat{\rm Vol}_{Y}(\Delta_{Y})\geq\widehat{\rm Vol}_{X}(\Delta)>0$. Since $-(K_{Y}+\Delta_{Y})$ is ample over $Z$, for any closed point $z\in Z$ we can choose some $0\leq H\sim_{\mathbb{R},Z}-(K_{Y}+\Delta_{Y})$ such that $(Y,\Delta_{Y}+H)\to Z\ni z$ is a log Calabi-Yau fibration germ (implicitly, here we replace $Z$ by an affine neighborhood of $z$ in order to talk about the fibration germ). By Lemma 3.15, we also have $\widehat{\rm Vol}_{Y}(\Delta_{Y}+H)\geq\widehat{\rm Vol}_{X}(\Delta)$. By Theorem 4.2, the Fano type fibration germ $Y\to Z\ni z$ belongs to a set depending only on $\dim X$ and $\widehat{\rm Vol}_{X}(\Delta)$ that is bounded up to Fano type degenerations. By Lemma 4.4, we conclude that the fiber $f^{-1}(z)$ of the extremal contraction belongs to a bounded family depending only on $(X,\Delta)$.
Next suppose that $(Y,\Delta_{Y})\to Z$ is a small contraction and $g\colon(Y^{\prime},\Delta^{\prime})\to Z$ is the flip. Let $H^{\prime}$ be the strict transform of $H$ on $Y^{\prime}$. Then by Lemma 3.11 and the above discussion we have $\widehat{\rm Vol}_{Y^{\prime}}(\Delta^{\prime}+H^{\prime})\geq\widehat{\rm Vol}_{Y}(\Delta_{Y}+H)\geq\widehat{\rm Vol}_{X}(\Delta)$. Thus the same argument above applies to the fibration $Y^{\prime}\to Z\ni z$ and we conclude that the fiber $g^{-1}(z)$ of the flip also belongs to a bounded family depending only on $(X,\Delta)$. ∎
4.2. Relative cone construction
We now turn to the proofs of Theorems 4.2 and 4.3. We first consider a special boundedness result for log Fano fibrations with a fixed finite set of rational coefficients.
Theorem 4.5.
Let $n\in\mathbb{N}$, $\varepsilon>0$, and let $I\subseteq[0,1]\cap\mathbb{Q}$ be a finite set. Let $\mathcal{S}$ be the set of log Fano fibration germs $(X,\Delta)\to Z\ni z$ with $\dim X=n$, $\mathrm{Coeff}(\Delta)\subseteq I$ such that $\widehat{\rm Vol}_{X,\Delta}(H)\geq\varepsilon$ for some $\mathbb{R}$-divisor $H\sim_{\mathbb{R},Z}-(K_{X}+\Delta)$. Then $\mathcal{S}$ is specially bounded (Definition 2.36).
When $X=Z$, this essentially follows from (by Lemma 3.12) the special boundedness of klt singularities proved in [XZ-local-bdd]. Combined with the argument from the previous subsection, Theorem 4.5 already implies the rational coefficient case of Theorem 4.1 for fibers of the extremal contractions.
The main idea of the proof is to consider the relative cone of the log Fano fibration and reduce to [XZ-local-bdd] after analyzing the local volume of the cone singularity. Before we do so, we make a digression to discuss some generality of the relative cone construction. Many results in this subsection are parallel to those of [Kol13]*Section 3.1, which deals with cones over projective pairs.
We start with some general setup and introduce some notation that will be fixed throughout this subsection. Let $f\colon X\to Z$ be a fibration and let $L$ be an $f$-ample $\mathbb{Q}$-Cartier Weil divisor on $X$. Let
$\mathcal{R}_{m}:=f_{*}\mathcal{O}_{X}(mL)\quad\mbox{and}\quad\mathcal{R}:=\oplus_{m\in\mathbb{N}}\mathcal{R}_{m}.$
We are interested in the relative (orbifold) cone $Y:=\mathrm{Spec}_{Z}(\mathcal{R})$. Next, let
$\widetilde{Y}:=\mathrm{Spec}_{X}\left(\oplus_{m\in\mathbb{N}}\mathcal{O}_{X}(mL)\right)$
and let $\pi\colon\widetilde{Y}\to X$ be the induced morphism. It has a section $E\subseteq\widetilde{Y}$ defined by the ideal $\oplus_{m\in\mathbb{N}^{+}}\mathcal{O}_{X}(mL)$. The morphism $\widetilde{Y}\setminus E\to X$ is a Seifert $\mathbb{G}_{m}$-bundle (in the sense of [Kol-Seifert-bundle]) and $E$ can be identified with the zero section. In particular, the divisor $E$ is $\mathbb{Q}$-Cartier and $E\cong X$. Locally on $Z$ we have $\mathcal{R}=H^{0}(\mathcal{O}_{\widetilde{Y}})$ (in other words $(f\circ\pi)_{*}\mathcal{O}_{\widetilde{Y}}=\mathcal{R}$). Hence we have an induced morphism $\varphi\colon\widetilde{Y}\to Y$ over $Z$ which is an isomorphism away from the zero section $E$ (by the ampleness of $L$ over $Z$), and $\varphi|_{E}=f$. Let $\Gamma:=\varphi(E)\cong Z$. It gives a section $\sigma\colon Z\to Y$ of the projection $g\colon Y\to Z$. The situation is summarized by the following diagram.
[…]
Note that if $f\colon X\to Z$ is birational, then $\varphi$ is small and induces an isomorphism $\varphi_{*}\colon\mathrm{Cl}(\widetilde{Y})\cong\mathrm{Cl}(Y)$. Otherwise $E$ is the only exceptional divisor of $\varphi$ and we get an isomorphism $\varphi_{*}\colon\mathrm{Cl}(\widetilde{Y})/[E]\cong\mathrm{Cl}(Y)$. We also have a natural pullback map $\pi^{*}\colon\mathrm{Cl}(X)\to\mathrm{Cl}(\widetilde{Y})$ since $\pi$ is equidimensional.
Lemma 4.6.
The induced map $\pi^{*}\colon\mathrm{Cl}(X)\to\mathrm{Cl}(\widetilde{Y})$ is an isomorphism. Moreover, we have $-E\sim\pi^{*}L$ and $K_{\widetilde{Y}}+E\sim\pi^{*}K_{X}$.
Proof.
Over the smooth locus $U$ of $X$ the divisor $L$ is Cartier, and $\pi^{-1}(U)$ is an $\mathbb{A}^{1}$-bundle over $U$, hence $\pi^{*}$ induces an isomorphism $\mathrm{Cl}(\pi^{-1}(U))\cong\mathrm{Cl}(U)$. Since $\pi$ is equidimensional, the complement of $\pi^{-1}U$ in $\widetilde{Y}$ has codimension at least $2$, thus $\mathrm{Cl}(X)\cong\mathrm{Cl}(U)$ and $\mathrm{Cl}(\widetilde{Y})\cong\mathrm{Cl}(\pi^{-1}(U))$. It follows that $\pi^{*}\colon\mathrm{Cl}(X)\to\mathrm{Cl}(\widetilde{Y})$ is also an isomorphism.
Any rational section of $\mathcal{O}_{X}(L)\subseteq\pi_{*}\mathcal{O}_{\widetilde{Y}}$ defines a rational function on $\widetilde{Y}$ and induces the linear equivalence relation $\pi^{*}L+E\sim 0$. If $z$ is the affine coordinate of the fiber of $\pi^{-1}(U)\to U$, then $\frac{dz}{z}$ is a nowhere zero section of $K_{\pi^{-1}(U)/U}+E$. This gives $K_{\pi^{-1}(U)/U}+E\sim 0$ and hence $K_{\widetilde{Y}}+E\sim\pi^{*}K_{X}$ as $\pi^{-1}(U)$ (resp. $U$) is a big open subset of $\widetilde{Y}$ (resp. $X$). ∎
For the next several lemmas, We say that the fibration $f\colon X\to Z$ is of fiber type if $\dim X>\dim Z$.
Lemma 4.7.
Let $G$ be a Weil divisor on $X$. Then $\varphi_{*}\pi^{*}G$ is Cartier if and only if either $f$ is birational and $G\sim_{Z}0$, or $f$ is of fiber type and $G\sim_{Z}mL$ for some $m\in\mathbb{Z}$. Moreover, $\pi^{*}G\sim_{Y}0$ if and only if $G\sim_{Z}0$.
Proof.
First consider the case when $f$ is birational. If $G=f^{*}G_{0}$ for some Cartier divisor $G_{0}$ on $Z$, then $\pi^{*}G\sim_{Y}0$ and $\varphi_{*}\pi^{*}G=g^{*}G_{0}$ is also Cartier. Conversely, if $G^{\prime}:=\varphi_{*}\pi^{*}G$ is Cartier then $\pi^{*}G=\varphi^{*}G^{\prime}$ as $\varphi$ is small. Restricting to the zero section $E\cong X$ gives $G=f^{*}\sigma^{*}G^{\prime}$, hence $G\sim_{Z}0$.
Next assume that $f$ is of fiber type. The proof is similar. If $G-mL=f^{*}G_{0}$ for some Cartier divisor $G_{0}$, then by Lemma 4.6 we have $\varphi_{*}\pi^{*}G=\varphi_{*}(\pi^{*}G+mE)\sim\varphi_{*}\pi^{*}(G-mL)=g^{*}G_{0}$, hence $\varphi_{*}\pi^{*}G$ is also Cartier. Conversely, if $G^{\prime}:=\varphi_{*}\pi^{*}G$ is Cartier then $\pi^{*}G+mE=\varphi^{*}G^{\prime}$ for some integer $m$. If in addition $\pi^{*}G\sim_{Y}0$ then $m=0$. Restricting to the zero section $E\cong X$ and noting that $-E|_{E}\cong L$ by Lemma 4.6, we obtain $G-mL\sim f^{*}\sigma^{*}G^{\prime}$, hence $G\sim_{Z}mL$. ∎
Next let $\Delta\geq 0$ be an $\mathbb{R}$-divisor on $X$ and consider the $\mathbb{R}$-divisor $D:=\varphi_{*}\pi^{*}\Delta\geq 0$ on $Y$. Geometrically, it is the relative cone over $\Delta$.
Lemma 4.8.
Assume that $f$ is birational and $0<\lambda\leq 1$ (resp. $0\leq\lambda\leq 1$). Then the pair $(Y,D+(1-\lambda)\Gamma)$ is klt (resp. log canonical) if and only if $K_{X}+\Delta+\lambda L\sim_{\mathbb{R},Z}0$ and $(X,\Delta)$ is klt (resp. log canonical).
Proof.
We only treat the klt case since the log canonical version is similar. By Lemma 4.6, we have
$K_{Y}+D+(1-\lambda)\Gamma\sim_{\mathbb{R}}\varphi_{*}(K_{\widetilde{Y}}+\pi^{*}\Delta+(1-\lambda)E)\sim_{\mathbb{R}}\varphi_{*}\pi^{*}(K_{X}+\Delta+\lambda L).$
Thus by Lemma 4.7, we see that $K_{Y}+D+(1-\lambda)\Gamma$ is $\mathbb{R}$-Cartier if and only if $K_{X}+\Delta+\lambda L\sim_{\mathbb{R},Z}0$, in which case $(\widetilde{Y},\pi^{*}\Delta+(1-\lambda)E)$ is the crepant pullback of $(Y,D+(1-\lambda)\Gamma)$.
By [Kol-Seifert-bundle]*Paragraphs 40-42, if $(Y,D+(1-\lambda)\Gamma)$ is klt, then so is $(X,\Delta)$ since the latter is the orbifold quotient of $(Y\setminus\Gamma,D|_{Y\setminus\Gamma})$ by the $\mathbb{G}_{m}$-action. Conversely, we show that $(Y,D+(1-\lambda)\Gamma)$ is klt if $(X,\Delta)$ is klt. To see this, it suffices to show that the crepant pullback $(\widetilde{Y},\pi^{*}\Delta+(1-\lambda)E)$ is klt. Since $(K_{\widetilde{Y}}+\pi^{*}\Delta+E)|_{E}\cong K_{X}+\Delta$ (where we identify $E$ with $X$) and $(X,\Delta)$ is klt, by inversion of adjunction we see that $(\widetilde{Y},\pi^{*}\Delta+E)$ is plt in a neighborhood of $E$. Since the natural $\mathbb{G}_{m}$-action on $\widetilde{Y}$ preserves the pair $(\widetilde{Y},\pi^{*}\Delta+E)$, every non-klt center intersects the $\mathbb{G}_{m}$-fixed locus $E$; this implies that $(\widetilde{Y},\pi^{*}\Delta+E)$ is plt everywhere. As $0<\lambda\leq 1$, we see that $(\widetilde{Y},\pi^{*}\Delta+(1-\lambda)E)$ is klt. ∎
By essentially the same argument, we also have the following.
Lemma 4.9.
Assume that $f$ is of fiber type. Then the pair $(Y,D)$ is klt (resp. log canonical) if and only if $K_{X}+\Delta+\lambda L\sim_{\mathbb{Q},Z}0$ for some $\lambda>0$ (resp. $\lambda\geq 0$) and $(X,\Delta)$ is klt (resp. log canonical). In such case we have $A_{Y,D}(E)=\lambda$.
Proof.
Since $K_{Y}+D=\varphi_{*}\pi^{*}(K_{X}+\Delta)$, the first part follows from Lemma 4.7 as in the proof of Lemma 4.9. It remains to show that $K_{X}+\Delta+\lambda L\sim_{\mathbb{R},Z}0$ where $\lambda:=A_{Y,D}(E)$. But this again follows from Lemma 4.7 as $\pi^{*}(K_{X}+\Delta+\lambda L)\sim_{\mathbb{R}}K_{\widetilde{Y}}+\pi^{*}\Delta+(1-\lambda)E=\varphi^{*}(K_{Y}+D)\sim_{\mathbb{R},Y}0$. ∎
4.3. Volume estimate
We next apply the above cone construction to log Fano fibrations and prove the key local volume estimate towards the proof of Theorem 4.5.
Let $f\colon(X,\Delta)\to Z\ni z$ be a log Fano fibration germ. We assume that the coefficients of $\Delta$ are rational. Let $r$ be a positive integer such that $r\Delta$ has integer coefficients, let
$\mathcal{R}=\oplus_{m\in\mathbb{N}}H^{0}(X,-mr(K_{X}+\Delta)),$
and let $Y=\mathrm{Spec}(\mathcal{R})$ be the relative cone with respect to the Weil divisor $L:=-r(K_{X}+\Delta)$. We keep the notation from the diagram (4.2). Let
(4.3) $D=\begin{cases}\varphi_{*}\pi^{*}\Delta+\left(1-\frac{1}{r}\right)\Gamma,&\quad\mbox{if }f\mbox{ is birational},\\ \varphi_{*}\pi^{*}\Delta,&\quad\mbox{if }f\mbox{ is of fiber type}.\end{cases}$
Note that $\Gamma$ is a divisor on $Y$ if and only if $f$ is birational. By Lemmas 4.8 and 4.9, the pair $(Y,D)$ is klt. Let $y=\sigma(z)\in Y$. The main result of this subsection is the following.
Theorem 4.10.
Under the above setup, we have
$\widehat{\rm vol}(y,Y,D)\geq\frac{\widehat{\rm Vol}_{X,\Delta}(H)}{r}$
for any $\mathbb{R}$-divisor $H\sim_{\mathbb{R},Z}-(K_{X}+\Delta)$.
Proof.
By Lemma 3.16, and since $\Delta$ has rational coefficient, it is enough to prove the statement when $H$ is a $\mathbb{Q}$-divisor. Since the question is local on $Z$, we may also shrink $Z$ around $z$ and assume that $H\sim_{\mathbb{Q}}-(K_{X}+\Delta)$. By definition, it suffices to show that
$\widehat{\rm vol}(y,Y,D)\geq\frac{(1-\varepsilon)^{n}\mathrm{vol}(V_{\bullet})}{r}$
for any $0<\varepsilon\ll 1$ and any admissible graded linear series $V_{\bullet}$ of $H$. Our plan is to find a sequence of subspaces $W_{m}\subseteq\mathcal{O}_{Y,y}$ of increasing dimensions such that
[…]
are reasonably bounded from below and apply Lemma 3.4. Here we use the shorthand notation $\mathrm{lct}_{y}(s)$ for $\mathrm{lct}_{y}(Y,D;\{s=0\})$ where $0\neq s\in\mathcal{O}_{Y,y}$. In the simplest case, if $-r(K_{X}+\Delta)$ is Cartier and $H=-(K_{X}+\Delta)$, then each $V_{mr}$ ($m\in\mathbb{N}$) is already a subspace of $\mathcal{R}_{m}\subseteq\mathcal{O}_{Y,y}$, and we can take $W_{m}$ to be the span of $V_{r},V_{2r},\cdots,V_{mr}$.
In general, let $k$ be a sufficiently divisible positive integer such that $kH$ and $k(K_{X}+\Delta)$ are both Cartier, $kH\sim-k(K_{X}+\Delta)$, and $V_{km}\neq 0$ for all $m\in\mathbb{N}$. Let $V^{\prime}_{\bullet}$ be the graded linear series of $kH$ given by $V^{\prime}_{m}=V_{km}$. Let $\alpha=\frac{1-\varepsilon}{k}$. Since $H\sim_{\mathbb{Q}}-(K_{X}+\Delta)$ is ample over $Z$, so is
$-mr(K_{X}+\Delta)-\lfloor\alpha mr\rfloor\cdot kH\sim-(mr-\lfloor\alpha mr\rfloor\cdot k)(K_{X}+\Delta)$
for any $m>0$; moreover, for any sufficiently positive integer $m$ (say, when $m\geq m_{0}$) we have $mr-\lfloor\alpha mr\rfloor\cdot k\geq\varepsilon mr\gg 0$ and thus by Lemma 2.9 we can find some
$u_{m}\in H^{0}(X,-mr(K_{X}+\Delta)-\lfloor\alpha mr\rfloor\cdot kH)$
such that $(X,\Delta+\frac{1}{\varepsilon mr}\{u_{m}=0\})$ is log canonical.
Multiplication by $u_{m}$ takes $V^{\prime}_{\lfloor\alpha mr\rfloor}$ into $\mathcal{R}_{m}=H^{0}(X,-mr(K_{X}+\Delta))\subseteq\mathcal{O}_{Y,y}$ and we take $W_{m}\subseteq\mathcal{O}_{Y,y}$ to be the span of $u_{j}\cdot V^{\prime}_{\lfloor\alpha jr\rfloor}$ for $j=m_{0},\dots,m$. We claim that
(4.4) $m\cdot\mathrm{lct}_{y}(s)\geq\frac{1}{r}$
for all $0\neq s\in W_{m}$. By the following Lemma 4.11, since $W_{m}$ is invariant under the $\mathbb{G}_{m}$-action on $Y$ and has a weight decomposition $W_{m}=\oplus_{j=m_{0}}^{m}u_{j}\cdot V^{\prime}_{\lfloor\alpha jr\rfloor}$, it suffices to prove the claim when $s=u_{j}\cdot w$ for some $m_{0}\leq j\leq m$ and some $w\in V^{\prime}_{\lfloor\alpha jr\rfloor}$.
Since $V_{\bullet}$ is admissible, we know that the pair $(X,\Delta+\frac{1}{\alpha jkr}\{w=0\})=(X,\Delta+\frac{1}{(1-\varepsilon)jr}\{w=0\})$ is log canonical. By construction $(X,\Delta+\frac{1}{\varepsilon jr}\{u_{j}=0\})$ is also log canonical. By taking convex combination, we deduce that
$\left(X,\Delta+(1-\varepsilon)\cdot\frac{1}{(1-\varepsilon)jr}\{w=0\}+\varepsilon\cdot\frac{1}{\varepsilon jr}\{u_{j}=0\}\right)$
is log canonical. In other words, the pair
$(X,\Delta+\Delta_{0}):=\left(X,\Delta+\frac{1}{jr}\{u_{j}w=0\}\right)$
is log canonical. On the other hand, we have $K_{X}+\Delta+\Delta_{0}\sim_{\mathbb{Q},Z}0$ as $u_{j}w\in H^{0}(X,-jr(K_{X}+\Delta))$. If $D_{0}:=\frac{1}{jr}\{s=0\}$ is the corresponding divisor on $Y$, then we also have
$D_{0}=\begin{cases}\varphi_{*}\pi^{*}\Delta_{0}+\frac{1}{r}\Gamma&\quad\mbox{if }f\mbox{ is birational},\\ \varphi_{*}\pi^{*}\Delta_{0}&\quad\mbox{if }f\mbox{ is of fiber type}.\end{cases}$
By Lemma 4.8 and 4.9, these imply that $(Y,D+D_{0})$ is log canonical at $y$, which gives $\mathrm{lct}_{y}(s)\geq\frac{1}{jr}\geq\frac{1}{mr}$ and hence the claim (4.4).
By Lemma 2.8 we have
$\dim W_{m}=\sum_{j=m_{0}}^{m}\dim V^{\prime}_{\lfloor\alpha jr\rfloor}=\sum_{j=m_{0}}^{m}\dim V_{\lfloor\alpha jr\rfloor\cdot k}=(\alpha kr)^{n}\mathrm{vol}(V_{\bullet})\frac{m^{n+1}}{(n+1)!}+o(m^{n+1}).$
Combined with (4.4) and Lemma 3.4 we see that
$\widehat{\rm vol}(y,Y,D)\geq\frac{c_{1}}{r^{n+1}}$
for any $c_{1}<(\alpha kr)^{n}\mathrm{vol}(V_{\bullet})$. It follows that
$\widehat{\rm vol}(y,Y,D)\geq\frac{(\alpha k)^{n}\cdot\mathrm{vol}(V_{\bullet})}{r}=\frac{(1-\varepsilon)^{n}\mathrm{vol}(V_{\bullet})}{r}.$
This concludes the proof. ∎
The following lemma is used in the above proof. We keep the same notation $\mathrm{lct}_{y}(s)$ for $\mathrm{lct}_{y}(Y,D;\{s=0\})$
Lemma 4.11.
Let $y\in(Y,D)$ be a klt singularity with a torus $\mathbb{T}$-action and let $V\subseteq\mathcal{O}_{Y,y}$ be a finite dimensional $\mathbb{T}$-invariant subspace. Assume that for some constant $c\geq 0$ we have $\mathrm{lct}_{y}(s)\geq c$ for all nonzero $\mathbb{T}$-eigenvectors $s\in V$. Then $\mathrm{lct}_{y}(s)\geq c$ for all nonzero $s\in V$.
Proof.
For any $0\neq s\in V$, there exists some one-parameter subgroup $\rho\colon\mathbb{G}_{m}\to\mathbb{T}$ such that the flat limit $D_{0}$ of $\rho(t)\cdot D$, where $D=\{s=0\}$ and when $t\to 0$, is a $\mathbb{T}$-invariant divisor. Indeed, if $s=\sum_{i=1}^{m}s_{i}$ is the weight decomposition where $s_{i}$ has weight $\alpha_{i}\in\mathrm{Hom}(\mathbb{T},\mathbb{G}_{m})$, and choose $\rho$ such that $\langle\rho,\alpha_{1}\rangle<\cdots<\langle\rho,\alpha_{m}\rangle$, then the flat limit is $\{s_{1}=0\}$. By assumption, we have $\mathrm{lct}_{y}(D_{0})\geq c$, hence by the lower semi-continuity of log canonical threshold [Amb-lct-lsc]*Corollary 1.10, we get $\mathrm{lct}_{y}(D)\geq c$ as well. ∎
By [XZ-local-bdd] and the stable degeneration theorem (see Theorem 2.27), we know that klt singularities whose local volumes are bounded away from zero are bounded up to special degeneration. Combined with Theorem 4.10, we are now ready to prove Theorem 4.5.
Proof of Theorem 4.5.
Choose some positive integer $r$ such that $I\subseteq\frac{1}{r}\mathbb{N}$. Let $f\colon(X,\Delta)\to Z\ni z$ be a log Fano fibration germ in $\mathcal{S}$. Let $y\in(Y,D)$ be the relative cone over $(X,\Delta)$ with respect to the ample Weil divisor $L=-r(K_{X}+\Delta)$, where the boundary $D$ is defined as in (4.3). By Theorem 4.10 and our assumption, we have $\widehat{\rm vol}(y,Y,D)\geq\frac{\varepsilon}{r}$. Note that $y\in(Y,D)$ carries a $\mathbb{G}_{m}$-action by the cone construction.
By Theorem 2.27, there exists an $\mathbb{R}$-Gorenstein family $\mathcal{B}\subseteq(\mathcal{Y},\mathcal{D})\to\mathcal{B}$ of klt singularities, depending only on $n,I$ and $\varepsilon$, with the action of an algebraic torus $\mathbb{T}$ over $\mathcal{B}$, such that $y\in(Y,D)$ admits a $\mathbb{G}_{m}$-equivariant special degeneration to $b\in(\mathcal{Y}_{b},\mathcal{D}_{b})$ for some $b\in\mathcal{B}$, and by specializing the $\mathbb{G}_{m}$-action to $b\in(\mathcal{Y}_{b},\mathcal{D}_{b})$, we get a one-parameter subgroup of $\mathbb{T}_{b}$.
By base change, we shall assume that on each of the connected components $\mathcal{B}_{1},\dots,\mathcal{B}_{k}$ of $\mathcal{B}$, we have $\mathbb{T}=\mathbb{T}_{i}\times\mathcal{B}$ for some torus $\mathbb{T}_{i}$. By Lemma 2.23, the weight monoid $S_{i}$ is constant over each $\mathcal{B}_{i}$. We claim that for each $i\in\{1,\dots,k\}$, there are only finitely many one-parameter subgroups of $\mathbb{T}_{i}$ that are possibly induced by the above construction. To see this, suppose that $b\in\mathcal{B}_{i}$ (so we may identify $\mathbb{T}_{b}$ with $\mathbb{T}_{i}$) and let $E$ be the divisor over $Y$ as in the diagram (4.2); that is, the zero section of the corresponding Seifert $\mathbb{G}_{m}$-bundle over $Z$. Then we have $\mathrm{wt}_{\rho}=\mathrm{ord}_{E}$ as valuations on $Y$, where we denote by $\rho\colon\mathbb{G}_{m}\to\mathrm{Aut}(Y,D)$ the one-parameter subgroup coming from the cone construction. We can identify the induced one-parameter subgroup of $\mathrm{Aut}(\mathcal{Y}_{b},\mathcal{D}_{b})$ with some primitive element of $S_{i}^{\vee}$, which we still denote by $\rho$. By (4.3) and Lemma 4.9, we know that $A_{Y,D}(E)=\frac{1}{r}$, thus by Lemma 2.24 we deduce that $A_{\mathcal{Y},\mathcal{D}}(\mathrm{wt}_{\rho})=A_{\mathcal{Y}_{b},\mathcal{D}_{b}}(\mathrm{wt}_{\rho})=\frac{1}{r}$. Since the log discrepancy function $\xi\mapsto A_{\mathcal{Y},\mathcal{D}}(\mathrm{wt}_{\xi})$ is positive (as $(\mathcal{Y},\mathcal{D})$ is klt) and linear on $S_{i}^{\vee}$ (by Lemma 2.21), we see that there are at most finitely many $\rho\in S_{i}^{\vee}$ such that $A_{\mathcal{Y},\mathcal{D}}(\mathrm{wt}_{\rho})=\frac{1}{r}$. This proves the claim.
Suppose that $y\in(Y,D)$ specially and $\mathbb{G}_{m}$-equivariantly degenerates to $b\in(\mathcal{Y}_{b},\mathcal{D}_{b})$. By the following Lemma 4.12, it induces a special degeneration of $(X,\Delta)\to Z\ni z$ to the log Fano fibration germ $(\mathcal{Y}_{b}\setminus\mathcal{Y}_{b}^{\mathbb{G}_{m}},\mathcal{D}_{b}\setminus\mathcal{D}_{b}^{\mathbb{G}_{m}})/\mathbb{G}_{m}\to\mathcal{Y}_{b}/\!\!/\mathbb{G}_{m}\ni z_{b}$, where $z_{b}$ is the image of $b\in\mathcal{Y}_{b}$ in the GIT quotient. Thus it remains to show that latter form an $\mathbb{R}$-Gorenstein family of log Fano fibrations over a finite type base. This is clear since there are only finitely many one-parameter subgroups of $\mathbb{T}_{i}$ to consider and $\mathcal{B}$ is of finite type. ∎
Lemma 4.12.
Let $(X,\Delta)\to Z\ni z$ be a log Fano fibration germ and let $y\in(Y,D)$ be the relative cone with respect to $L=-r(K_{X}+\Delta)$ as above for some $r\in\mathbb{N}$ such that $r\Delta$ has integer coefficients. Let $(\mathcal{Y}=\mathrm{Spec}(\mathcal{R}),\Delta_{\mathcal{Y}})\to\mathbb{A}^{1}$ be a $\mathbb{G}_{m}$-equivariant special test configuration of the singularity $y\in(Y,D)$. Then it induces a special test configuration
$(\mathcal{X},\mathcal{D})\to\mathcal{Z}$
of $(X,\Delta)\to Z\ni z$ such that $(\mathcal{Y},\Delta_{\mathcal{Y}})$ is the relative cone of $(\mathcal{X},\mathcal{D})$ over $\mathcal{Z}$ with respect to $\mathcal{L}:=-r(K_{\mathcal{X}}+\mathcal{D})$.
Proof.
Since $(\mathcal{Y},\Delta_{\mathcal{Y}})\to\mathbb{A}^{1}$ is a $\mathbb{G}_{m}$-equivariant test configuration, we have a weight decomposition $\mathcal{R}=\oplus_{m\in\mathbb{N}}\mathcal{R}_{m}$ with respect to the fiberwise $\mathbb{G}_{m}$-action. Note that there are no negative weights by Lemma 2.23. Let $\mathcal{X}:=\mathrm{Proj}(\mathcal{R})$ and $\mathcal{Z}:=\mathrm{Spec}(\mathcal{R}_{0})$. Then the natural morphism $\mathcal{X}\to\mathcal{Z}$ over $\mathbb{A}^{1}$ is a test configuration of the fibration $X\to Z$. Since $\mathcal{Y}\to\mathbb{A}^{1}$ is a test configuration of the singularity $y\in Y$, the closure $C$ of $\{y\}\times(\mathbb{A}^{1}\setminus\{0\})$ in $\mathcal{Y}$ is a section of the morphism $\mathcal{Y}\to\mathbb{A}^{1}$. The latter factors through $\mathcal{Z}$, hence the closure of $\{z\}\times(\mathbb{A}^{1}\setminus\{0\})$ in $\mathcal{Z}$ coincides with the image of $C$ and is a section of $\mathcal{Z}\to\mathbb{A}^{1}$. This implies that $\mathcal{Z}\to\mathbb{A}^{1}$ is also a test configuration of the singularity $z\in Z$ and therefore $\mathcal{X}\to\mathcal{Z}$ is a test configuration of the fibration germ $X\to Z\ni z$.
Let $\mathcal{D}$ be the closure of $\Delta\times(\mathbb{A}^{1}\setminus\{0\})$ in $\mathcal{X}$. We next show that the Weil divisor $\mathcal{L}=-r(K_{\mathcal{X}}+\mathcal{D})$ is ample $\mathbb{Q}$-Cartier and $(\mathcal{Y},\Delta_{\mathcal{Y}})$ is the relative orbifold cone of $(\mathcal{X},\mathcal{D})$ over $\mathcal{Z}$ respect to $\mathcal{L}$. Let $\mathcal{W}=\mathcal{Y}^{\mathbb{G}_{m}}$ be the fixed locus of the fiberwise $\mathbb{G}_{m}$-action. Then $\mathcal{Y}\setminus\mathcal{W}\to\mathcal{X}$ is a Seifert $\mathbb{G}_{m}$-bundle. We claim that it has reduced (hence smooth) fibers over all codimension one points of $\mathcal{X}$. In fact, away from $\mathcal{X}_{0}$ the Seifert $\mathbb{G}_{m}$-bundle is given by the Weil divisor $-r(K_{\mathcal{X}}+\mathcal{D})$, hence the codimension one fibers are reduced by [Kol-Seifert-bundle]*Theorem 7. The fiber over the generic point of $\mathcal{X}_{0}$ is also reduced since $(\mathcal{Y},\Delta_{\mathcal{Y}})\to\mathbb{A}^{1}$ is a special test configuration and in particular $\mathcal{Y}_{0}$ is reduced. This proves the claim. By [Kol-Seifert-bundle]*Theorem 7, we deduce that the Seifert $\mathbb{G}_{m}$-bundle $\mathcal{Y}\setminus\mathcal{W}\to\mathcal{X}$ is given by some Weil divisor $\mathcal{L}$ on $\mathcal{X}$. We necessarily have $\mathcal{L}=\mathcal{O}_{\mathcal{X}}(1)$ and thus $\mathcal{L}$ is $\mathbb{Q}$-Cartier and ample. By construction, $\mathcal{L}|_{\mathcal{X}\setminus\mathcal{X}_{0}}\sim-r(K_{\mathcal{X}}+\mathcal{D})|_{\mathcal{X}\setminus\mathcal{X}_{0}}$. Since $\mathcal{Y}_{0}$ is integral, so is $\mathcal{X}_{0}$. As $\mathcal{X}_{0}\sim 0$, this implies that $\mathcal{L}\sim-r(K_{\mathcal{X}}+\mathcal{D})$ and therefore $-r(K_{\mathcal{X}}+\mathcal{D})$ is $\mathbb{Q}$-Cartier and ample. Replacing $\mathcal{L}$ by $-r(K_{\mathcal{X}}+\mathcal{D})$, we conclude that $(\mathcal{Y},\Delta_{\mathcal{Y}})$ is the relative orbifold cone of $(\mathcal{X},\mathcal{D})$ over $\mathcal{Z}$ with respect to $\mathcal{L}=-r(K_{\mathcal{X}}+\mathcal{D})$. Since $(\mathcal{Y},\mathcal{Y}_{0}+\Delta_{\mathcal{Y}})$ is plt by assumption, the base $(\mathcal{X},\mathcal{X}_{0}+\mathcal{D})$ is also plt by [Kol-Seifert-bundle]*Paragraphs 40-42 as in the proof of Lemma 4.8. This implies that $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ is a special test configuration and we finish the proof. ∎
4.4. Finiteness of models
To go from the special boundedness of log Fano fibration germs in Theorem 4.5 to the log special boundedness of Fano type fibration germs in Theorems 4.2 and 4.3, we employ the following proof strategy. For simplicity, suppose we are in the setting of Theorem 4.2. Let $X\to Z\ni z$ be a Fano type fibration germ. Then the ample model $X\dashrightarrow X^{\prime}$ of $-K_{X}$ exists over $Z$. Assume that $\widehat{\rm Vol}_{X}(\Delta)\geq\varepsilon$ for some $\mathbb{R}$-divisor $0\leq\Delta\sim_{\mathbb{R},Z}-K_{X}$ as in Theorem 4.2. From Lemma 3.11, it is not hard to see that $\widehat{\rm Vol}_{X^{\prime}}(\Delta^{\prime})\geq\varepsilon$ where $\Delta^{\prime}\sim_{\mathbb{R},Z}-K_{X^{\prime}}$ is the strict transform of $\Delta$. Since $X^{\prime}\to Z\ni z$ is a log Fano fibration germ, it belongs to a specially bounded set by Theorem 4.5. Next we observe that $X$ can be recovered from $X^{\prime}$ up to finite ambiguity. In fact, all the exceptional divisors of the birational map $X\dashrightarrow X^{\prime}$ have non-positive discrepancies over $X^{\prime}$. Since $X^{\prime}$ is log Fano over $Z$, there are only finitely many birational models $X$ with this property. With this in mind, we prove a family version of such finiteness of models (Theorem 4.13). As a result, if $X^{\prime}\to Z$ belongs to a bounded (resp. specially bounded) set, then so is $X\to Z$ (Theorem 4.22).
Turning to more details, we first prove in this subsection the following statement that roughly says a set of Fano type fibrations is bounded if their ample models are bounded. The stronger version, where we replace boundedness by special boundedness, will be established in the next subsection.
Theorem 4.13.
Let $\mathcal{S}$ be a bounded set of log Fano fibrations (Definition 2.32). Then there exists a bounded set $\mathcal{T}$ of Fano type fibrations such that for any $(X,\Delta)\to Z$ in $\mathcal{S}$ and any birational map $f\colon X^{\prime}\dashrightarrow X$ over $Z$ such that $f^{*}(K_{X}+\Delta)=K_{X^{\prime}}+\Delta^{\prime}$ for some $\Delta^{\prime}\geq 0$, we have $(X^{\prime},\Delta^{\prime})\to Z$ belongs to $\mathcal{T}$.
This can be viewed as a local version of [HX-CY-bdd]*Section 2.1 and we essentially mimic the strategy of loc. cit.. The first step is to prove a local version of [dFH-deform-canonical-pair]*Lemma 6.6. Namely, given a family of Fano type fibrations, the relative Picard group is a constant local system possibly after a base change (Corollary 4.17). For this we need a few auxiliary results. For any fibration $f\colon X\to Y$, we let $N^{1}(X/Y):=\mathbb{Q}\otimes(\mathrm{Pic}(X)/f^{*}\mathrm{Pic}(Y))$, and let $H_{j}(X/Y,\mathbb{Q})\subseteq H_{j}(X,\mathbb{Q})$ be the subspace generated by the images of $H_{j}(X_{y},\mathbb{Q})\to H_{j}(X,\mathbb{Q})$ where $y\in Y$ (cf. [KM-flip-classification]*Definition (12.1.2)). We also let $H_{i}(X/Y,\mathbb{Q})_{\mathrm{alg}}\subseteq H_{i}(X/Y,\mathbb{Q})$ be the subspace generated by the algebraic cycles in the fiber of $f$.
Lemma 4.14.
Let $f\colon X\to Y$ be a fibration. Assume that $R^{1}f_{*}\mathcal{O}_{X}=0$ and $X$ has rational singularities. Then the intersection pairing $N^{1}(X/Y)\times H_{2}(X/Y,\mathbb{Q})\to\mathbb{Q}$ induces an inclusion $\iota\colon N^{1}(X/Y)\hookrightarrow H_{2}(X/Y,\mathbb{Q})_{\mathrm{alg}}^{\vee}$. In particular, $\dim N^{1}(X/Y)<+\infty$. If in addition $X$ is smooth, then $\iota$ is an isomorphism.
Proof.
The exponential exact sequence gives rise to a homomorphism $\mathrm{Pic}(X)\to H^{2}(X,\mathbb{Z})$ and hence a natural pairing $\mathrm{Pic}(X)_{\mathbb{Q}}\times H_{2}(X/Y,\mathbb{Q})\to H^{2}(X,\mathbb{Q})\times H_{2}(X,\mathbb{Q})\to\mathbb{Q}$. It descends to $N^{1}(X/Y)$ as $\langle f^{*}\mathrm{Pic}(Y),H_{2}(X/Y,\mathbb{Q})\rangle=0$ by definition. By [KM-flip-classification]*Proposition (12.1.4), we see that the induced map $\iota$ is an inclusion. Suppose that $X$ is smooth and let $\overline{X}$ be a smooth projective compactification of $X$. If $\iota$ is not surjective, then there exists a nonzero algebraic class $\alpha\in H_{2}(X/Y,\mathbb{Q})_{\mathrm{alg}}$ such that $(L\cdot\alpha)=0$ for all line bundle $L$ on $\overline{X}$. As $\overline{X}$ is smooth, this implies that the image of $\alpha$ in $H_{2}(\overline{X},\mathbb{Q})$ is zero. But by (the dual statement of) [Voi-Hodge-book-II]*Proposition 4.23, the composition $H_{2}(X/Y,\mathbb{Q})\hookrightarrow H_{2}(X,\mathbb{Q})\to H_{2}(\overline{X},\mathbb{Q})$ is injective, a contradiction. ∎
Lemma 4.15.
Let $f\colon X\to Y$ be a fibration over a reduced base $T$. Let $\mathcal{F}$ be a coherent sheaf on $X$ such that $R^{i}f_{*}\mathcal{F}=0$ for some $i\in\mathbb{N}$. Then there exists a dense open subset $U$ of $T$ such that $R^{i}f_{*}\mathcal{F}_{t}=0$ for all $t\in U$.
Proof.
[…]
Lemma 4.16.
Let $\varphi\colon\mathcal{X}\to\mathcal{Z}$ be a fibration over a variety $\mathcal{B}$. Assume that $\mathcal{X}$ has rational singularities and $R^{1}\varphi_{*}\mathcal{O}_{\mathcal{X}}=R^{2}\varphi_{*}\mathcal{O}_{\mathcal{X}}=0$. Then there exists a generically finite dominant morphism $\mathcal{B}^{\prime}\to\mathcal{B}$ such that after base change $\mathcal{X}^{\prime}=\mathcal{X}\times_{\mathcal{B}}\mathcal{B}^{\prime}$, $\mathcal{Z}^{\prime}=\mathcal{Z}\times_{\mathcal{B}}\mathcal{B}^{\prime}$, the restriction map
$N^{1}(\mathcal{X}^{\prime}/\mathcal{Z}^{\prime})\to N^{1}(\mathcal{X}^{\prime}_{b}/\mathcal{Z}^{\prime}_{b})$
is an isomorphism for every $b\in\mathcal{B}^{\prime}$.
Proof.
By Lemma 4.15 and shrinking $\mathcal{B}$ if necessary, we shall assume that $\mathcal{B}$ is smooth, $\mathcal{X}_{b}$ has rational singularities, and $R^{1}\varphi_{*}\mathcal{O}_{\mathcal{X}_{b}}=R^{2}\varphi_{*}\mathcal{O}_{\mathcal{X}_{b}}=0$ for all $b\in\mathcal{B}$ throughout the proof. First consider the special case when $\mathcal{X}$ is smooth over $\mathcal{B}$. By Lemma 4.14 and [KM-flip-classification]*Theorem (12.1.3), we have
$N^{1}(\mathcal{X}/\mathcal{Z})\cong H_{2}(\mathcal{X}/\mathcal{Z},\mathbb{Q})_{\mathrm{alg}}^{\vee}\cong\mathrm{coker}(H^{2}(\mathcal{Z},\mathbb{Q})\to H^{2}(\mathcal{X},\mathbb{Q}))$
and similarly
$N^{1}(\mathcal{X}_{b}/\mathcal{Z}_{b})\cong\mathrm{coker}(H^{2}(\mathcal{Z}_{b},\mathbb{Q})\to H^{2}(\mathcal{X}_{b},\mathbb{Q})).$
Denote the maps $\mathcal{X}\to\mathcal{B}$, $\mathcal{Z}\to\mathcal{B}$ by $p$ and $q$ respectively. By [Ver-Whitney-stratification]*Corollaire (5.1), there exists a dense, Zariski open subset $U\subseteq\mathcal{B}$ such that $p^{-1}(U)\to U$ and $q^{-1}(U)\to U$ are locally topologically trivial fibrations (in the analytic topology). After replacing $\mathcal{B}$ by $U$, we may therefore assume that $\varphi$ induces a map $\varphi^{*}\colon R^{2}q_{*}\mathbb{Q}\to R^{2}p_{*}\mathbb{Q}$ of local systems.
By the previous discussions, each $N^{1}(\mathcal{X}_{b}/\mathcal{Z}_{b})$ is a fiber of the local system $\mathrm{coker}(\varphi^{*})$. As they also form the local system $\mathcal{G}\mathcal{N}^{1}(\mathcal{X}/B)$ defined in [KM-flip-classification]*Definition (12.2.4), we deduce from [KM-flip-classification]*Proposition (12.2.5) that it has finite monodromy. We can then trivialize it by some dominant generically finite cover $\mathcal{B}^{\prime}\to\mathcal{B}$ to produce the desired isomorphism
$N^{1}(\mathcal{X}^{\prime}/\mathcal{Z}^{\prime})|_{\mathcal{X}^{\prime}_{b}}\cong N^{1}(\mathcal{X}^{\prime}_{b}/\mathcal{Z}^{\prime}_{b}).$
Thus the lemma holds in this case.
In general, let $\pi\colon\mathcal{Y}\to\mathcal{X}$ be a log resolution and let $\psi\colon\mathcal{Y}\to\mathcal{Z}$ the induced map. By shrinking $\mathcal{B}$, we may assume that $\mathcal{Y}$ is smooth over $\mathcal{B}$. As $\mathcal{X}$ has rational singularities, we have $R^{i}\pi_{*}\mathcal{O}_{\mathcal{Y}}=0$ for all $i>0$ and hence $R^{i}\psi_{*}\mathcal{O}_{\mathcal{Y}}=R^{i}\varphi_{*}\mathcal{O}_{\mathcal{X}}$ for all $i\in\mathbb{N}$. Thus from the smooth case treated above, we know that $N^{1}(\mathcal{Y}_{b}/\mathcal{Z}_{b})$ and $N^{1}(\mathcal{Y}_{b}/\mathcal{X}_{b})$ each form a local system with finite monodromy possibly after shrinking $\mathcal{B}$. Since $N^{1}(\mathcal{X}_{b}/\mathcal{Z}_{b})$ is the kernel of the natural restriction map $N^{1}(\mathcal{Y}_{b}/\mathcal{Z}_{b})\to N^{1}(\mathcal{Y}_{b}/\mathcal{X}_{b})$, we deduce that $N^{1}(\mathcal{X}_{b}/\mathcal{Z}_{b})$ also form a local system with finite monodromy and the lemma follows as before. ∎
The following provides the first step in the proof of Theorem 4.13.
Corollary 4.17.
Let $\varphi\colon\mathcal{X}\to\mathcal{Z}$ be a Fano type fibration over a variety $\mathcal{B}$. Then there exists a generically finite dominant morphism $\mathcal{B}^{\prime}\to\mathcal{B}$ such that after base change $\mathcal{X}^{\prime}=\mathcal{X}\times_{\mathcal{B}}\mathcal{B}^{\prime}$, $\mathcal{Z}^{\prime}=\mathcal{Z}\times_{\mathcal{B}}\mathcal{B}^{\prime}$, the natural restriction map
$N^{1}(\mathcal{X}^{\prime}/\mathcal{Z}^{\prime})\to N^{1}(\mathcal{X}^{\prime}_{b}/\mathcal{Z}^{\prime}_{b})$
is an isomorphism for every $b\in\mathcal{B}^{\prime}$.
Proof.
By [KM98]*Theorem 5.22 and Kawamata-Viehweg vanishing, we know that $\mathcal{X}$ has rational singularities and $R^{1}\varphi_{*}\mathcal{O}_{\mathcal{X}}=R^{2}\varphi_{*}\mathcal{O}_{\mathcal{X}}=0$. The statement then follows directly from Lemma 4.16. ∎
By essentially the same argument, we also get the following result.
Lemma 4.18.
Let $\varphi\colon\mathcal{X}\to\mathcal{Z}$ be a Fano type fibration over a variety $\mathcal{B}$. Then there exist a generically finite dominant morphism $\mathcal{B}^{\prime}\to\mathcal{B}$ and a small birational modification $\mathcal{Y}\to\mathcal{X}^{\prime}=\mathcal{X}\times_{\mathcal{B}}\mathcal{B}^{\prime}$ such that $\mathcal{Y}_{b}$ is $\mathbb{Q}$-factorial for every $b\in\mathcal{B}^{\prime}$.
Proof.
Let $\pi\colon\widetilde{\mathcal{X}}\to\mathcal{X}$ be a log resolution. Note that for all $i>0$, we have $R^{i}\pi_{*}\mathcal{O}_{\widetilde{\mathcal{X}}}=0$ as $\mathcal{X}$ has rational singularities by [KM98]*Theorem 5.22, and $R^{i}\varphi_{*}\mathcal{O}_{\mathcal{X}}=0$ by Kawamata-Viehweg vanishing, hence $R^{i}(\varphi\circ\pi)_{*}\mathcal{O}_{\widetilde{\mathcal{X}}}=0$ for all $i>0$ as well. Thus by Lemmas 4.14 and 4.16, possibly after a generically finite base change, we may assume that $\mathcal{B}$ is smooth, $\widetilde{\mathcal{X}}$ is smooth over $\mathcal{B}$, and there exist finitely many line bundles $\mathcal{L}_{1},\dots,\mathcal{L}_{r}$ on $\widetilde{\mathcal{X}}$ whose restriction to the fibers $\widetilde{\mathcal{X}}_{b}$ generate $N^{1}(\widetilde{\mathcal{X}}_{b}/\mathcal{Z}_{b})$ for all $b\in\mathcal{B}$. It follows that the class group $\mathrm{Cl}(\mathcal{X}_{b})$ is generated by $\varphi^{*}\mathrm{Pic}(\mathcal{Z}_{b})$ and $\pi_{*}\mathcal{L}_{i,b}$ ($i=1,\dots,r$). As $\mathcal{X}$ is of Fano type over $\mathcal{Z}$, it has a small $\mathbb{Q}$-factorial modification $\mathcal{Y}\to\mathcal{X}$ by [BCHM, Corollary 1.4.3]. By shrinking $\mathcal{B}$ again, we may assume that the induced map $\mathcal{Y}_{b}\to\mathcal{X}_{b}$ is small and $\mathrm{Cl}(\mathcal{Y}_{b})\cong\mathrm{Cl}(\mathcal{X}_{b})$ is generated by $\mathrm{Pic}(\mathcal{Z}_{b})$ and $\mathcal{L}^{\prime}_{i,b}$ where $\mathcal{L}^{\prime}_{i}$ is the pushforward of $\mathcal{L}_{i}$ to $\mathcal{Y}$. Since $\mathcal{Y}$ is $\mathbb{Q}$-factorial, every $\mathcal{L}^{\prime}_{i}$ is $\mathbb{Q}$-Cartier. This implies that $\mathcal{Y}_{b}$ is $\mathbb{Q}$-factorial. ∎
The next step (Lemma 4.20) in the proof of Theorem 4.13 is to show that the relative pseudo-effective cone in $N^{1}(\mathcal{X}_{b}/\mathcal{Z}_{b})_{\mathbb{R}}$ is also locally constant. Before we prove this, we need an auxiliary result.
Lemma 4.19.
Let $(X,\Delta)$ be a klt pair over $T$. Assume that $\Delta$ has rational coefficients, $\Delta\geq A\geq 0$ for some ample $\mathbb{Q}$-divisor $A$. Then there exist a log resolution $\pi\colon X^{\prime}\to X$, a projective compactification $X^{\prime}\subseteq\overline{X}$ over $T$, and an SNC $\mathbb{Q}$-divisor $\overline{\Delta}$ on $\overline{X}$ such that $\overline{X}\setminus X^{\prime}\subseteq\mathrm{Supp}(\overline{\Delta})$, $\overline{\Delta}\geq\bar{A}\geq 0$ for some ample $\mathbb{Q}$-divisor $\bar{A}$, $(\overline{X},\overline{\Delta})$ is klt, and
$K_{X^{\prime}}+\Delta^{\prime}\sim_{\mathbb{Q}}\pi^{*}(K_{X}+\Delta)+E$
for some $\pi$-exceptional $\mathbb{Q}$-divisor $E\geq 0$, where $\Delta^{\prime}=\overline{\Delta}|_{X^{\prime}}$.
Proof.
By assumption, we may write $\Delta=A+B$ where $B\geq 0$. First consider the case when $(X,\Delta)$ is log smooth. Then there exists a projective compactification $X\subseteq X_{1}$ over $T$ such that $A$ extends to an ample divisor $A_{1}$ on $X_{1}$. Choose some log resolution $\varphi\colon\overline{X}\to X_{1}$ that is an isomorphism over $X$ such that the exceptional locus supports a $\varphi$-ample divisor $-F$ and $D=\overline{X}\setminus X$ is simple normal crossing. Then $\varphi^{*}A_{1}-\varepsilon F$ is ample when $0<\varepsilon\ll 1$ ($\varepsilon\in\mathbb{Q}$). Let $\overline{B}$ be the closure of $B$ in $\overline{X}$. Passing to a higher model we may assume that $\overline{B}$ has SNC support. By Bertini’s theorem, we may choose some $\mathbb{Q}$-divisor $0\leq\overline{A}\sim_{\mathbb{Q}}\varphi^{*}A_{1}-\varepsilon F$ such that $(\overline{X},\overline{\Delta}:=\overline{A}+\overline{B}+\frac{1}{2}D)$ is log smooth and klt. By construction, $\overline{\Delta}$ supports the ample divisor $\overline{A}$ and $\overline{\Delta}|_{X}\sim_{\mathbb{Q}}A+B=\Delta$. Hence the lemma holds in this case (with $X^{\prime}=X$).
In general, write $\Delta=A+B$ as before and take a log resolution $\pi\colon X^{\prime}\to X$ of $(X,\Delta)$ whose exceptional locus supports a $\pi$-ample divisor $-F$. By the negativity lemma we have $F\geq 0$. We may write
$K_{X^{\prime}}+B^{\prime}=\pi^{*}(K_{X}+B)+E$
for some $\mathbb{Q}$-divisors $B^{\prime},E\geq 0$ such that $B^{\prime}$ and $E$ have no common components, and $E$ is $\varphi$-exceptional. Since $(X,B)$ is klt, the coefficients of $B^{\prime}$ are less than $1$, thus the log smooth pair $(X^{\prime},B^{\prime})$ is also klt. Choose some $0<\varepsilon\ll 1$ such that $\varphi^{*}A-\varepsilon F$ is ample and $(X^{\prime},B^{\prime}+\varepsilon F)$ remains klt. By Bertini’s theorem, we may also choose some $0\leq A^{\prime}\sim_{\mathbb{Q}}\pi^{*}A-\varepsilon F$ such that $(X^{\prime},\Delta^{\prime}:=A^{\prime}+B^{\prime}+\varepsilon F)$ is still log smooth and klt. By construction, we see that $K_{X^{\prime}}+\Delta^{\prime}\sim_{\mathbb{Q}}\pi^{*}(K_{X}+\Delta)+E$ and $\Delta^{\prime}$ supports the ample $\mathbb{Q}$-divisor $A^{\prime}$. Thus we may replace $(X,\Delta)$ by $(X^{\prime},\Delta^{\prime})$ and reduce to the log smooth case treated before. ∎
The following gives the second main ingredient in the proof of Theorem 4.13.
Lemma 4.20.
Let $\mathcal{Z}\to\mathcal{B}$ be a dominant morphism of varieties and let $\mathcal{X}\to\mathcal{Z}$ be a Fano type fibration. Then there exists a dense open subset $U\subseteq\mathcal{B}$ such that for any line bundle $\mathcal{L}$ on $\mathcal{X}$ and any closed point $b\in U$, we have $\mathcal{L}_{b}$ is pseudo-effective (resp. big) over $\mathcal{Z}_{b}$ if and only if $\mathcal{L}_{\eta}$ is pseudo-effective (resp. big) over $\mathcal{Z}_{\eta}$, where $\eta$ is the generic point of $\mathcal{B}$.
Proof.
The statement is local on $\mathcal{Z}$, thus we may assume that $\mathcal{Z}$ is affine, in which case a divisor on $\mathcal{X}$ is big (resp. ample) over $\mathcal{Z}$ if and only if it is big (resp. ample) by itself. Since $\mathcal{X}\to\mathcal{Z}$ is a Fano type fibration, by shrinking $\mathcal{Z}$ again we may assume that there exists some $\mathbb{Q}$-divisor $\mathcal{D}\geq 0$ on $\mathcal{X}$ such that $(\mathcal{X},\mathcal{D})$ is klt, $K_{\mathcal{X}}+\mathcal{D}\sim_{\mathbb{Q}}0$, and $\mathcal{D}\geq A\geq 0$ for some ample $\mathbb{Q}$-divisor $A$. Since a divisor $\mathcal{L}$ on $\mathcal{X}$ is big if and only if $\mathcal{L}-\varepsilon A$ is pseudo-effective for some $\varepsilon>0$, the pseudo-effective version of the lemma implies the big version. Thus it suffices to consider the pseudo-effective case of the statement.
Since $\mathcal{X}\to\mathcal{Z}$ is a Fano type fibration, by applying [BCHM]*Corollary 1.3.2 to a small $\mathbb{Q}$-factorialization of $\mathcal{X}$, we know that there are finitely many divisors $\mathcal{D}_{i}\geq 0$ on $\mathcal{X}$ that generate the (relative) pseudo-effective cone; in other words, for any $\mathbb{R}$-divisor $\mathcal{L}$ on $\mathcal{X}$ that is pseudo-effective over $\mathcal{Z}$ (equivalently, $\mathcal{L}_{\eta}$ is pseudo-effective over $\mathcal{Z}_{\eta}$) we have $\mathcal{L}\sim_{\mathbb{R},\mathcal{Z}}\sum a_{i}\mathcal{D}_{i}$ for some $a_{i}\geq 0$. By shrinking $\mathcal{B}$ we may assume that $\mathcal{B}$ is smooth and affine, $\mathcal{X}$ is flat over $\mathcal{B}$, $\mathcal{X}_{b}$ is normal, and $\mathcal{D}_{i}$ does not contain $\mathcal{X}_{b}$ for all $i$ and all $b\in\mathcal{B}$. In particular, the restriction $\mathcal{D}_{i,b}$ is well-defined and we have $\mathcal{L}_{b}\sim_{\mathbb{R},\mathcal{Z}_{b}}\sum a_{i}\mathcal{D}_{i,b}$, hence $\mathcal{L}_{b}$ is pseudo-effective over $\mathcal{Z}_{b}$ whenever $\mathcal{L}_{\eta}$ is pseudo-effective over $\mathcal{Z}_{\eta}$.
It remains to show that the converse direction holds after possibly further shrinking $\mathcal{B}$. For this, we first note that since $\mathcal{Z}$ is affine, we can factorize $\mathcal{Z}\to\mathcal{B}$ into an open immersion $\mathcal{Z}\to\overline{\mathcal{Z}}$ followed by a projective morphism $\overline{\mathcal{Z}}\to\mathcal{B}$ such that there is an ample Cartier divisor $\mathcal{M}_{0}$ on $\overline{\mathcal{Z}}$ with $\mathrm{Supp}(\mathcal{M}_{0})=\overline{\mathcal{Z}}\setminus\mathcal{Z}$. By Lemma 4.19, there exist a log resolution $\pi\colon\mathcal{X}^{\prime}\to\mathcal{X}$, a projective compactification $\mathcal{X}^{\prime}\subseteq\overline{\mathcal{X}}$ over $\overline{\mathcal{Z}}$ and an SNC $\mathbb{Q}$-divisor $\overline{\mathcal{D}}$ on $\overline{\mathcal{X}}$ such that $\overline{\mathcal{X}}\setminus\mathcal{X}^{\prime}\subseteq\mathrm{Supp}(\overline{\mathcal{D}})$, $\overline{\mathcal{D}}=\overline{A}+\overline{B}$ where $\overline{A}\geq 0$ is an ample $\mathbb{Q}$-divisor and $\overline{B}\geq 0$, $(\overline{\mathcal{X}},\overline{\mathcal{D}})$ is klt, and
$K_{\mathcal{X}^{\prime}}+\mathcal{D}^{\prime}\sim_{\mathbb{Q}}\pi^{*}(K_{\mathcal{X}}+\mathcal{D})+\mathcal{E}$
for some $\pi$-exceptional $\mathbb{Q}$-divisor $\mathcal{E}\geq 0$, where $\mathcal{D}^{\prime}=\overline{\mathcal{D}}|_{\mathcal{X}^{\prime}}$. Let $U\subseteq\mathcal{B}$ be an affine open subset such that $(\mathcal{X}_{b},\mathcal{D}_{b})$ is klt (in particular, this implies that $\mathcal{X}_{b}\to\mathcal{Z}_{b}$ is a Fano type fibration) for all $b\in U$, and $(\overline{\mathcal{X}},\overline{\mathcal{D}})$ is log smooth over $U$. Let $\mathcal{M}$ be the pullback of $\mathcal{M}_{0}$ to $\overline{\mathcal{X}}$. By construction, $\overline{\mathcal{X}}$ is projective over $\mathcal{B}$, the divisor $\mathcal{M}$ is semiample over $\mathcal{B}$, and $\mathrm{Supp}(\mathcal{M})=\overline{\mathcal{X}}\setminus\mathcal{X}^{\prime}$.
Now let $\mathcal{L}$ be a line bundle on $\mathcal{X}$ such that $\mathcal{L}_{b}$ is pseudo-effective over $\mathcal{Z}_{b}$. Then $H^{0}(\mathcal{X}_{b},\ell\mathcal{L}_{b})\neq 0$ for some positive integer $\ell$ since $\mathcal{X}_{b}$ is of Fano type over $\mathcal{Z}_{b}$ and $\mathcal{Z}_{b}$ is affine. As $\overline{\mathcal{X}}$ is smooth, we may extend $\mathcal{L}^{\prime}:=\pi^{*}\mathcal{L}$ to a line bundle $\overline{\mathcal{L}}$ on $\overline{\mathcal{X}}$. Note that for all sufficiently divisible positive integer $m$ we have
$m(K_{\mathcal{X}^{\prime}}+\mathcal{D}^{\prime})+\ell\mathcal{L}^{\prime}\sim m(\pi^{*}(K_{\mathcal{X}}+\mathcal{D})+\mathcal{E})+\ell\pi^{*}\mathcal{L}\sim m\mathcal{E}+\ell\pi^{*}\mathcal{L}.$
As $H^{0}(\mathcal{X}_{b},\ell\mathcal{L}_{b})\neq 0$ and $\mathcal{E}\geq 0$, this implies that
$H^{0}(\mathcal{X}^{\prime}_{b},m(K_{\mathcal{X}^{\prime}_{b}}+\mathcal{D}^{\prime}_{b})+\ell\mathcal{L}^{\prime}_{b})\neq 0.$
Since $\mathcal{E}$ is $\pi$-exceptional, it also implies that $\mathcal{L}$ is pseudo-effective as long as $m(K_{\mathcal{X}^{\prime}}+\mathcal{D}^{\prime})+\ell\mathcal{L}^{\prime}$ is pseudo-effective for some positive integers $\ell,m$. Choose some sufficiently large and divisible integer $m$ (which depends on $\ell$) such that $m\overline{A}$ and $m\overline{B}$ are Cartier, and $m\overline{A}+\ell\overline{\mathcal{L}}$ is ample. Since $H^{0}(\mathcal{X}^{\prime}_{b},m(K_{\mathcal{X}^{\prime}_{b}}+\mathcal{D}^{\prime}_{b})+\ell\mathcal{L}^{\prime}_{b})\neq 0$ and $\mathrm{Supp}(\mathcal{M})=\overline{\mathcal{X}}\setminus\mathcal{X}^{\prime}$, we may take a sufficiently large integer $r$ (depending on $\ell$ and $m$) such that
(4.5) $H^{0}(\overline{\mathcal{X}}_{b},m(K_{\overline{\mathcal{X}}_{b}}+\overline{\mathcal{D}}_{b})+\ell\overline{\mathcal{L}}_{b}+r\mathcal{M}_{b})\neq 0.$
By construction, we have
$m(K_{\overline{\mathcal{X}}}+\overline{\mathcal{D}})+\ell\overline{\mathcal{L}}+r\mathcal{M}=m(K_{\overline{\mathcal{X}}}+\overline{B})+(m\overline{A}+\ell\overline{\mathcal{L}}+r\mathcal{M}).$
Note that $(\overline{\mathcal{X}},\overline{B})$ is log smooth over $\mathcal{B}$, and $m\overline{A}+\overline{\mathcal{L}}+r\mathcal{M}$ is ample, thus by (4.5) and [HMX-BirAut]*Theorem 1.8, we deduce that
$H^{0}(\overline{\mathcal{X}},m(K_{\overline{\mathcal{X}}}+\overline{\mathcal{D}})+\ell\overline{\mathcal{L}}+r\mathcal{M})\neq 0.$
By restriction to $\mathcal{X}^{\prime}$ we also get a section of $m(K_{\mathcal{X}^{\prime}}+\mathcal{D}^{\prime})+\ell\mathcal{L}^{\prime}$. From the previous discussion, we deduce that $\mathcal{L}$ is pseudo-effective and this concludes the proof. ∎
We now return to the proof of Theorem 4.13. We first consider a special case.
Lemma 4.21.
Let $\mathcal{S}$ be a bounded set of Fano type fibrations. Then there exists a bounded set $\mathcal{T}$ of Fano type fibrations such that for any $(X,\Delta)\to Z$ in $\mathcal{S}$ and any birational contraction $f\colon X\dashrightarrow X^{\prime}$ over $Z$, we have $(X^{\prime},\Delta^{\prime}=f_{*}\Delta)\to Z$ belongs to $\mathcal{T}$.
Proof.
We may assume that $\mathcal{S}$ is given by the fibers of a Fano type fibration $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ over a finite type base $\mathcal{B}$. By induction on $\dim\mathcal{B}$, it suffices to prove the statement after some generically finite dominant base change $\mathcal{B}^{\prime}\to\mathcal{B}$. Thus we may shrink $\mathcal{B}$ freely as needed. In particular we assume that $\mathcal{B}$ is smooth and connected throughout the proof.
By Lemma 4.18, possibly after a generically finite dominant base change and replacing $\mathcal{X}$ by a small birational modification, we may assume that $\mathcal{X}_{b}$ is $\mathbb{Q}$-factorial for all $b\in\mathcal{B}$. Note that a small birational modification does not change the set of birational contractions. By Corollary 4.17, possibly after another generically finite dominant base change, we may also assume that $N^{1}(\mathcal{X}_{b}/\mathcal{Z}_{b})$ form a trivial local system. By [BCHM]*Corollary 1.3.1, there are only finitely many birational contractions $f_{i}\colon\mathcal{X}\dashrightarrow\mathcal{X}_{i}$ ($i=1,\dots,m$) over $\mathcal{Z}$. Let $\mathcal{E}_{i}$ be the sum of all the exceptional divisors of $f_{i}$. By shrinking the base $\mathcal{B}$ again, we can assume that the following holds simultaneously for all $i$ and all $b\in\mathcal{B}$:
  • $\mathcal{X}_{i,b}$ is normal,
  • $\mathcal{E}_{i,b}\neq 0$ whenever $\mathcal{E}_{i}\neq 0$,
  • the map $f_{i}$ is defined at the generic point of $\mathcal{X}_{b}$,
  • the induced map $f_{i,b}\colon\mathcal{X}_{b}\dashrightarrow\mathcal{X}_{i,b}$ is a birational contraction,
  • $\mathcal{E}_{i,b}$ is contained in the exceptional locus of $f_{i,b}$, and
  • $f_{i,b}$ is small whenever $f_{i}$ is small.
Let $\mathcal{X}_{b}\dashrightarrow Y$ be a birational contraction over $\mathcal{Z}_{b}$. Let $H$ be a very ample Cartier divisor on $Y$ and let $\mathcal{L}_{b}$ be its birational pullback to $\mathcal{X}_{b}$. Then $\mathcal{L}_{b}$ is big and movable (i.e. the base locus of $|\mathcal{L}_{b}|$ has codimension at least $2$ in $\mathcal{X}_{b}$). It is also $\mathbb{Q}$-Cartier since $\mathcal{X}_{b}$ is $\mathbb{Q}$-factorial. By Lemma 4.20, possibly after replacing $\mathcal{L}_{b}$ by a multiple we may extend it to a big line bundle $\mathcal{L}$ on $\mathcal{X}$. The ample model of $\mathcal{L}$ over $\mathcal{Z}$ necessarily coincides with one of the $f_{i}\colon\mathcal{X}\dashrightarrow\mathcal{X}_{i}$. We claim that $Y\cong\mathcal{X}_{i,b}$.
To see this, first note that the $\mathcal{L}$-MMP over $\mathcal{Z}$ coincides with some $f_{j}\colon\mathcal{X}\dashrightarrow\mathcal{X}_{j}$. We first show that $f_{j}$ is small. Let $\mathcal{L}_{j}=f_{j*}\mathcal{L}$ which is nef over $\mathcal{Z}$. By the negativity lemma (see e.g. [KM98]*Lemma 3.38), we have $\mathcal{L}=f_{j}^{*}\mathcal{L}_{j}+\mathcal{E}$ for some divisor $\mathcal{E}\geq 0$ whose support contains every exceptional divisor of $f_{j}$. If $f_{j}$ is not small, then $\mathcal{E}\neq 0$. Restricting to the fiber we see that $\mathcal{L}_{b}=f_{j,b}^{*}\mathcal{L}_{j,b}+\mathcal{E}_{b}$ and hence the base locus of $\mathcal{L}_{b}$ contains the divisor $\mathcal{E}_{b}\neq 0$, a contraction as $\mathcal{L}_{b}$ is movable. Thus $f_{j}$ is small and so is $f_{j,b}$ (by the last bullet point above). The ample model $\mathcal{X}_{i}$ (resp. $Y$) of $\mathcal{L}$ (resp. $\mathcal{L}_{b}$) is therefore the same as the ample model of $\mathcal{L}_{j}$ (resp. $\mathcal{L}_{j,b}$) over $\mathcal{Z}$ (resp. $\mathcal{Z}_{b}$). Since $\mathcal{L}_{j}$ is nef over $\mathcal{Z}$, its ample model is a morphism $\mathcal{X}_{j}\to\mathcal{X}_{i}$; combined with the fact that $\mathcal{X}_{i,b}$ is normal, we see that $\mathcal{X}_{i,b}$ is also the ample model of $\mathcal{L}_{j,b}$. This proves that $Y\cong\mathcal{X}_{i,b}$. It follows that every birational contraction of $(\mathcal{X}_{b},\mathcal{D}_{b})$ over $\mathcal{Z}_{b}$ is a fiber of the finitely many families $(\mathcal{X}_{i},\mathcal{D}_{i}=f_{i*}\mathcal{D})\to\mathcal{Z}$ over $\mathcal{B}$. ∎
Proof of Theorem 4.13.
By Lemma 2.33, we may assume that $\mathcal{S}$ is given by the fibers of an $\mathbb{R}$-Gorenstein family $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ of log Fano fibrations over a finite type base $\mathcal{B}$. By the proof of [HX-CY-bdd]*Proposition 2.5 (or by Lemma 4.18 and loc. cit.), possibly after a base change via some surjective morphism $\mathcal{B}^{\prime}\to\mathcal{B}$, there exists a crepant birational morphism $(\mathcal{Y},\mathcal{E})\to(\mathcal{X},\mathcal{D})$ whose restriction to the fiber over $b\in\mathcal{B}$ extracts every divisor of non-positive discrepancy over $(\mathcal{X}_{b},\mathcal{D}_{b})$. Note that $(\mathcal{Y},\mathcal{E})$ is also of Fano type over $\mathcal{Z}$ by Lemma 2.29, and every birational model $f\colon X^{\prime}\dashrightarrow\mathcal{X}_{b}$ such that $f^{*}(K_{\mathcal{X}_{b}}+\mathcal{D}_{b})=K_{X^{\prime}}+\Delta^{\prime}$ for some $\Delta^{\prime}\geq 0$ can be obtained by a birational contraction $\mathcal{Y}_{b}\dashrightarrow X^{\prime}$. The theorem thus follows from Lemma 4.21. ∎
4.5. Special boundedness
To prove Theorems 4.2 and 4.3, we still need to upgrade Theorem 4.13 to allow specially bounded input. The precise statement is the following.
Theorem 4.22.
Let $\mathcal{S}$ be a specially bounded set of log Fano fibration germs. Let $\mathcal{T}$ be the set of Fano type fibration germs $(X^{\prime},\Delta^{\prime})\to Z\ni z$ for which there exist some $(X,\Delta)\to Z\ni z$ in $\mathcal{S}$ and a birational map $f\colon X^{\prime}\dashrightarrow X$ over $Z$ such that $f^{*}(K_{X}+\Delta)\geq K_{X^{\prime}}+\Delta^{\prime}$. Then $\mathcal{T}$ is log specially bounded.
To prove this, we need a few auxiliary lemmas.
Lemma 4.23.
Let $(X,G+D)$ be a log canonical pair where $G$ is a normal Weil divisor and $D\geq 0$, and let $X\to Z$ a projective morphism such that $-(K_{X}+G+D)$ is semiample over $Z$. Let $f\colon X^{\prime}\dashrightarrow X$ be a birational map over $Z$ where $X^{\prime}$ is projective over $Z$. Let $G^{\prime}$ be the strict transform of $G$ on $X^{\prime}$ and let $D^{\prime}$ be an $\mathbb{R}$-divisor on $X^{\prime}$ such that
$f^{*}(K_{X}+G+D)\geq K_{X^{\prime}}+G^{\prime}+D^{\prime}.$
Assume that $G^{\prime}$ is normal and Cartier on $X^{\prime}$. Then for the induced birational map $f_{G}\colon G^{\prime}\dashrightarrow G$, we have
$f_{G}^{*}(K_{G}+\mathrm{Diff}_{G}(D))\geq K_{G^{\prime}}+D^{\prime}|_{G^{\prime}}.$
Proof.
By enlarging $D^{\prime}$ we may assume that $f^{*}(K_{X}+G+D)=K_{X^{\prime}}+G^{\prime}+D^{\prime}$. The statement is local on $Z$, hence we may also assume that $-(K_{X}+G+D)$ is semiample. By Bertini’s theorem, we can choose some $\mathbb{R}$-divisor $0\leq H\sim_{\mathbb{R}}-(K_{X}+G+D)$ such that $H$ (resp. $H|_{G}$) does not contain any exceptional divisor of $f$ (resp. $f_{G}$) in its support. In particular, we may assume that $f$ is an isomorphism at every generic point of $\mathrm{Supp}(H)\cap G$, hence $f^{*}H=H^{\prime}$ and $f_{G}^{*}(H|_{G})\leq H^{\prime}|_{G^{\prime}}$, where $H^{\prime}$ is the strict transform of $H$ on $X^{\prime}$. As $K_{X}+G+D+H\sim_{\mathbb{R}}0$, we also have $f^{*}(K_{X}+G+D+H)=K_{X^{\prime}}+G^{\prime}+D^{\prime}+H^{\prime}\sim_{\mathbb{R}}0$ and this is a crepant pullback. By adjunction, we deduce that
$f_{G}^{*}(K_{G}+\mathrm{Diff}_{G}(D+H))=K_{G^{\prime}}+\mathrm{Diff}_{G^{\prime}}(D^{\prime}+H^{\prime}).$
Note that $\mathrm{Diff}_{G}(D+H)=\mathrm{Diff}_{G}(D)+H|_{G}$ since $H$ is $\mathbb{R}$-Cartier and $\mathrm{Diff}_{G^{\prime}}(D^{\prime}+H^{\prime})=(D^{\prime}+H^{\prime})|_{G^{\prime}}$ since $G^{\prime}$ is Cartier. Thus
$f_{G}^{*}(K_{G}+\mathrm{Diff}_{G}(D))=K_{G^{\prime}}+(D^{\prime}+H^{\prime})|_{G^{\prime}}-f_{G}^{*}(H|_{G})\geq K_{G^{\prime}}+D^{\prime}|_{S^{\prime}}$
as desired. ∎
Lemma 4.24.
Let $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ be a special test configuration of a log Fano fibration $(X,\Delta)\to Z$. Let $(X^{\prime},\Delta^{\prime})$ be a projective pair over $Z$ and $f\colon X^{\prime}\dashrightarrow X$ a birational map over $Z$ such that
(4.6) $f^{*}(K_{X}+\Delta)\geq K_{X^{\prime}}+\Delta^{\prime}.$
Then there exist a Fano type test configuration $(\mathcal{X}^{\prime},\mathcal{D}^{\prime})\to\mathcal{Z}$ of $(X^{\prime},\Delta^{\prime})\to Z$ and a birational map $\mathcal{X}^{\prime}\dashrightarrow\mathcal{X}$ over $\mathcal{Z}$ extending $f$, such that $\mathcal{X}^{\prime}_{0}$ is the strict transform of $\mathcal{X}_{0}$ and the induced birational map $f_{0}\colon\mathcal{X}^{\prime}_{0}\dashrightarrow\mathcal{X}_{0}$ over $\mathcal{Z}_{0}$ satisfies
(4.7) $f_{0}^{*}(K_{\mathcal{X}_{0}}+\mathcal{D}_{0})\geq K_{\mathcal{X}^{\prime}_{0}}+\mathcal{D}^{\prime}_{0}.$
Proof.
Let $E_{1},\dots,E_{r}$ be the prime divisors on $X^{\prime}$ that are contracted by $f$. Through the isomorphism $\mathcal{X}\setminus\mathcal{X}_{0}\cong X\times(\mathbb{A}^{1}\setminus\{0\})$, each $E_{i}\times(\mathbb{A}^{1}\setminus\{0\})$ is identified with a divisor $\mathcal{E}_{i}$ over $\mathcal{X}$. By assumption, $a(E_{i},X,\Delta)\leq 0$, thus $a(\mathcal{E}_{i},\mathcal{X},\mathcal{D})\leq 0$. By [BCHM]*Corollary 1.4.3, we know that there exists a projective birational morphism $h\colon\mathcal{Y}\to\mathcal{X}$ such that $\mathcal{Y}$ is $\mathbb{Q}$-factorial and the exceptional divisors are exactly given by $\mathcal{E}_{1},\dots,\mathcal{E}_{r}$. Since $\mathcal{E}_{i}$ is invariant under the $\mathbb{G}_{m}$-action, we also know that $h$ is $\mathbb{G}_{m}$-equivariant. By construction, $\mathcal{Y}_{0}$ is irreducible (as all the $h$-exceptional divisors dominate $\mathbb{A}^{1}$), and the induced birational map $\mathcal{Y}\setminus\mathcal{Y}_{0}\dashrightarrow X^{\prime}\times(\mathbb{A}^{1}\setminus\{0\})$ is a birational contraction.
Let $L$ be an ample line bundle on $X^{\prime}$ and choose some Weil divisor $\mathcal{L}$ on $\mathcal{Y}$ whose restriction to $\mathcal{Y}\setminus\mathcal{Y}_{0}$ is the birational pullback of $L\times(\mathbb{A}^{1}\setminus\{0\})$. Since $\mathcal{Y}$ is $\mathbb{Q}$-factorial, we know that $\mathcal{L}$ is $\mathbb{Q}$-Cartier. Note that $\mathcal{Y}$ is of Fano type over $\mathcal{Z}$ by Lemma 2.29, hence the ample model $g\colon\mathcal{Y}\dashrightarrow\mathcal{X}^{\prime}$ of $\mathcal{L}$ over $\mathcal{Z}$ exists and is automatically $\mathbb{G}_{m}$-equivariant. By construction we have $\mathcal{X}^{\prime}\setminus\mathcal{X}^{\prime}_{0}\cong X^{\prime}\times(\mathbb{A}^{1}\setminus\{0\})$ and the induced map $\widetilde{f}\colon\mathcal{X}^{\prime}\dashrightarrow\mathcal{X}$ extends $f$. Note that as $\mathcal{Y}_{0}$ is irreducible and semiample, it is not in the base locus of $\mathcal{L}$, thus $g$ is an isomorphism at the generic point of $\mathcal{Y}_{0}$. It follows that $\mathcal{X}^{\prime}_{0}$ is the strict transform of $\mathcal{Y}_{0}$ (in particular, it is irreducible), and we have an induced birational map $f_{0}\colon\mathcal{X}^{\prime}_{0}\dashrightarrow\mathcal{X}_{0}$.
Let $\mathcal{D}^{\prime}$ be the closure of $\Delta^{\prime}\times(\mathbb{A}^{1}\setminus\{0\})$ in $\mathcal{X}^{\prime}$. By (4.6), we have
$\widetilde{f}^{*}(K_{\mathcal{X}}+\mathcal{D})\geq K_{\mathcal{X}^{\prime}}+\mathcal{D}^{\prime}.$
As $(\mathcal{X},\mathcal{D})\to\mathcal{Z}$ is a special test configuration, we see that $(\mathcal{X}^{\prime},\mathcal{D}^{\prime})\to\mathcal{Z}$ is a Fano type test configuration of $(X^{\prime},\Delta^{\prime})\to Z$ by Lemma 2.35. Note that we also have $\widetilde{f}^{*}\mathcal{X}_{0}=\mathcal{X}^{\prime}_{0}$, hence $\widetilde{f}^{*}(K_{\mathcal{X}}+\mathcal{X}_{0}+\mathcal{D})\geq K_{\mathcal{X}^{\prime}}+\mathcal{X}^{\prime}_{0}+\mathcal{D}^{\prime}$. Thus we obtain (4.7) by Lemma 4.23. ∎
We are ready to prove Theorem 4.22.
Proof of Theorem 4.22.
By assumption, there exists a bounded set $\mathcal{S}_{0}$ of log Fano fibrations such that any fibration germ $(X,\Delta)\to Z\ni z$ in $\mathcal{S}$ specially degenerates to some $(X_{0},\Delta_{0})\to Z_{0}\ni z_{0}$ in $\mathcal{S}_{0}$. If $f\colon X^{\prime}\dashrightarrow X$ is a birational map over $Z$ such that $f^{*}(K_{X}+\Delta)\geq K_{X^{\prime}}+\Delta^{\prime}$ for some $\Delta^{\prime}\geq 0$, then by Lemma 4.24 we get an induced Fano type test configuration of $(X^{\prime},\Delta^{\prime})\to Z\ni z$ whose central fiber is of the form $(X^{\prime}_{0},\Delta^{\prime}_{0})\to Z_{0}\ni z_{0}$ for some birational model $f_{0}\colon X^{\prime}_{0}\dashrightarrow X_{0}$ such that $f_{0}^{*}(K_{X_{0}}+\Delta_{0})\geq K_{X^{\prime}_{0}}+\Delta^{\prime}_{0}$. Write $f_{0}^{*}(K_{X_{0}}+\Delta_{0})=K_{X^{\prime}_{0}}+D^{\prime}_{0}$. Then $D^{\prime}_{0}\geq\Delta^{\prime}_{0}\geq 0$ and $\mathrm{Supp}(\Delta^{\prime}_{0})\subseteq\mathrm{Supp}(D^{\prime}_{0})$. By Theorem 4.13, the set of such fibrations $(X^{\prime}_{0},D^{\prime}_{0})\to Z_{0}$ is bounded, hence the set of $(X^{\prime},\Delta^{\prime})\to Z\ni z$ is log specially bounded. ∎
We now have all the ingredients for the proof of Theorems 4.2 and 4.3.
Proof of Theorem 4.2.
Let $X\to Z\ni z$ be a Fano type fibration germ such that $\widehat{\rm Vol}_{X}(\Delta)\geq\varepsilon$ for some $\mathbb{R}$-divisor $0\leq\Delta\sim_{\mathbb{R},Z}-K_{X}$. By [BCHM]*Corollary 1.4.3, a small $\mathbb{Q}$-factorial modification $X^{\prime\prime}\to X$ exists. By Lemma 2.29 and [BCHM]*Corollary 1.3.2, the ample model $X^{\prime\prime}\dashrightarrow X^{\prime}$ of $-K_{X^{\prime\prime}}$ over $Z$ also exists. Note that $X^{\prime}\to Z$ is a log Fano fibration by construction. Let $h\colon X\dashrightarrow X^{\prime}$ be the induced birational contraction and let $\Delta^{\prime}$ be the strict transform of $\Delta$ on $X^{\prime}$. Since $K_{X}+\Delta\sim_{\mathbb{R},Z}0$, we have $K_{X^{\prime}}+\Delta^{\prime}\sim_{\mathbb{R},Z}0$ as well. In particular, the birational map $(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ is crepant. By Lemma 3.11, we have $\widehat{\rm Vol}_{X^{\prime}}(\Delta^{\prime})\geq\widehat{\rm Vol}_{X}(\Delta)\geq\varepsilon$, hence by Theorem 4.5, we know that $X^{\prime}\to Z\ni z$ belongs to a specially bounded set of log Fano fibration germs.
By the negativity lemma, we have $-K_{X}\geq h^{*}(-K_{X^{\prime}})$, or $h^{*}K_{X^{\prime}}\geq K_{X}$. By Theorem 4.22, we see that the Fano type fibration germ $X\to Z\ni z$ belongs to a log specially bounded set. By definition, this implies that the set $\mathcal{S}$ of such fibrations is bounded up to Fano type degenerations. ∎
As explained in Section 4.1, this also completes the proof of Theorem 4.1.
Proof of Theorem 4.3.
By perturbing $\delta$, we may assume that $\delta\in\mathbb{Q}$. Let $(X,\Delta)\to Z\ni z$ be a log Calabi-Yau fibration germ in $\mathcal{S}$. Let $\Delta_{0}=\frac{\delta}{2}\mathrm{Supp}(\Delta)$. Since $X$ is of Fano type and $(X,\Delta)$ is log Calabi-Yau over $Z$, it is not hard to see that their convex combination $(X,\frac{1}{2}\Delta)$ is of Fano type over $Z$. Therefore, since $\Delta_{0}\leq\frac{1}{2}\Delta$, we see that $(X,\Delta_{0})$ is of Fano type over $Z$ as well. Similar to the above proof of Theorem 4.2, let $h\colon X\dashrightarrow X^{\prime}$ be the ample model of $-(K_{X}+\Delta_{0})$ which exists by [BCHM]*Corollary 1.3.2. Let $\Delta^{\prime}=h_{*}\Delta$ and $\Delta^{\prime}_{0}=h_{*}\Delta_{0}$. We have $\widehat{\rm Vol}_{X^{\prime}}(\Delta^{\prime})\geq\varepsilon$ by Lemma 3.11, and $h^{*}(K_{X^{\prime}}+\Delta_{0}^{\prime})\geq K_{X}+\Delta_{0}$ by the negativity lemma. By Lemma 3.14, we also have $\widehat{\rm Vol}_{X^{\prime},\Delta^{\prime}_{0}}(\Delta^{\prime}-\Delta^{\prime}_{0})\geq\frac{\varepsilon}{2^{n}}$ since $\Delta^{\prime}_{0}\leq\frac{1}{2}\Delta^{\prime}$ by construction. Note that $K_{X^{\prime}}+\Delta^{\prime}\sim_{\mathbb{R},Z}0$ and thus $\Delta^{\prime}-\Delta^{\prime}_{0}\sim_{\mathbb{R},Z}-(K_{X^{\prime}}+\Delta^{\prime}_{0})$. By Theorem 4.5, it follows that the log Fano fibration germ $(X^{\prime},\Delta^{\prime}_{0})\to Z\ni z$ belongs to a specially bounded set. By Theorem 4.22, the Fano type fibrations $(X,\Delta_{0})\to Z\ni z$ belong to a log specially bounded set. As $\mathrm{Supp}(\Delta_{0})=\mathrm{Supp}(\Delta)$, the set $\mathcal{S}$ is also log specially bounded. ∎
5. Termination in low dimension
In this section, we apply the results obtained in Section 3 to prove some effective termination result for general type MMPs in dimensions at most $5$. In particular, we prove Theorem 1.3.
5.1. Termination
We first prove a non-effective version of termination in dimensions four and five.
Theorem 5.1.
Let $(X,\Delta)$ be a klt pair. Assume that $\Delta$ is big, and either
  1. (1)
    $\dim X=4$, or
  2. (2)
    $\dim X=5$ and $K_{X}+\Delta\sim_{\mathbb{R}}D\geq 0$ for some $\mathbb{R}$-divisor $D$.
Then any $(K_{X}+\Delta)$-MMP terminates.
Before proving this result, we give an immediate corollary.
Corollary 5.2.
Let $(X,\Delta)$ be a klt pair of dimension at most $5$ and $\pi:X\rightarrow T$ a projective morphism. Assume that at least one of the following holds.
  1. (1)
    $K_{X}+\Delta$ is big over $T$.
  2. (2)
    $K_{X}+\Delta$ is pseudo-effective over $T$ and $\Delta$ is big over $T$.
  3. (3)
    $\dim X=4$ and $\Delta$ is big over $T$.
Then any $(K_{X}+\Delta)$-MMP over $T$ terminates.
Proof.
By Lemma 2.18, we may assume that $\Delta$ is big in all cases. Then (3) follows from Theorem 5.1(1), and (1) follows from (2), so we only need to prove (2) when $\dim X=5$.
By [BCHM]*Theorem D, there exists an $\mathbb{R}$-divisor $D\geq 0$ on $X$, such that $K_{X}+\Delta\sim_{\mathbb{R},T}D$, or equivalently, there exists an $\mathbb{R}$-divisor $G_{T}$ on $T$ such that $K_{X}+\Delta\sim_{\mathbb{R}}D+\pi^{*}G_{T}$. Let $H_{T}\geq 0$ be an ample divisor on $T$ such that $H_{T}+G_{T}\geq 0$. By Bertini’s theorem, there exists an $\mathbb{R}$-divisor $0\leq\Delta^{\prime}\sim_{\mathbb{R}}\Delta+\pi^{*}H_{T}$ on $X$ such that $(X,\Delta^{\prime})$ is klt. Then $K_{X}+\Delta^{\prime}\sim_{\mathbb{R}}K_{X}+\Delta+\pi^{*}H_{T}\sim_{\mathbb{R}}D+\pi^{*}(G_{T}+H_{T})\geq 0$ and any $(K_{X}+\Delta)$-MMP over $T$ is a also a $(K_{X}+\Delta^{\prime})$-MMP, thus we conclude (2) by Theorem 5.1(2). ∎
In the rest of this subsection we prove Theorem 5.1. We first consider the 4-fold case. In view of Theorem 3.2, the key is to prove the following statement.
Lemma 5.3.
Let $(X,\Delta)$ be a klt pair of dimension at most $4$ and let $S\subseteq\mathbb{R}$ be a discrete set. Let
[…]
be a face-contracting $(K_{X}+\Delta)$-MMP (Definition 2.14) such that $a(E,X_{i},\Delta_{i})\in S$ for all $i$ and all prime divisors $E$ over $X$. Then this face-contracting MMP terminates.
Proof.
Since $(X_{i},\Delta_{i})$ is klt, by [KM98, Proposition 2.36] there are only finitely many prime divisors $E$ over $X$ with $a(E,X_{i},\Delta_{i})\leq 0$. The number $l_{i}$ of such divisors is non-increasing in the MMP, thus after possibly truncating the MMP we may assume that this number stabilizes, i.e. $l_{0}=l_{1}=\cdots=l$ for some $l\in\mathbb{N}$. Let $E_{1},\dots,E_{l}$ be the corresponding prime divisors over $X$ with $a(E_{j},X_{i},\Delta_{i})\leq 0$ for all $i$ and $j$.
Since $S$ is discrete, $S\cap(-1,0]$ is a finite set. Thus as $a(E_{j},X_{i},\Delta_{i})\in S\cap(-1,0]$ by assumption and $a(E_{j},X_{i+1},\Delta_{i+1})\geq a(E_{j},X_{i},\Delta_{i})$ for all $i$ and $j$, after another truncation of the MMP we may assume that
$a(E_{j},X_{0},\Delta_{0})=a(E_{j},X_{1},\Delta_{1})=\cdots=a(E_{j},X_{i},\Delta_{i})=\cdots$
for all $j=1,\cdots,l$. Let $\pi:(Y,D)\rightarrow(X,\Delta)$ be a $\mathbb{Q}$-factorial terminalization of $(X,\Delta)$. By Lemma 2.17, there exists a sequence of birational maps
[…]
such that each $(Y_{i},D_{i})$ is a $\mathbb{Q}$-factorial terminalization of $(X_{i},\Delta_{i})$ and each $\psi_{i}$ is composed of a sequence of steps of a $(K_{Y_{i}}+D_{i})$-MMP. As $\dim X\leq 4$, this sequence terminates by Lemma 2.20, hence the original face-contracting $(K_{X}+\Delta)$-MMP also terminates. ∎
Proof of Theorem 5.1 when $\dim X=4$.
By Theorem 3.1, the Cartier index of $\mathbb{Q}$-Cartier Weil divisors is uniformly bounded in the MMP, thus by Lemma 2.2, the discrepancy of primes divisors over $X$ takes values in some discrete set in the MMP. We then conclude by Lemma 5.3. ∎
We move on to the $5$-fold case of Theorem 5.1. A key observation is that since the Cariter index of $\mathbb{Q}$-Cartier Weil divisors are bounded from above by Theorem 3.2, in Birkar’s inductive approach to termination [Bir07], we only need termination of flips in lower dimensions when the index is bounded.
Lemma 5.4.
Let $(Y,D)$ be a $\mathbb{Q}$-factorial dlt pair of dimension $5$. Let
[…]
be a $(K_{Y}+D)$-MMP. Assume that there exists some integer $r\in\mathbb{N}$ such that the Cartier index of every $\mathbb{Q}$-Cartier Weil divisor on $Y_{i}$ is at most $r$ for every $i$. Then $\psi_{i}$ is an isomorphism near $\lfloor D_{i}\rfloor$ for any $i\gg 0$.
Remark 5.5.
As will be clear from the proof, if Lemma 5.3 holds in dimension $n-1$, then the above lemma holds in dimension $n$.
Proof.
We only outline the main steps since this essentially follows from the same proof of the special termination theorem, see e.g. [Fuj07a]*Proof of Theorem 4.2.1. First note that by the negativity lemma as in Step 1 of loc. cit., after truncating the MMP we may assume that it is an isomorphism at the generic point of every log canonical center of $(Y,D)$. Let $W\subseteq\mathrm{Supp}\lfloor D\rfloor$ be a log canonical center of $(Y,D)$. Let $W_{i}$ be its strict transform on $Y_{i}$ and define $B_{i}$ by adjunction $(K_{Y_{i}}+D_{i})|_{W_{i}}=K_{W_{i}}+B_{i}$. We next prove by induction on $\dim W$ that the induced birational map $\varphi_{i}\colon W_{i}\dashrightarrow W_{i+1}$ induces an isomorphism of pairs $(W_{i},B_{i})\cong(W_{i+1},B_{i+1})$ when $i\gg 0$. The base case $\dim W=0$ is clear, hence we may assume that $\dim W\geq 1$. Let $B=B_{0}$. By Lemma 2.19 and the induction hypothesis, we see that the birational map $W_{i}\dashrightarrow W_{i+1}$ is an isomorphism near $\lfloor B_{i}\rfloor$ for any $i\gg 0$. By [Fuj07a]*Proposition 4.2.14 and Lemma 4.2.15, after another truncation of the MMP we may assume the induced sequence of birational maps
[…]
is a face-contracting $(K_{W}+B)$-MMP. Let $U=W\setminus\lfloor B\rfloor$ and $B_{U}=B|_{U}$. Then we also get a sequence of $(K_{U}+B_{U})$-MMP
[…]
where $U_{i}=W_{i}\setminus\lfloor B_{i}\rfloor$ and $B_{U_{i}}=B_{i}|_{U_{i}}$. Note that $(U,B_{U})$ is klt and $\dim U\leq 4$. Since the index of every $\mathbb{Q}$-Cartier Weil divisor on $Y_{i}$ is at most $r$ for every $i$, by Lemma 2.2 and Lemma 5.3 we deduce that (5.2) becomes an isomorphism when $i\gg 0$. This completes the induction. Since every component of $\lfloor D_{i}\rfloor$ is a log canonical center of $(Y_{i},D_{i})$, the statement now follows from another application of Lemma 2.19. ∎
To apply Lemma 5.4, we need to verify the uniform boundedness of the Cartier index when we lift an MMP to some dlt modification.
Lemma 5.6.
Let $(X,B)$ be a log canonical pair and let $\pi\colon(Y,D)\to(X,B)$ be a crepant birational morphism such that $\mathrm{mult}_{E}D>0$ for all $\pi$-exceptional divisor $E$. Assume that there exists a big $\mathbb{R}$-divisor $\Delta\geq 0$ on $X$ such that $(X,\Delta)$ is klt and $K_{X}+B\sim_{\mathbb{R}}\mu(K_{X}+\Delta)$ for some $\mu>0$. Then there exists a positive integer $r$, such that for any MMP type contraction $\psi\colon(Y,D)\dashrightarrow(Y^{\prime},D^{\prime})$ with $D^{\prime}=\psi_{*}D$, the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor on $Y^{\prime}$ is at most $r$.
Proof.
Since $\Delta$ is big, after possibly replacing $\Delta$ by some other $\mathbb{R}$-divisor $0\leq\Delta_{0}\sim_{\mathbb{R}}\Delta$ we may assume by Lemma 2.18 that $\Delta\geq A$ where $A\geq 0$ is an ample $\mathbb{R}$-divisor. Since $(X,\Delta)$ is klt and $\mathrm{mult}_{E}D>0$ for any $\pi$-exceptional prime divisor $E$, we can choose some $0<\varepsilon\ll 1$ such that
  • the pair $(X,\Delta^{\prime}:=(1-\varepsilon)B+\varepsilon\Delta)$ is klt, and
  • $\mathrm{mult}_{E}G>0$ for any $\pi$-exceptional prime divisor $E$, where the $\mathbb{R}$-divisor $G$ is defined by crepant pullback: $K_{Y}+G=\pi^{*}(K_{X}+\Delta^{\prime})$.
In particular, $\mathrm{Supp}(\pi^{*}A)\subseteq\mathrm{Supp}(G)$ and hence $G$ is big. We have
$K_{Y}+G=(1-\varepsilon)\pi^{*}(K_{X}+B)+\varepsilon\pi^{*}(K_{X}+\Delta)\sim_{\mathbb{R}}(1-\varepsilon+\varepsilon\mu^{-1})(K_{Y}+D).$
Thus as $\psi\colon(Y,D)\dashrightarrow(Y^{\prime},D^{\prime})$ is an MMP type contraction, by Lemma 2.13 it also induces an MMP type contraction $(Y,G)\dashrightarrow(Y,G^{\prime})$ with $G^{\prime}=\psi_{*}G$. We then conclude by Theorem 3.2. ∎
Lemma 5.7.
Let $(X,\Delta)$ be a $\mathbb{Q}$-factorial klt pair of dimension $5$ and let $B$ be an $\mathbb{R}$-divisor on $X$ such that $(X,B)$ is log canonical. Assume that $\Delta$ is big and $K_{X}+B\sim_{\mathbb{R}}\mu(K_{X}+\Delta)$ for some $\mu>0$. Let
[…]
be a $(K_{X}+B)$-MMP. Then $\varphi_{i}$ is an isomorphism near $\mathrm{Nklt}(X_{i},B_{i})$ for any $i\gg 0$.
Proof.
By [Kol13, Corollary 1.36], there exists a $\mathbb{Q}$-factorial dlt pair $(Y,D)$ and a crepant birational morphism $\pi\colon(Y,D)\to(X,B)$ such that every $\pi$-exceptional divisor has coefficient $1$ in $D$. By Lemma 2.15, we can lift (5.3) to a sequence
[…]
such that each $\psi_{i}$ is a sequence of steps of a $(K_{Y_{i}}+D_{i})$-MMP, and there are crepant birational morphisms $\pi_{i}:(Y_{i},D_{i})\to(X_{i},B_{i})$. We claim that $\mathrm{Ex}(\pi_{i})\subseteq\lfloor D_{i}\rfloor$ for all $i$. In fact, since $X_{i}$ is $\mathbb{Q}$-factorial, $\mathrm{Ex}(\pi_{i})$ is of pure codimension one in $Y_{i}$ by [KM98, Corollary 2.63]. In particular, the claim is clear when $i=0$, so we may assume that $i\geq 1$. For any irreducible component $E$ of $\mathrm{Ex}(\pi_{i})$, we have
$a(E,X_{i},\Delta_{i})=-\mathrm{mult}_{E}D_{i}=-\mathrm{mult}_{E}D_{i-1}=a(E,X_{i-1},\Delta_{i-1}),$
so $E$ cannot be the strict transform of some $\varphi_{i-1}$-exceptional divisor (otherwise we get a contradiction to the negativity lemma). It follows that the strict transform of $E$ on $Y_{i-1}$ is contained in $\mathrm{Ex}(\pi_{i-1})$, hence $\mathrm{Ex}(\pi_{i})=(\psi_{i-1})_{*}\mathrm{Ex}(\pi_{i-1})$, and the claim follows by induction on $i$.
By Lemma 5.6, there exists some positive integer $r$ such that the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor on $Y_{i}$ is at most $r$ for all $i$. Thus by Lemma 5.4, we see that $\psi_{i}$ is an isomorphism near $\lfloor D_{i}\rfloor$ for any $i\gg 0$. Since $(Y_{i},D_{i})$ is dlt and $\pi_{i}$ is crepant, combined with the previous claim $\mathrm{Ex}(\pi_{i})\subseteq\lfloor D_{i}\rfloor$ we see that $\lfloor D_{i}\rfloor=\pi_{i}^{-1}(\mathrm{Nklt}(X_{i},B_{i}))$. Therefore, $\varphi_{i}$ is also an isomorphism near $\mathrm{Nklt}(X_{i},B_{i})$ for any $i\gg 0$; otherwise, by the negativity lemma, for some $i\gg 0$ we can find a prime divisor $E$ over $X$ whose center is contained in $\mathrm{Nklt}(X_{i},B_{i})$ such that
$a(E,Y_{i},D_{i})=a(E,X_{i},B_{i})<a(E,X_{i+1},B_{i+1})=a(E,Y_{i+1},D_{i+1}),$
which is a contradiction as the center of $E$ on $Y_{i}$ is contained in $\lfloor D_{i}\rfloor$ and $\psi_{i}$ is an isomorphism around $\lfloor D_{i}\rfloor$. ∎
Proof of Theorem 5.1 when $\dim X=5$.
Let
[…]
be a $(K_{X}+\Delta)$-MMP. Let $D_{i}$ be the strict transform of $D$ on $X_{i}$, and let $t_{0}=\mathrm{lct}(X,\Delta;D)$. Then $(X,\Delta+t_{0}D)$ is log canonical and $K_{X}+\Delta+t_{0}D\sim_{\mathbb{R}}(1+t_{0})(K_{X}+\Delta)$. Since $K_{X}+\Delta\sim_{\mathbb{R}}D$, the MMP is an isomorphism away from $\mathrm{Supp}(D_{i})$ and any $(K_{X}+\Delta)$-MMP is also a $(K_{X}+\Delta+t_{0}D)$-MMP. By Lemma 5.7, after finitely many steps, the MMP is an isomorphism near $\mathrm{Nklt}(X_{i},\Delta_{i}+t_{0}D_{i})$. In other words, after truncation it induces a sequence of $(K_{U}+\Delta_{U})$-MMP where $U=X\setminus\mathrm{Nklt}(X,\Delta+t_{0}D)$ and $\Delta_{U}=\Delta|_{U}$. Now $(U,\Delta_{U}+t_{0}D_{U})$ is klt (where $D_{U}=D|_{U}$) and hence $t_{1}:=\mathrm{lct}(U,\Delta_{U};D_{U})>t_{0}$. Replacing $(X,\Delta)$, $D$ with $(U,\Delta_{U})$, $D_{U}$ and repeat the same argument, we see that the MMP is again an isomorphism near $\mathrm{Nklt}(X_{i},\Delta_{i}+t_{1}D_{i})$ after finitely many steps, and thus we can inductively construct, if the MMP does not terminate, a strictly increasing sequence $t_{0}<t_{1}<\dots$ of log canonical thresholds, where each $t_{i}$ has the form $\mathrm{lct}_{x_{j}}(X_{j},\Delta_{j};D_{j})$ for some $j\in\mathbb{N}$ and some point $x_{j}\in X_{j}$. This contradicts the ACC of log canonical thresholds [HMX-ACC]*Theorem 1.1. It follows that the MMP must terminate. ∎
5.2. Effective termination
We next upgrade the termination result from the previous subsection to an effective one and give some further applications.
Theorem 5.8.
Let $(X,\Delta)$ be a klt pair of dimension $\leq 5$ and $X\to T$ a projective morphism. Assume that at least one of the following holds.
  1. (1)
    $K_{X}+\Delta$ is big over $T$,
  2. (2)
    $K_{X}+\Delta$ is pseudo-effective over $T$, and $\Delta$ is big over $T$,
  3. (3)
    $\dim X\leq 4$ and $\Delta$ is big over $T$.
Then there are finitely many birational maps $\phi_{i}:X\dashrightarrow Y_{i}$ over $T$, $1\leq i\leq m$, such that for any sequences of steps of a $(K_{X}+\Delta)$-MMP $\phi:X\dashrightarrow Y$ over $T$, we have $\phi=\phi_{i}$ for some $i$. In particular, any $(K_{X}+\Delta)$-MMP over $T$ terminates in at most $m$ steps.
This essentially follows from Theorem 5.1 and Kőnig’s lemma: there are only finitely many extremal rays we can possibly contract at each step of the MMP, so if the MMP can have an arbitrary large number of steps, we would be able to construct an MMP that does not terminate, which contradicts Corollary 5.2. We spell out the details in the next two lemmas.
Lemma 5.9.
Let $(X,\Delta)$ be a $\mathbb{Q}$-factorial klt pair and $X\to T$ a projective morphism such that $K_{X}+\Delta$ or $\Delta$ is big over $T$. Then there are only finitely many possible $(K_{X}+\Delta)$-divisorial contractions or flips over $T$.
Proof.
By Lemma 2.18 we may assume that $\Delta=\Delta_{0}+A$, where $\Delta_{0}\geq 0$ and $A\geq 0$ is an ample $\mathbb{R}$-divisor. By the cone theorem [KM98, Theorem 3.7(2)], there are only finitely many $(K_{X}+\Delta)$-negative rays over $T$. In particular, there are only finitely many possible $(K_{X}+\Delta)$-divisorial contractions or flips over $T$. ∎
Lemma 5.10.
Let $(X,\Delta)$ be a $\mathbb{Q}$-factorial klt pair and $X\to T$ a projective morphism such that $K_{X}+\Delta$ or $\Delta$ is big over $T$. Assume that any $(K_{X}+\Delta)$-MMP over $T$ terminates. Then there are finitely many birational maps $\phi_{i}:X\dashrightarrow Y_{i}$ over $T$, $1\leq i\leq m$, such that for any sequences of steps of a $(K_{X}+\Delta)$-MMP $\phi:X\dashrightarrow Y$ over $T$, we have $\phi=\phi_{i}$ for some $i$. In particular, any $(K_{X}+\Delta)$-MMP over $T$ terminates in at most $m$ steps.
Proof.
By Lemma 5.9, there are only finitely many possible $(K_{X}+\Delta)$-divisorial contractions or flips $\phi_{i}:(X_{0},\Delta_{0}):=(X,\Delta)\dashrightarrow(X_{1}^{(j)},\Delta_{1}^{(j)})$, $1\leq j\leq m_{1}$, where $\Delta_{1}^{(j)}$ is the strict transform of $\Delta_{0}$ on $X_{1}^{(j)}$. If the lemma does not hold for $(X,\Delta)$, then the lemma does not hold for $(X_{1}^{(j)},\Delta_{1}^{(j)})$ for some $j$. Let $(X_{1},\Delta_{1}):=(X_{1}^{(j)},\Delta_{1}^{(j)})$ and repeat the same argument with $(X_{1},\Delta_{1})$ in place of $(X_{0},\Delta_{0})$. Inductively, we may construct $(X_{i},\Delta_{i})$, $i\geq 1$, such that
$(X_{0},\Delta_{0})\dashrightarrow(X_{1},\Delta_{1})\dashrightarrow\cdots\dashrightarrow(X_{i},\Delta_{i})\dashrightarrow(X_{i+1},\Delta_{i+1})\dashrightarrow\cdots$
is an infinite sequence of $(K_{X}+\Delta)$-MMP over $T$, and the lemma does not hold for each $(X_{i},\Delta_{i})$. This is impossible as any $(K_{X}+\Delta)$-MMP over $T$ terminates. ∎
Proof of Theorem 5.8.
This follows from Corollary 5.2 and Lemma 5.10. ∎
Proof of Theorem 1.3.
This follows from Theorem 5.8.∎
As another application of the termination result in this section, we prove the strong Sarkisov program conjecture ([He23]*Conjecture 3.2) in dimension $5$:
Corollary 5.11.
The strong Sarkisov program holds in dimension $5$.
We refer to [He23] and the reference therein for more details of the strong Sarkisov program.
Proof.
It follows from Corollary 5.2 and [He23]*Corollary 3.5. ∎
6. Explicit termination
In this section, we give explicit bounds (see Corollary 6.9, Theorems 6.14, 6.15, and 6.16) on the number of steps in various types of MMP in dimension at most $4$. This improves Theorem 5.8, which only gives the existence of an inexplicit upper bound.
6.1. Invariants
Before we state the result, we need to introduce a few invariants that we shall use to formulate the explicit bound and recall some of their basic properties.
Definition 6.1 (Picard number).
Let $X$ be a smooth variety. Let $j\colon X\hookrightarrow\overline{X}$ be a smooth projective compactification. We set $N^{1}(X):=\mathrm{Pic}(X)/j^{*}\mathrm{Pic}^{0}(\overline{X})$ and define the Picard number of $X$ as
$\rho(X):=\dim N^{1}(X)_{\mathbb{R}}.$
This is well-defined since $j^{*}\mathrm{Pic}^{0}(\overline{X})$ is independent of the compactification. In general, for any normal variety $X$, we define $\rho(X):=\rho(U)$ where $U$ is the smooth locus of $X$.
Lemma 6.2.
Let $\pi\colon X\rightarrow Z$ be a projective birational morphism of normal varieties. Then $\rho(X)=\rho(Z)+\ell$, where $\ell$ is the number of $\pi$-exceptional divisors.
Proof.
If $X$ is smooth, this follows from [AHK07, Lemma 1.6(1)]. The general case can be reduced to this by considering a common resolution of $X$ and $Z$. ∎
Definition 6.3 (Algebraic Betti number).
Let $X$ be a variety and let $m\in\mathbb{N}$. We denote by $H^{\mathrm{alg}}_{2m}(X)$ the subspace in the singular homology group $H_{2m}(X,\mathbb{R})$ generated by the the classes of $m$-dimensional complete algebraic subvarieties. We denote by $h^{\mathrm{alg}}_{2m}(X)$ the dimension of $H^{\mathrm{alg}}_{2m}(X)$.
Lemma 6.4.
Let $m\in\mathbb{N}$ and let $\pi\colon X\rightarrow Z$ be a projective birational morphism of normal varieties such that $\dim\mathrm{Ex}(\pi)\leq m$. Then
(6.1) $h^{\mathrm{alg}}_{2m}(Z)+\min\{1,\ell\}\leq h^{\mathrm{alg}}_{2m}(X)\leq h^{\mathrm{alg}}_{2m}(Z)+\ell$
where $\ell$ is the number of $m$-dimensional complete irreducible components of $\mathrm{Ex}(\pi)$.
Proof.
For simplicity, we abbreviate all singular homology groups $H_{*}(-,\mathbb{R})$ as $H_{*}(-)$ in this proof.
As $\pi$ is projective, every complete subvariety of $Z$ is the image of some complete subvariety of $X$ of the same dimension. Thus the natural homomorphism $\pi_{*}\colon H_{2m}(X)\to H_{2m}(Z)$ induces a surjection $H^{\mathrm{alg}}_{2m}(X)\twoheadrightarrow H^{\mathrm{alg}}_{2m}(Z)$. In particular, $h^{\mathrm{alg}}_{2m}(Z)\leq h^{\mathrm{alg}}_{2m}(X)$. For any complete subvariety $V\subseteq X$ of dimension $m$ and any projective compactification $\overline{X}$ of $X$, we have $[V]\neq 0\in H^{\mathrm{alg}}_{2m}(\overline{X})$, hence $[V]\neq 0\in H^{\mathrm{alg}}_{2m}(X)$ as well. If $\pi$ contracts at least one complete subvariety of dimension $m$ (i.e. $\ell\geq 1$), then it follows that $\ker(\pi_{*})\neq 0$ and we get $h^{\mathrm{alg}}_{2m}(Z)\leq h^{\mathrm{alg}}_{2m}(X)-1$. This proves the first inequality in (6.1).
Let $W=\mathrm{Ex}(\pi)$. By assumption, $\dim\pi(W)\leq m-1$, thus $H_{i}(\pi(W))=0$ for any $i>2m-2$. By the long exact sequence of homology we deduce that $H_{2m}(Z)\cong H_{2m}(Z,\pi(W))$ where $H_{2m}(Z,\pi(W))$ is the relative homology. Since $X\setminus W\cong Z\setminus\pi(W)$, by excision we also have $H_{2m}(Z,\pi(W))\cong H_{2m}(X,W)$. Combining with the long exact sequence of homology we see that
$H_{2m}(W)\to H_{2m}(X)\to H_{2m}(X,W)\cong H_{2m}(Z)$
is exact. Note that $\dim H_{2m}(W)=\ell$ by the Mayer–Vietoris sequence and the fact that for any variety $V$ of dimension at most $m$, we have
$\dim H_{2m}(V)=\begin{cases}1,&\mbox{if $V$ is complete and }\dim V=m,\\ 0,&\mbox{otherwise}.\end{cases}$
It follows that $\ker(\pi_{*})\leq\ell$ and hence $h^{\mathrm{alg}}_{2m}(X)\leq h^{\mathrm{alg}}_{2m}(Z)+\ell$. This proves the other inequality in (6.1). ∎
Definition 6.5.
Let $(X,\Delta)$ be a klt pair of dimension $n$ and let $\widetilde{B}$ be the normalization of $B:=\mathrm{Supp}(\Delta)$. We define
(6.2) $\rho(X,\Delta):=\rho(X)+\rho(\widetilde{B})+h^{\mathrm{alg}}_{2n-4}(X).$
Following [AHK07], we call an exceptional divisor $E$ over $X$ with $a(E,X,\Delta)<1$ an echo of $(X,\Delta)$ if its center $c_{X}(E)$ is a codimension one point in the smooth locus of $B$. We also define
(6.3) $d(X,\Delta):=\#\left\{\begin{tabular}[]{c}\mbox{Exceptional prime divisor $E$ over $X$ with}\\ \mbox{$a(E,X,\Delta)<1$ that is not an echo of $(X,\Delta)$}\end{tabular}\right\}.$
Note that if $(X,\Delta)$ is terminal, then $d(X,\Delta)<+\infty$ by [K+-flip-abundance]*Lemma (4.14.2.1) or [AHK07]*Lemma 1.5. Finally, we define
(6.4) $s(X,\Delta):=\min_{(Y,D)}\{\rho(Y,D)+d(Y,D)\}$
where the minimum runs over all terminal pairs $(Y,D)$ with a projective birational morphism $\pi\colon(Y,D)\to(X,\Delta)$ such that $K_{Y}+D=\pi^{*}(K_{X}+\Delta)+E$ for some $\pi$-exceptional $\mathbb{R}$-divisor $E\geq 0$.
Since the last invariant $s(X,\Delta)$ will play a major role in controlling the number of steps in the MMP, we give an explicit upper bound of $s(X,\Delta)$ at least for log smooth pairs. We also bound the invariant $e_{+}(X,\Delta)$ defined after (3.1).
Lemma 6.6.
Let $(X,\Delta=\sum b_{i}B_{i})$ be a log smooth pair of dimension $n$. Let $N$ be a positive integer. Assume that $b_{i}\leq 1-\frac{1}{N}$ for all $i$, $(X,\Delta)$ has at most $N$ strata, each stratum has Picard number at most $N$, and $h_{2n-4}^{\mathrm{alg}}(X)\leq N$. Then
$e_{+}(X,\Delta)<N^{n+1}\quad\mathrm{and}\quad s(X,\Delta)<3N^{2n+3}.$
Proof.
The result is clear when $N=1$: in this case $\Delta=0$ and $\rho(X)=h_{2n-4}^{\mathrm{alg}}(X)=1$ by assumption, which gives $e_{+}(X,\Delta)=d(X,\Delta)=0$ (as $(X,\Delta)$ is already terminal) and $s(X,\Delta)=2$ by definition. Thus we may assume that $N\geq 2$.
A $\mathbb{Q}$-factorial terminalization $(Y,D)$ of $(X,\Delta)$ can be constructed by repeatedly blowing up the center of some prime exceptional divisor $E$ over $X$ with $a(E,X,\Delta)\leq 0$ until all such divisors have been extracted. Note that the center of such a divisor $E$ is necessarily a stratum at each step (this follows from e.g. the calculations in [KM98, Corollary 2.31]). For each stratum $W\subseteq B_{J}$, every divisor $E$ with center $W$ and $a(E,X,\Delta)\leq 0$ can also be obtained by a weighted blowup with some weights $m_{j}\in\mathbb{N}_{+}$ along $B_{j}$. Since
$a(E,X,\Delta)=-1+\sum_{j\in J}(1-b_{j})m_{j}$
for the exceptional divisor $E$ of such a weighted blowup, and $b_{i}\leq 1-\frac{1}{N}$ by assumption, we see that $m_{j}\leq N$ and hence for each stratum $W$ there are no more than $N^{n}$ divisors $E$ with center $W$ and $a(E,X,\Delta)\leq 0$. As there at no more than $N$ strata and one of them is $X$, we deduce that the $\mathbb{R}$-divisor $D$ on the terminalization has less than $N^{n+1}$ components, and we only need to blowup less than $N^{n+1}$ times to obtain the terminalization $Y$. In particular, $e_{+}(X,\Delta)<N^{n+1}$.
At the initial stage (i.e. on $X$), the intersection of any two strata has at most $N$ connected components, and this number does not increase under the blowup process. This implies that after each stratum blowup, the Picard number of a stratum can increase at most $N$, hence after $\ell$ blowups, the Picard number of any stratum is at most $(\ell+1)N$. It follows that every stratum of $(Y,D)$ has Picard number at most $N^{n+2}$; in particular, $\rho(Y)\leq N^{n+2}$, and if $\widetilde{B}_{Y}$ is the normalization of $B_{Y}:=\mathrm{Supp}(D)$, then
$\rho(\widetilde{B}_{Y})\leq N^{n+2}(N^{n+1}-1)$
as $D$ has less than $N^{n+1}$ components. On the other hand, by [Voi-Hodge-book-I, Theorem 7.31], if we blowup a stratum of Picard number $\rho$, then $h_{2n-4}^{\mathrm{alg}}$ increases by at most $\rho+1$. By assumption, $h_{2n-4}^{\mathrm{alg}}(X)\leq N$, hence
$h_{2n-4}^{\mathrm{alg}}(Y)\leq h_{2n-4}^{\mathrm{alg}}(X)+(N+1)+(2N+1)+\dots+((N^{n+1}-1)N+1)<N^{2n+3}.$
Putting the above together we see that
$\rho(Y,D)=\rho(Y)+\rho(\widetilde{B}_{Y})+h_{2n-4}^{\mathrm{alg}}(Y)<N^{n+2}+N^{n+2}(N^{N+1}-1)+N^{2n+3}=2N^{2n+3}.$
Next observe that by [Kol-flops, Corollary 3.2(iii)], every divisor $E$ over $Y$ with $a(E,Y,D)=a(E,X,\Delta)<1$ that is not an echo of $(Y,D)$ is necessarily obtained by a weighted blowup along some stratum of $(X,\Delta)$, as the assumptions of loc. cit. is satisfied after blowing up some strata. For each fixed stratum $W\subseteq B_{J}$, if $m_{j}$ ($j\in J$) are the weights along $B_{j}$, then
$\sum_{j\in J}(1-b_{j})m_{j}=1+a(E,Y,D)<2.$
As $b_{i}\leq 1-\frac{1}{N}$, we have $m_{j}<2N$, hence there are at most $(2N)^{n}$ divisors $E$ with $a(E,X,\Delta)<1$ that are centered at any given stratum. Since there are no more than $N$ strata, we see that $d(Y,D)\leq 2^{n}N^{n+1}$ and (since $N\geq 2$)
$s(X,\Delta)\leq\rho(Y,D)+d(Y,D)<2N^{2n+3}+2^{n}N^{n+1}<3N^{2n+3},$
which concludes the proof. ∎
6.2. Terminal case
Next we prove an explicit bound on the number of steps in a special type of MMP on terminal pairs, which includes all terminal flips in dimension $3$ or $4$.
We start with some general setup. For a terminal pair $(X,\Delta)$ of dimension $n$, write $\Delta=\sum_{i=1}^{k}b_{i}B_{i}$ where $10<b_{k}<\dots<b_{1}<$ 1 and the $B_{i}$’s are reduced divisors without common components, and let $\widetilde{B}_{i}$ be the normalization of $B_{i}$. Set $b_{0}:=1$, $b_{k+1}:=0$, and let
(6.5) $S:=[0,1]\cap\sum_{i=0}^{k}b_{i}\mathbb{N}\subseteq\mathbb{R}.$
In particular, $S=\{0,1\}$ if $\Delta=0$. We define a sequence of topological invariants as follows.
(6.6) $\rho_{i}:=\rho_{i}(X,\Delta):=\begin{cases}\rho(X),&\mbox{if }i=0,\\ \rho(\widetilde{B}_{i}),&\mbox{if }1\leq i\leq k,\mbox{ and}\\ h^{\mathrm{alg}}_{2n-4}(X),&\mbox{if }i=k+1.\end{cases}$
By definition, $\rho(X,\Delta)=\sum_{i=0}^{k+1}\rho_{i}(X,\Delta)$ (cf. (6.2)). For each $i=1,\dots,k+1$, we also define the $i$-th (weighted) difficulty invariant by
(6.7) $d_{i}:=d_{i}(X,\Delta):=\sum_{\eta\in S,\,\eta\geq b_{i}}\#\left\{\begin{tabular}[]{c}\mbox{Exceptional divisors $E$ over $X$ with}\\ \mbox{$a(E,X,\Delta)<1-\eta$ whose center $c_{X}(E)$ is not a}\\ \mbox{codimension one point in the smooth locus of $\mathrm{Supp}(\Delta)$}\end{tabular}\right\}.$
Clearly $d_{1}\leq d_{2}\leq\cdots\leq d_{k+1}\leq|S|\cdot d(X,\Delta)$, where $d(X,\Delta)$ is defined in (6.3). The importance of these invariants comes from the following result.
Theorem 6.7.
Let $(X,\Delta)$ be a $\mathbb{Q}$-factorial terminal pair of dimension $n\in\{3,4\}$. Then in an $(K_{X}+\Delta)$-MMP that does not contract any component of $\Delta$, the sequence $(\rho_{0},d_{1},\rho_{1},d_{2},\dots,d_{k+1},\rho_{k+1})$ is strictly decreasing in lexicographic order.
Proof.
This follows from the proof [Fuj04, Fuj05] that such MMP terminates. ∎
This suggests that if we define an “$M$-adic difficulty” by
(6.8) $\delta_{M}(X,\Delta)$ $:=\sum_{i=0}^{k+1}M^{2k+2-2i}\rho_{i}(X,\Delta)+\sum_{i=1}^{k+1}M^{2k+3-2i}d_{i}(X,\Delta)$
$=M^{2k+2}\rho_{0}+M^{2k+1}d_{1}+M^{2k}\rho_{1}+\dots+Md_{k+1}+\rho_{k+1},$
where $M$ is a positive integer, then it is reasonable to expect that $\delta_{M}(X,\Delta)$ is also strictly decreasing in the MMP in Theorem 6.7 when $M\gg 0$. The main result of this subsection is an explicit form of such a statement.
Theorem 6.8.
Let $(X,\Delta=\sum_{i=1}^{k}b_{i}B_{i})$ be a $\mathbb{Q}$-factorial terminal pair of dimension $n$ with $b_{k}<\dots<b_{1}$. Set $b_{0}=1$, $b_{k+1}=0$ and $S=[0,1]\cap\sum_{i=0}^{k}b_{i}\mathbb{N}$ as above. Then for any
$M\geq\left(2+\left\lceil\frac{1}{b_{k}}\right\rceil\cdot\left\lceil\frac{1}{1-b_{1}}\right\rceil\right)\cdot|S|,$
the invariant $\delta_{M}(X,\Delta)$ is non-increasing in any $(K_{X}+\Delta)$-MMP that does not contract any component of $\Delta$. Moreover, it strictly decreases in the MMP if $n\in\{3,4\}$.
Corollary 6.9.
Let $(X,\Delta=\sum_{i=1}^{k}b_{i}B_{i})$ be a $\mathbb{Q}$-factorial terminal pair of dimension $n\in\{3,4\}$ with $b_{k}<\dots<b_{1}$. Let $b_{0}=1$, $b_{k+1}=0$ and $S=[0,1]\cap\sum_{i=0}^{k}b_{i}\mathbb{N}$. Then every $(K_{X}+\Delta)$-MMP that does not contract any component of $\Delta$ terminates after at most
$\left(2+\left\lceil\frac{1}{b_{k}}\right\rceil\cdot\left\lceil\frac{1}{1-b_{1}}\right\rceil\right)^{2k+2}|S|^{2k+2}(\rho(X,\Delta)+d(X,\Delta))$
steps, where $\rho(X,\Delta)$ and $d(X,\Delta)$ are defined in Definition 6.5.
The remaining part of this subsection is devoted to proving the two results above.
Proof of Theorem 6.8.
By induction, it suffices to prove that $\delta_{M}(X,\Delta)\geq\delta_{M}(X^{\prime},\Delta^{\prime})$ for any step $\varphi\colon(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ of the $(K_{X}+\Delta)$-MMP which does not contract any component of $\Delta$, and that the inequality is strict if $n\in\{3,4\}$. Let $\rho^{\prime}_{i}:=\rho_{i}(X^{\prime},\Delta^{\prime})$ and $d^{\prime}_{i}:=d_{i}(X^{\prime},\Delta^{\prime})$.
First assume that $\varphi$ is a flip. Let $f\colon X\to Z$ be the flipping contraction and $f^{\prime}\colon X^{\prime}\to Z$ the flip. Let $B^{\prime}_{i}$ be the strict transform of $B_{i}$ on $X^{\prime}$ and $\widetilde{B}^{\prime}_{i}$ its normalization, and let $\bar{B}_{i}$ be the normalization of $f(B_{i})=f^{\prime}(B^{\prime}_{i})\subseteq Z$. For ease of notation we also denote the induced morphism $\widetilde{B}_{i}\to\bar{B}_{i}$ and $\widetilde{B}^{\prime}_{i}\to\bar{B}_{i}$ by $f$ and $f^{\prime}$.
[…]
We have $\rho_{0}=\rho^{\prime}_{0}$ since $\varphi$ is small. To compare the other invariants $\rho^{\prime}_{i},d^{\prime}_{i}$ with $\rho_{i},d_{i}$, we introduce two intermediate sequences as follows.
(6.9) $\bar{\rho}_{i}:=\begin{cases}\rho(\bar{B}_{i}),&\mbox{if }1\leq i\leq k,\\ h^{\mathrm{alg}}_{2n-4}(Z),&\mbox{if }i=k+1.\end{cases}$
(6.10) $\bar{d}_{i}:=\sum_{\eta\in S,\eta\geq b_{i}}\#\left\{\begin{tabular}[]{c}\mbox{Exceptional divisors $E$ over $X^{\prime}$ with}\\ \mbox{$a(E,X^{\prime},\Delta^{\prime})<1-\eta$ whose center $c_{X}(E)$ on $X$ is not a}\\ \mbox{codimension one point in the smooth locus of $\mathrm{Supp}(\Delta)$}\end{tabular}\right\}.$
Note that in the definition of $\bar{d}_{i}$, the discrepancy is computed with respect to $(X^{\prime},\Delta^{\prime})$ while the center is taken on $X$. It is clear that $\rho_{i}\geq\bar{\rho}_{i}$ for all $i=1,\dots,k+1$. Since $a(E,X,\Delta)\leq a(E,X^{\prime},\Delta^{\prime})$ for all prime divisors $E$ over $X$ by the negativity lemma, we also have $d_{i}\geq\bar{d}_{i}$ for all $i=1,\dots,k+1$. We claim that
(6.11) $d^{\prime}_{i}-\bar{d}_{i}\leq(M-1)\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j}),\,\mbox{ and }$
(6.12) $\rho^{\prime}_{i}-\bar{\rho}_{i}\leq(M-1)\left(d_{i}-\bar{d}_{i}+\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})\right).$
Roughly speaking, the claim says that the sequence $(\rho_{0},d_{1},\rho_{1},d_{2},\dots,d_{k+1},\rho_{k+1})$ behaves like the digits of $M$-adic numbers: each time a higher digit drops by $1$, the lower digits increase by at most $M-1$. Taking (6.11) and (6.12) for granted, Lemma 6.10 then implies that $\delta_{M}(X,\Delta)\geq\delta_{M}(X^{\prime},\Delta^{\prime})$. Moreover, equality holds if and only if $d^{\prime}_{i}=d_{i}$ for all $i=1,\dots,k+1$ and $\rho^{\prime}_{j}=\rho_{j}$ for all $j=0,1,\dots,k+1$. This is not possible in dimension $3$ and $4$ by Theorem 6.7, thus $\delta_{M}(X,\Delta)>\delta_{M}(X^{\prime},\Delta^{\prime})$ when $n\in\{3,4\}$.
We now focus on the proof of (6.11) and (6.12).
Proof of (6.11).
Let $E$ be an exceptional divisor that has more contributions to $d^{\prime}_{i}$ than to $\bar{d}_{i}$. Let $\xi=c_{X}(E)$ (resp. $\xi^{\prime}=c_{X^{\prime}}(E)$) be its center on $X$ (resp. $X^{\prime}$). Then $\xi^{\prime}$ is not a codimension one point in the smooth locus of $\mathrm{Supp}(\Delta^{\prime})$ and $a(E,X^{\prime},\Delta^{\prime})<1-b_{i}$, otherwise $E$ has no contribution to $d^{\prime}_{i}$. On the other hand, since $a(E,X,\Delta)\leq a(E,X^{\prime},\Delta^{\prime})$, we see that $\xi$ must be a codimension one point in the smooth locus of $\mathrm{Supp}(\Delta)$, otherwise it cannot have more contributions to $d^{\prime}_{i}$. Thus $\xi$ is contained in the exceptional locus of $f$. We also have $a(E,X,\Delta)\leq a(E,X^{\prime},\Delta^{\prime})<1-b_{i}$, hence $\xi\in B_{j}$ for some $j<i$ by e.g. [AHK07]*Lemma 1.5. By Lemma 6.2, there are at most $\rho_{j}-\bar{\rho}_{j}$ components of dimension $n-2$ in the exceptional locus of $f|_{B_{j}}$, and by [AHK07]*Lemma 1.5, over the generic point of every such component that is contained in the smooth locus of $B_{j}$ there are at most $\left\lceil\frac{1}{1-b_{j}}\right\rceil\leq\left\lceil\frac{1}{1-b_{1}}\right\rceil$ divisors $E$ with $a(E,X,\Delta)<1$. It follows that there are at most
$\left\lceil\frac{1}{1-b_{1}}\right\rceil\cdot\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})$
exceptional divisors $E$ that contribute positively to $d^{\prime}_{i}-\bar{d}_{i}$. Note that every such divisor $E$ is counted at most $|S|$ times in $d^{\prime}_{i}$. Thus
$d^{\prime}_{i}-\bar{d}_{i}\leq|S|\cdot\left\lceil\frac{1}{1-b_{1}}\right\rceil\cdot\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})\leq(M-1)\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j}),$
which gives (6.11). ∎
Proof of (6.12).
For notational convenience, set $\widetilde{B}^{\prime}_{k+1}=B^{\prime}_{k+1}=X^{\prime}$. By Lemmas 6.2 and 6.4, we know that $\rho^{\prime}_{i}-\bar{\rho}_{i}$ is at most the number of irreducible components of dimension $n-2$ in the exceptional locus of $f^{\prime}|_{\widetilde{B}^{\prime}_{i}}$. We first count the number of irreducible components of dimension $n-2$ in the exceptional locus of $f^{\prime}|_{B^{\prime}_{i}}$ (that is, before passing to the normalization). Let $W$ be one such component. Note that $(X^{\prime},\Delta^{\prime})$ is terminal, hence $X^{\prime}$ is smooth at the generic point of $W$. Let $E$ be the exceptional divisor obtained by blowing up $W$ near the generic point of $W$ on $X^{\prime}$. Then by [Fuj04]*Lemma 2.2, we have
(6.13) $a(E,X,\Delta)<a(E,X^{\prime},\Delta^{\prime})=1-\eta$
for some $b_{i}\leq\eta\in S$.
Let $\xi=c_{X}(E)$ be the center of $E$ on $X$. Then $\xi\in\mathrm{Ex}(f)$ since $c_{X^{\prime}}(E)\in W\subseteq\mathrm{Ex}(f^{\prime})$. If $\xi$ is not a codimension one point in the smooth locus of $\mathrm{Supp}(\Delta)$, then (6.13) implies that $E$ has more contributions to $d_{i}$ than to $\bar{d}_{i}$, and there are at most $d_{i}-\bar{d}_{i}$ such divisors. If $\xi$ is a codimension one point in the smooth locus of $\mathrm{Supp}(\Delta)$, then by (6.13) and [AHK07]*Lemma 1.5 we see that $\xi\in B_{j}$ for some $1\leq j<i$. Since $\xi\in\mathrm{Ex}(f)$, it is also the generic point of a component $V$ of dimension $n-2$ in the exceptional locus of $f$. Combined with Lemma 6.2, we deduce that there are at most $\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})$ such components $V$. Moreover, once $V$ is fixed, there are at most $\left\lceil\frac{1}{1-b_{j}}\right\rceil\leq\left\lceil\frac{1}{1-b_{1}}\right\rceil$ divisors $E$ centered at the generic point of $V$ such that $a(E,X,\Delta)<1$. Putting these together, and since $W$ is uniquely determined by $E$ as its center, we see that there are at most
$d_{i}-\bar{d}_{i}+\left\lceil\frac{1}{1-b_{1}}\right\rceil\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})$
components $W$ of dimension $n-2$ in the exceptional locus of $f^{\prime}|_{B^{\prime}_{i}}$. When $i=k+1$ this already gives an upper bound of $\rho^{\prime}_{i}-\bar{\rho}_{i}$ by Lemma 6.4. If $1\leq i\leq k$, then since $(X^{\prime},\Delta^{\prime})$ is terminal, we have $\mathrm{mult}_{W}\Delta^{\prime}<1$ and hence $\mathrm{mult}_{W}B^{\prime}_{i}<\frac{1}{b_{i}}$. It follows that the preimage of $W$ in the normalization $\widetilde{B}^{\prime}_{i}$ has at most $\left\lceil\frac{1}{b_{i}}\right\rceil\leq\left\lceil\frac{1}{b_{k}}\right\rceil$ components. In other words, the contribution to $\rho^{\prime}_{i}-\bar{\rho}_{i}$ by each component $W$ is at most $\left\lceil\frac{1}{b_{k}}\right\rceil$. Therefore,
$\rho^{\prime}_{i}-\bar{\rho}_{i}$ $\leq\left\lceil\frac{1}{b_{k}}\right\rceil\left(d_{i}-\bar{d}_{i}+\left\lceil\frac{1}{1-b_{1}}\right\rceil\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})\right)$
$\leq(M-1)\left(d_{i}-\bar{d}_{i}+\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})\right)$
as desired. ∎
As discussed before, this concludes the proof of the theorem when $\varphi$ is a flip.
Next assume that $\varphi\colon(X,\Delta)\to(X^{\prime},\Delta^{\prime})$ is a divisorial contraction. Then clearly $\rho^{\prime}_{0}=\rho_{0}-1$ and $\rho^{\prime}_{i}\leq\rho_{i}$ for all $i\geq 1$. Since none of the components of $\Delta$ is contracted, the pair $(X^{\prime},\Delta^{\prime})$ remains terminal. Set $\bar{\rho}_{i}:=\rho^{\prime}_{i}$ and $\bar{d}_{i}:=d_{i}$. Similar to the previous case, we claim that
(6.14) $d^{\prime}_{i}-\bar{d}_{i}\leq(M-1)\sum_{0\leq j<i}(\rho_{j}-\bar{\rho}_{j}),$
which implies, again by Theorem 6.7 and Lemma 6.10, that $\delta_{M}(X,\Delta)\geq\delta_{M}(X^{\prime},\Delta^{\prime})$, with strict inequality when $n\in\{3,4\}$.
To see this claim, let $F$ be the (unique) exceptional divisor of $\varphi$ and let $E$ be an exceptional divisor over $X^{\prime}$ that contributes more to $d^{\prime}_{i}$ than to $\bar{d}_{i}=d_{i}$. Let $\xi=c_{X}(E)$ (resp. $\xi^{\prime}=c_{X^{\prime}}(E)$) be the center of $E$ on $X$ (resp. $X^{\prime}$). Then $\xi^{\prime}$ is not a codimension one point of the smooth locus of $\mathrm{Supp}(\Delta^{\prime})$, and $a(E,X^{\prime},\Delta^{\prime})<1-b_{i}$. Since $a(E,X,\Delta)\leq a(E,X^{\prime},\Delta^{\prime})$, in order for the divisor $E$ to contribute less to $d_{i}$, there are only two possibilities: either $E=F$ (so it is not exceptional over $X$), or $E$ is exceptional over $X$ but its center $\xi$ is a codimension one point of the smooth locus of $\mathrm{Supp}(\Delta)$. In both cases we necessarily have $\xi\in\mathrm{Ex}(\varphi)=F$. Let us analyze the latter situation more carefully.
Since $a(E,X,\Delta)\leq a(E,X^{\prime},\Delta^{\prime})<1-b_{i}$, by [AHK07]*Lemma 1.5 we know that $\xi\in B_{j}$ for some $1\leq j<i$. Note that $\xi^{\prime}=\varphi(\xi)$. If $\xi^{\prime}$ is a point of codimension at least $3$ in $X^{\prime}$, then $W=\overline{\{\xi\}}$ is a component of dimension $n-2$ in the exceptional locus of $\varphi|_{B_{j}}$. There are at most $\rho_{j}-\rho^{\prime}_{j}=\rho_{j}-\bar{\rho}_{j}$ such components by Lemma 6.2. Once $W$ is fixed, by [AHK07]*Lemma 1.5, there are at most $\left\lceil\frac{1}{1-b_{j}}\right\rceil\leq\left\lceil\frac{1}{1-b_{1}}\right\rceil$ exceptional divisors $E$ with center $W$ and $a(E,X,\Delta)<1$. As $j$ varies these give at most
$\left\lceil\frac{1}{1-b_{1}}\right\rceil\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})$
candidates of the divisor $E$ in total. If $\xi^{\prime}$ has codimension $2$ in $X^{\prime}$, then $\xi^{\prime}$ is the generic point of $\varphi(F)$ (because $\xi\in F$ and $\varphi(F)$ has codimension at least $2$ in $X$) and $\varphi|_{B_{j}}$ is finite around $\xi$. As $(X^{\prime},\Delta^{\prime})$ is terminal, we get $\mathrm{mult}_{\xi^{\prime}}\Delta^{\prime}<1$ and thus $\mathrm{mult}_{\xi^{\prime}}(B^{\prime}_{1}+\dots+B^{\prime}_{j})<\frac{1}{b_{j}}\leq\frac{1}{b_{k}}$. It follows that the preimage of $\xi^{\prime}$ in the normalization $\widetilde{B}_{1}\cup\dots\cup\widetilde{B}_{j}$ has at most $\left\lceil\frac{1}{b_{k}}\right\rceil$ components, giving rise to at most $\left\lceil\frac{1}{b_{k}}\right\rceil$ possibilities of $\xi$. For each such choice of $\xi$, there are at most $\left\lceil\frac{1}{1-b_{j}}\right\rceil\leq\left\lceil\frac{1}{1-b_{1}}\right\rceil$ divisors $E$ with center $\xi$ such that $a(E,X,\Delta)<1$ (by [AHK07]*Lemma 1.5 as before). Putting these together, we conclude that there are at most
$\left\lceil\frac{1}{1-b_{1}}\right\rceil\cdot\left\lceil\frac{1}{b_{k}}\right\rceil+\left\lceil\frac{1}{1-b_{1}}\right\rceil\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})=\left\lceil\frac{1}{1-b_{1}}\right\rceil\cdot\left(\left\lceil\frac{1}{b_{k}}\right\rceil+\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})\right)$
exceptional divisors over $X$ that contribute more to $d^{\prime}_{i}$ than to $d_{i}$. Recall that $\rho_{0}-\rho^{\prime}_{0}=1$ and there is also the option that $E=F$. As every divisor $E$ is counted at most $|S|$ times in $d^{\prime}_{i}$, we finally deduce that
$d^{\prime}_{i}-\bar{d}_{i}\leq|S|\cdot\left(1+\left\lceil\frac{1}{1-b_{1}}\right\rceil\cdot\left(\left\lceil\frac{1}{b_{k}}\right\rceil+\sum_{1\leq j<i}(\rho_{j}-\bar{\rho}_{j})\right)\right)\leq(M-1)\sum_{0\leq j<i}(\rho_{j}-\bar{\rho}_{j}).$
This proves the claim and hence of the theorem as well. ∎
The following elementary lemma has been used in the proof above.
Lemma 6.10.
Let $a_{i},b_{i},c_{i}\in\mathbb{R}\,\,(i=0,\dots,m)$ and $M>1$ be such that $a_{i}\geq b_{i}$ and $c_{i}-b_{i}\leq(M-1)\sum_{0\leq j<i}(a_{j}-b_{j})$ for all $i$. Then
$\sum_{i=0}^{m}M^{-i}c_{i}\leq\sum_{i=0}^{m}M^{-i}a_{i}$
with equality if and only if $a_{i}=b_{i}=c_{i}$ for all $i$.
In the applications above, the sequence $a_{i}$ (resp. $b_{i}$, resp. $c_{i}$) represents the sequence $(\rho_{0},d_{1},\dots,d_{k+1},\rho_{k+1})$ (resp. $(\bar{\rho}_{0},\bar{d}_{1},\dots,\bar{d}_{k+1},\bar{\rho}_{k+1})$, resp. $(\rho^{\prime}_{0},d^{\prime}_{1},\dots,d^{\prime}_{k+1},\rho^{\prime}_{k+1})$).
Proof.
We may assume that $b_{i}=0$ by replacing $a_{i}$ (resp. $c_{i}$) with $a_{i}-b_{i}$ (resp. $c_{i}-b_{i}$). The assumption then becomes $a_{i}\geq 0$ and $c_{i}\leq(M-1)\sum_{0\leq j<i}a_{j}$. It follows that $M^{-i}c_{i}\leq(1-\frac{1}{M})\sum_{0\leq j<i}M^{1-i}a_{j}$, hence
$\sum_{i=0}^{m}M^{-i}c_{i}$ $\leq\left(1-\frac{1}{M}\right)\sum_{i=0}^{m}\sum_{0\leq j<i}M^{1-i}a_{j}\leq\left(1-\frac{1}{M}\right)\sum_{j=0}^{m-1}\sum_{i=j}^{+\infty}M^{-i}a_{j}$
$=\sum_{j=0}^{m-1}M^{-j}a_{j}\leq\sum_{j=0}^{m}M^{-j}a_{j}.$
The second and the last equality hold if and only if $a_{i}=0$ for all $i$. If this is the case, then the first equality holds if and only if $c_{i}=(M-1)\sum_{0\leq j<i}a_{j}=0$ for all $i$. ∎
To prove Corollary 6.9, we need another elementary statement.
Lemma 6.11.
We have $\rho(X,\Delta)+d(X,\Delta)\leq\delta_{M}(X,\Delta)$. Assume that $M\geq|S|+1$. Then we also have
(6.15) $\delta_{M}(X,\Delta)\leq M^{2k+2}(\rho(X,\Delta)+d(X,\Delta)).$
Proof.
The first inequality is clear from the definition (6.8) of $\delta_{M}(X,\Delta)$. If $M\geq|S|+1$, then as $d_{i}\leq|S|\cdot d(X.\Delta)$ for all $i$, from (6.8) we also have
$\delta_{M}(X,\Delta)$ $\leq M^{2k+2}\rho(X,\Delta)+(M^{2k+1}+M^{2k-1}+\dots+M)\cdot|S|\cdot d(X,\Delta)$
$\leq M^{2k+2}\rho(X,\Delta)+(M^{2k+1}+M^{2k-1}+\dots+M)\cdot(M-1)\cdot d(X,\Delta)$
$\leq M^{2k+2}(\rho(X,\Delta)+d(X,\Delta)),$
which gives the second inequality. ∎
Proof of Corollary 6.9.
Since $\delta_{M}(X,\Delta)\in\mathbb{N}$ by definition, we deduce from Theorem 6.8 that the MMP terminates in at most $\delta_{M}(X,\Delta)$ steps, where
$M:=\left(2+\left\lceil\frac{1}{b_{k}}\right\rceil\cdot\left\lceil\frac{1}{1-b_{1}}\right\rceil\right)\cdot|S|.$
The result then follows from Lemma 6.11. ∎
6.3. MMP with discrete discrepancies
Our next goal is to give explicit bounds for klt MMP in dimension $3$, or in dimension $4$ when one of $K_{X}+\Delta$ or $\Delta$ is big. A common feature in these cases is that the log discrepancies belong to some fixed finite set during the MMP, see Section 3. To give a more uniform treatment, we first prove an explicit termination statement by fixing this set. See Definition 6.5 for the definition of $s(X,\Delta)$ and $e(X,\Delta)$ that appear in the statement below.
Lemma 6.12.
Let $S\subseteq[0,1]$ be a finite set such that $0,1\in S$ and for any $a,b\in S$ with $a+b\leq 1$, we have $a+b\in S$. Let $(X,\Delta)$ be a klt pair of dimension $n\in\{3,4\}$. Let
[…]
be an $(K_{X}+\Delta)$-MMP such that for any divisor $E$ over $X$ and any $i$ with $a(E,X_{i},\Delta_{i})<0$, we have $-a(E,X_{i},\Delta_{i})\in S$. Then this MMP terminates after at most
$M^{2(k+1)^{2}|S|}s(X,\Delta)$
steps, where $k=e_{+}(X,\Delta)$ and $M=\left(2+\left\lceil\frac{1}{\min S\setminus\{0\}}\right\rceil\cdot\left\lceil\frac{1}{1-\max S\setminus\{1\}}\right\rceil\right)\cdot|S|$.
Before proving this we make an elementary observation.
Lemma 6.13.
Let $(X,\Delta)$ be a terminal pair and let $0\leq D\leq\Delta$ be an $\mathbb{R}$-divisor such that $K_{X}+D$ is $\mathbb{R}$-Cartier. Then $\rho(X,D)\leq\rho(X,\Delta)$ and $d(X,D)\leq d(X,\Delta)$.
Proof.
As $0\leq D\leq\Delta$, every component of $D$ is also a component of $\Delta$, hence the first inequality is clear from the definition. Since $a(E,X,\Delta)\leq a(E,X,D)$ holds for any prime divisor $E$ over $X$, to prove the second inequality it suffices to show that if $E$ is an exceptional divisor with $a(E,X,D)<1$ and it is not an echo of $(X,D)$, then it is also not an echo of $(X,\Delta)$. Suppose that $E$ is an echo for $(X,\Delta)$, then its center $\xi$ is a codimension one smooth point of $\mathrm{Supp}(\Delta)$. Note that in this case $\xi\in\mathrm{Supp}(D)$, otherwise $a(E,X,D)=a(E,X)\geq 1$ since $X$ is terminal and hence smooth at $\xi$. It follows that $\xi$ is also a codimension one smooth point of $\mathrm{Supp}(D)$ and hence an echo for $(X,D)$, a contradiction. This gives the second inequality. ∎
Proof of Lemma 6.12.
Let $\pi\colon(Y,D)\to(X,\Delta)$ be a projective birational morphism from a $\mathbb{Q}$-factorial terminal pair such that $K_{Y}+D=\pi^{*}(K_{X}+\Delta)+E$ for some $\pi$-exceptional $\mathbb{R}$-divisor $E\geq 0$. In other words, $\mathrm{mult}_{F}(D)\geq\min\{0,-a(F,X,\Delta)\}$ for all prime divisors $F$ on $Y$. We need to prove that the MMP (6.16) terminates after at most $M^{2(k+1)^{2}|S|}(\rho(Y,D)+d(Y,D))$ steps. By Lemma 6.13, we may decrease the coefficients of $D$ and assume that
$\mathrm{mult}_{F}(D)=\min\{0,-a(F,X,\Delta)\}$
for all prime divisors $F$ on $Y$. In particular, we may assume that $\mathrm{Coeff}(D)\subseteq S$.
By Lemma 2.17, we can lift the given $(K_{X}+\Delta)$-MMP to a diagram
[…]
where each $\pi_{i}$ ($i\geq 1$) is a $\mathbb{Q}$-factorial terminalization of $(X_{i},\Delta_{i})$ and each $\psi_{i}$ ($i\geq 0$) is composed of a sequence of $(K_{Y_{i}}+B_{i})$-MMP where $B_{i}=(\psi_{i}^{-1})_{*}D_{i+1}\leq D_{i}$. In particular, $\psi_{i}$ does not contract any component of $B_{i}$. Let $\delta_{i}:=\delta_{M}(Y_{i},D_{i})$ which is defined by (6.8). At each step we have two possibilities:
  1. (1)
    $B_{i}=D_{i}$. In this case $\delta_{i+1}\leq\delta_{i}-1$ by Theorem 6.8.
  2. (2)
    $B_{i}\neq D_{i}$. By Lemma 6.13 and Lemma 6.11, we have $\delta_{M}(Y_{i},B_{i})\leq M^{2k+2}\delta_{M}(Y_{i},D_{i})$, hence Theorem 6.8 gives $\delta_{i+1}\leq\delta_{M}(Y_{i},B_{i})\leq M^{2k+2}\delta_{i}$ (note that in this case it might be possible that $Y_{i+1}=Y_{i}$).
By construction $\mathrm{Supp}(D_{i})$ has only $k$ components, and $\mathrm{Coeff}(B_{i})\cup\mathrm{Coeff}(D_{i})\subseteq S$ for all $i$, hence the second case can only happen at most $k\cdot|S|$ times since we always have $0\leq B_{i}\leq D_{i}$. If we rescale the sequence $\delta_{0},\delta_{1},\dots$ by setting $\tilde{\delta}_{i}:=M^{(2k+2)\ell_{i}}\delta_{i}$ where $\ell_{i}=\#\{j\geq i\,|\,B_{j}\neq D_{j}\}$, then the above analysis gives $\tilde{\delta}_{i+1}\leq\tilde{\delta}_{i}$ for all $i$ and the equality holds at most $k\cdot|S|$ times. From here we see that the MMP (6.16) must terminate after at most
$\tilde{\delta}_{0}\leq M^{(2k+2)k\cdot|S|}\delta_{0}+k\cdot|S|\leq M^{2(k+1)^{2}|S|}(\rho(Y,D)+d(Y,D))$
steps, where the last inequality is by Lemma 6.11. This proves the desired result. ∎
We now put together the results obtained in this section and give some concrete applications. We shall only consider MMP that starts with a log smooth pair, since for any klt pair $(X,\Delta)$, we can always choose a log resolution $\pi\colon Y\to X$ and a log smooth klt pair $(Y,D)$ such that $K_{Y}+D\geq\pi^{*}(K_{X}+\Delta)+E$ for some $\pi$-exceptional $\mathbb{R}$-divisor $E\geq 0$, and any $(K_{X}+\Delta)$-MMP can be lifted to a $(K_{Y}+D)$-MMP by Lemma 2.15. Moreover, if $K_{X}+\Delta$ (resp. $\Delta$) is big, then we can choose the log smooth model $(Y,D)$ so that $K_{Y}+D$ (resp. $D$) is big.
The first application is about fourfold general type MMPs.
Theorem 6.14.
Let $N$ be a positive integer and let $(X,\Delta)$ be a projective log smooth klt pair of dimension $4$ such that $N\Delta$ has integer coefficients. Assume that
  1. (1)
    $(X,\Delta)$ has at most $N$ strata, each stratum has Picard number at most $N$, and $h_{4}^{\mathrm{alg}}(X)\leq N$,
  2. (2)
    $K_{X}+\Delta$ is big, $\mathrm{vol}(K_{X}+\Delta)\geq\frac{1}{N}$, and $((K_{X}+\Delta)\cdot H^{3})\leq N$ for some very ample divisor $H$ on $X$.
Then any $(K_{X}+\Delta)$-MMP terminates after at most $M^{M}$ steps, where $M:=(2N)^{9}!$.
Proof.
By Corollary 3.20, we know that in any $(K_{X}+\Delta)$-MMP, the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor is at most $(2^{8}N^{9})!$. Since $N\Delta$ has integer coefficients, we see that the assumptions of Lemma 6.12 are satisfied if we set
$S:=[0,1]\cap\frac{1}{N\cdot(2^{8}N^{9})!}\mathbb{R}.$
Let $N_{1}:=N\cdot(2^{8}N^{9})!$. Then $|S|=N_{1}+1$ and $\min(S\setminus\{0\})=1-\max(S\setminus\{1\})=\frac{1}{N_{1}}$. By Lemma 6.6, we also have $e(X,\Delta)\leq N^{5}-1$ and $s(X,\Delta)\leq 3N^{11}$. Thus by Lemma 6.12 we see that the MMP terminates after at most
$(2N_{1}+N_{1}^{3})^{2N^{10}(N_{1}+1)}\cdot 3N^{11}\leq M^{M}$
steps. ∎
The same argument also works when the boundary $\Delta$ is big.
Theorem 6.15.
Let $N$ be a positive integer and let $(X,\Delta)$ be a projective log smooth klt pair of dimension $4$ such that $N\Delta$ has integer coefficients. Assume that
  1. (1)
    $(X,\Delta)$ has at most $N$ strata, each stratum has Picard number at most $N$, and $h_{4}^{\mathrm{alg}}(X)\leq N$,
  2. (2)
    $\Delta$ is big, $\mathrm{vol}(\Delta)\geq\frac{1}{N}$, and $(\Delta\cdot H^{3})\leq N$ for some very ample divisor $H$ on $X$.
Then any $(K_{X}+\Delta)$-MMP terminates after at most $M^{M}$ steps, where $M:=(2N)^{9}!$.
Proof.
This follows from the same proof of Theorem 6.14, replacing Corollary 3.20 by Corollary 3.18. ∎
Finally, we give an explicit bound on the length of threefold MMPs. Is is not clear if such an explicit termination should be expected in dimension $\geq 4$, see [Sho-mld-conj, Example 12] for some discussion.
Theorem 6.16.
Let $N$ be a positive integer and let $(X,\Delta)$ be a projective log smooth klt pair of dimension $3$ such that $N\Delta$ has integer coefficients, $(X,\Delta)$ has at most $N$ strata, and each stratum has Picard number at most $N$. Then any $(K_{X}+\Delta)$-MMP terminates after at most
$3(N\cdot M!)^{8N^{8}(M!+1)}$
steps, where $M:=((N^{5}+2N)!)^{2^{N^{4}}}$.
Proof.
Since $N\Delta$ has integer coefficients, any positive discrepancy of $(X,\Delta)$ is at least $\frac{1}{N}$. For smooth threefolds we also have $h^{\mathrm{alg}}_{2}(X)=\rho(X)$ (indeed, the invariant $h^{\mathrm{alg}}_{2}(X)$ does not play any role in proving the termination of $3$-fold MMP, but we don’t need this fact). By Lemma 6.6, we see that $e(X,\Delta)\leq N^{4}-1$ and $s(X,\Delta)\leq 3N^{9}$. Thus by Corollary 3.22, the Cartier index of any $\mathbb{Q}$-Cartier Weil divisor is at most $M$. Then the assumptions of Lemma 6.12 are satisfied by $S:=[0,1]\cap\frac{1}{N\cdot M!}\mathbb{R}$, and it follows that the MMP terminates after at most
$((2+(N\cdot M!)^{2})\cdot N\cdot M!)^{2N^{8}(M!+1)}\cdot 3N^{9}\leq 3(N\cdot M!)^{8N^{8}(M!+1)}$
steps. ∎
Appendix A Boundedness of models implies termination
In this appendix, we prove that if the pairs that appear in a log canonical MMP belong to a bounded family, then the MMP must terminate after finitely many steps. As a corollary, we show that in any sequence of a log canonical MMP the pairs are mutually non-isomorphic, answering [Kol-KM98exercise]*Problem 82.
Theorem A.1.
Let $(X,\Delta)$ be a log canonical pair and let
[…]
be a sequence of steps of a $(K_{X}+\Delta)$-MMP. Assume that the set of pairs $\{(X_{i},\Delta_{i})\,|\,i\in\mathbb{N}\}$ is bounded. Then this MMP terminates after finitely many steps.
Here the boundedness assumption means that there exists an $\mathbb{R}$-Gorenstein family of pairs (see Definition 2.22) $(\mathcal{X},\mathcal{D})\to\mathcal{B}$ that contains every $(X_{i},\Delta_{i})$ as a fiber.
Corollary A.2.
Let $(X,\Delta)$ be a log canonical pair and let $(X,\Delta)\dashrightarrow(X^{\prime},\Delta^{\prime})$ be a sequence of steps of some $(K_{X}+\Delta)$-MMP. Then $(X^{\prime},\Delta^{\prime})$ is not isomorphic to $(X,\Delta)$.
Before we give the proof, let us recall Batyrev’s stringy $E$-function (see e.g. [Batyrev-stringy-E-function]*Definition 3.7 or [Veys-stringy-survey]*Definition 7.7). First let $(X,\Delta)$ be a klt log smooth sub-pair. Let $D=\cup_{i\in I}D_{i}$ be a simple normal crossing divisor on $X$ that contains every irreducible component of $\mathrm{Supp}(\Delta)$, and write $\Delta=\sum_{i\in I}a_{i}D_{i}$. Note that $a_{i}>-1$ for all $i$ by the klt assumption. Then we define the stringy $E$-function of the sub-pair $(X,\Delta)$ as
$E_{\mathrm{st}}(X,\Delta;u,v):=\sum_{J\subseteq I}H(D_{J}^{\circ};u,v)\prod_{j\in J}\frac{uv-1}{(uv)^{1+a_{j}}-1},$
where $D_{J}^{\circ}:=\cap_{j\in J}D_{j}\setminus(\cup_{j\not\in J}D_{j})$ (with the understanding that $D_{\emptyset}^{\circ}=X\setminus\mathrm{Supp}(D)$ and $\prod_{j\in J}=1$ if $J=\emptyset$), and $H(D_{J}^{\circ};u,v)$ is the Hodge-Deligne polynomial of $D_{J}^{\circ}$, which is defined by
$H(U;u,v):=\sum_{i,p,q}(-1)^{i}h^{p,q}(H^{i}_{c}(U,\mathbb{C}))u^{p}v^{q}$
for any algebraic variety $U$. It is not hard to see that the definition of the stringy $E$-function is independent of the choice of $D$.
In general, let $(X,\Delta)$ be a log sub-pair, let $\mathfrak{a}\subseteq\mathcal{O}_{X}$ be an ideal sheaf, and let $c\in\mathbb{R}$ be such that $(X,\Delta+\mathfrak{a}^{c})$ is klt, in the sense that
$A_{X,\Delta+\mathfrak{a}^{c}}(F):=A_{X,\Delta}(F)-c\cdot\mathrm{ord}_{E}(\mathfrak{a})>0$
for all prime divisor $F$ over $X$. Note that we allow $c$ and the coefficient of $\Delta$ to be negative. Let $\pi\colon(Y,D)\to(X,\Delta+\mathfrak{a})$ be a crepant log resolution. In other words,
$\mathrm{Ex}(\pi)+\mathrm{Supp}(\pi^{-1}_{*}\Delta)+\pi^{-1}\mathrm{Supp}(\mathcal{O}_{X}/\mathfrak{a})$
is a simple normal crossing divisor on $Y$, and
$-\mathrm{mult}_{F}D=a(F,X,\Delta+\mathfrak{a}^{c}):=a(F,X,\Delta)-c\cdot\mathrm{ord}_{F}(\mathfrak{a})$
for every prime divisor $F$ on $Y$. Then we define the stringy $E$-function of the klt sub-pair $(X,\Delta+\mathfrak{a}^{c})$ as
$E_{\mathrm{st}}(X,\Delta+\mathfrak{a}^{c};u,v):=E_{\mathrm{st}}(Y,D;u,v).$
By [Batyrev-stringy-E-function]*Theorems 3.6 and 3.8, the definition is independent of the crepant log resolution $(Y,D)$.
Lemma A.3.
Let $(X,\Delta+\mathfrak{a}^{c})$ and $(X^{\prime},\Delta^{\prime}+\mathfrak{b}^{d})$ be klt sub-pairs and let $\varphi\colon X\dashrightarrow X^{\prime}$ be a birational map. Assume that:
  1. (1)
    There exist proper birational morphisms $f\colon Y\to X$ and $f^{\prime}\colon Y\to X^{\prime}$ such that $f^{\prime}=\varphi\circ f$, and
  2. (2)
    $A_{X,\Delta+\mathfrak{a}^{c}}(F)\leq A_{X^{\prime},\Delta^{\prime}+\mathfrak{b}^{d}}(F)$ for all divisors $F$ over $X$ and the inequality is strict for some $F$.
Then for $u,v\gg 0$ we have
$E_{\mathrm{st}}(X,\Delta+\mathfrak{a}^{c};u,v)>E_{\mathrm{st}}(X^{\prime},\Delta^{\prime}+\mathfrak{b}^{d};u,v).$
Proof.
Let $Y$ be a common log resolution of $(X,\Delta+\mathfrak{a}^{c})$ and $(X^{\prime},\Delta^{\prime}+\mathfrak{b}^{d})$ and let $D=\cup_{i\in I}D_{i}$ be an SNC divisor on $Y$ that contains the strict transforms of $\Delta$ and $\Delta^{\prime}$, the preimage of $\mathrm{Supp}(\mathcal{O}_{X}/\mathfrak{a})$ and $\mathrm{Supp}(\mathcal{O}_{X^{\prime}}/\mathfrak{b})$, and all the exceptional divisors over $X$ or $X^{\prime}$. Let $a_{i}=A_{X,\Delta+\mathfrak{a}^{c}}(D_{i})$ and $a^{\prime}_{i}=A_{X^{\prime},\Delta^{\prime}+\mathfrak{b}^{d}}(D_{i})$. By assumption we have $a^{\prime}_{i}\geq a_{i}>0$ and the inequality is strict for some $i$. Without loss of generality we may assume that $a^{\prime}_{1}>a_{1}$. Note that the Hodge-Deligne polynomials have positive leading coefficients by [PS-mix-Hodge, Theorem 5.39]. Then for any $u,v\gg 0$ and any $J\subseteq I$, we have
$H(D_{J}^{\circ};u,v)\prod_{j\in J}\frac{uv-1}{(uv)^{a_{j}}-1}\geq H(D_{J}^{\circ};u,v)\prod_{j\in J}\frac{uv-1}{(uv)^{a^{\prime}_{j}}-1}.$
Moreover, the inequality is strict when $J=\{1\}$. By the definition of the stringy $E$-function, this implies that
$E_{\mathrm{st}}(X,\Delta+\mathfrak{a}^{c};u,v)>E_{\mathrm{st}}(X^{\prime},\Delta^{\prime}+\mathfrak{b}^{d};u,v)$
when $u,v\gg 0$. ∎
Lemma A.4.
Let $\overline{\mathcal{X}}\to\mathcal{B}$ be a projective morphism. Let $\overline{\mathcal{D}}$ be an $\mathbb{R}$-divisor on $\overline{\mathcal{X}}$ such that $(\overline{\mathcal{X}},\overline{\mathcal{D}})$ is log smooth and klt, and every stratum of $(\overline{\mathcal{X}},\overline{\mathcal{D}})$ is smooth over $\mathcal{B}$. Let $\mathcal{X}\subseteq\overline{\mathcal{X}}$ be an open subset that is a union of some open strata of $(\overline{\mathcal{X}},\overline{\mathcal{D}})$, and let $\mathcal{D}=\overline{\mathcal{D}}|_{\mathcal{X}}$. Then the stringy $E$-function $E_{\mathrm{st}}(\mathcal{X}_{b},\mathcal{D}_{b};u,v)$ is constant in $b\in\mathcal{B}$.
Proof.
It suffices to show that for any open stratum $\mathcal{W}^{\circ}$ of $(\overline{\mathcal{X}},\overline{\mathcal{D}})$, the Hodge-Deligne polynomial $H(\mathcal{W}^{\circ}_{b};u,v)$ of $\mathcal{W}^{\circ}_{b}:=\mathcal{W}^{\circ}\cap\overline{\mathcal{X}}_{b}$ is independent of $b\in\mathcal{B}$. We prove this by Noetherian induction. To this end, let $\mathcal{W}:=\overline{\mathcal{W}^{\circ}}$ which is a stratum of $(\overline{\mathcal{X}},\overline{\mathcal{D}})$. There is nothing to prove when $\mathcal{W}=\emptyset$. For the inductive step, note that by Deligne’s mixed Hodge theory, we have
(A.1) $H(X;u,v)=H(X\setminus Z;u,v)+H(Z;u,v)$
whenever $Z$ is a closed subvariety of $X$ (see e.g. [Sri-Hodge, Theorem 1.1]). Since $\mathcal{W}\setminus\mathcal{W}^{\circ}$ is a union of open strata of smaller dimensions, we deduce from the induction hypothesis and (A.1) that $H(\mathcal{W}_{b}\setminus\mathcal{W}^{\circ}_{b};u,v)$ is constant in $b$. By assumption, $\mathcal{W}$ is smooth over $\mathcal{B}$, hence $H(\mathcal{W}_{b};u,v)$ is also constant in $b$. Thus another application of (A.1) implies that $H(\mathcal{W}^{\circ}_{b};u,v)$ is constant in $b$. This completes the proof. ∎
Proof of Theorem A.1.
First assume that $(X,\Delta)$ is klt. By the negativity lemma, we know that for any $i\in\mathbb{N}$ we have $a(E,X_{i},\Delta_{i})\leq a(E,X_{i+1},\Delta_{i+1})$ for all divisors $E$ over $X$, and the inequality is strict for some $E$. By Lemma A.3, we see that
$E_{\mathrm{st}}(X_{i},\Delta_{i};u,v)>E_{\mathrm{st}}(X_{i+1},\Delta_{i+1};u,v)$
when $u,v\gg 0$. It follows that
(A.2) $E_{\mathrm{st}}(X_{i},\Delta_{i};u,v)\neq E_{\mathrm{st}}(X_{j},\Delta_{j};u,v)$
whenever $i\neq j$. If there is an $\mathbb{R}$-Gorenstein family of pairs $(\mathcal{X},\mathcal{D})\to\mathcal{B}$ that contains every $(X_{i},\Delta_{i})$ as a fiber, then after possibly stratifying $\mathcal{B}$, we may assume that there is a crepant log resolution $\pi\colon(\mathcal{Y},\mathcal{G})\to(\mathcal{X},\mathcal{D})$, a projective morphism $\overline{\mathcal{Y}}\to\mathcal{B}$ with an open immersion $\mathcal{Y}\subseteq\overline{\mathcal{Y}}$ over $\mathcal{B}$, and an $\mathbb{R}$-divisor $\overline{\mathcal{G}}$ on $\overline{\mathcal{Y}}$ such that $\mathcal{G}=\overline{\mathcal{G}}|_{\mathcal{Y}}$, the sub-pair $(\overline{\mathcal{Y}},\overline{\mathcal{G}})$ is log smooth, every stratum of $(\overline{\mathcal{Y}},\overline{\mathcal{G}})$ is smooth over $\mathcal{B}$, and $\mathcal{Y}$ is a union of some open strata of $(\overline{\mathcal{Y}},\overline{\mathcal{G}})$. By Lemma A.4, we see that the stringy $E$-function $E_{\mathrm{st}}(\mathcal{Y}_{b},\mathcal{G}_{b};u,v)$ ($b\in\mathcal{B}$) is constant on each connected component of $\mathcal{B}$, hence there are only finitely many possible stringy $E$-functions $E_{\mathrm{st}}(X_{i},\Delta_{i};u,v)$. It follows that the MMP must terminate in finitely many steps, otherwise we get a contradiction to (A.2).
In general, let $W_{i}$ be the reduced non-klt locus of $(X_{i},\Delta_{i})$. We claim that for any $i\in\mathbb{N}$ and any sufficiently small $0<\varepsilon\ll 1$ (which may depend on $i$), we have
(A.3) $a(E,X_{i},\Delta_{i}-\mathcal{I}_{W_{i}}^{\varepsilon})\leq a(E,X_{i+1},\Delta_{i+1}-\mathcal{I}_{W_{i+1}}^{\varepsilon})$
for all divisors $E$ over $X$ and the inequality is strict for some $E$. It is enough to prove this for prime divisors on some common log resolution $Y$ of $(X_{i},\Delta_{i}+\mathcal{I}_{W_{i}})$ and $(X_{i+1},\Delta_{i+1}+\mathcal{I}_{W_{i+1}})$. Let $E$ be one such divisor. If the center of $E$ on $X_{i}$ is contained in the exceptional locus of the birational map $X_{i}\dashrightarrow X_{i+1}$ (there is at least one such divisor), then
$a(E,X_{i},\Delta_{i})<a(E,X_{i+1},\Delta_{i+1})$
by the negativity lemma, hence the strict inequality in (A.3) holds for sufficiently small $\varepsilon$. As there are only finitely many such divisors $E$ on $Y$, we can make a uniform choice of $\varepsilon$. If the center $\xi=c_{X_{i}}(E)$ of $E$ on $X_{i}$ is not contained in the exceptional locus, then the map $X_{i}\dashrightarrow X_{i+1}$ is an isomorphism at $\xi$ and identifies the non-klt locus $W_{i}$ with $W_{i+1}$ around $\xi$. It follows that $a(E,X_{i},\Delta_{i})=a(E,X_{i+1},\Delta_{i+1})$ and $\mathrm{ord}_{E}(\mathcal{I}_{W_{i}})=\mathrm{ord}_{E}(\mathcal{I}_{W_{i+1}})$, hence (A.3) also holds. This proves the claim.
Note that the non-klt locus $W_{i}$ is determined by the pair $(X_{i},\Delta_{i})$. If there exists an $\mathbb{R}$-Gorenstein family $(\mathcal{X},\mathcal{D})\to\mathcal{B}$ that contains every pair $(X_{i},\Delta_{i})$ as a fiber, then possibly after stratifying $\mathcal{B}$, we may assume that the non-klt locus $\mathcal{W}$ of $(\mathcal{X},\mathcal{D})$ restricts to $W_{i}$ on $X_{i}$. By passing to log resolutions and compactifications over $\mathcal{B}$ as in the klt case and using Lemma A.4, we deduce that there are only finitely many possibilities for the functions $E_{\mathrm{st}}(X_{i},\Delta_{i}-\mathcal{I}_{W_{i}}^{\varepsilon};u,v)$ (in the variables $\varepsilon,u,v$). On the other hand, by (A.3) and Lemma A.3 we also have
$E_{\mathrm{st}}(X_{i},\Delta_{i}-\mathcal{I}_{W_{i}}^{\varepsilon};u,v)>E_{\mathrm{st}}(X_{i+1},\Delta_{i+1}-\mathcal{I}_{W_{i+1}}^{\varepsilon};u,v)$
when $0<\varepsilon\ll 1$ and $u,v\gg 0$, hence
$E_{\mathrm{st}}(X_{i},\Delta_{i}-\mathcal{I}_{W_{i}}^{\varepsilon};u,v)\neq E_{\mathrm{st}}(X_{j},\Delta_{j}-\mathcal{I}_{j}^{\varepsilon};u,v)$
whenever $i\neq j$. Thus the MMP terminates in finitely many steps by the same argument as in the klt case. ∎
Proof of Corollary A.2.
If $(X^{\prime},\Delta^{\prime})\cong(X,\Delta)$, then by repeating the same sequence of $(K_{X}+\Delta)$-MMP steps, we get an infinite sequence of $(K_{X}+\Delta)$-MMP that only involves finitely many isomorphism classes of pairs. This contradicts Theorem A.1. ∎